### v1

# On the Identifiability of Steering Vectors in Large Language Models

Abstract
Activation steering methods, such as persona vectors, are widely used to control large language model behavior and increasingly interpreted as revealing meaningful internal representations. This interpretation implicitly assumes steering directions are identifiable and uniquely recoverable from input–output behavior. We formalize steering as an intervention on internal representations and prove that, under realistic modeling and data conditions, steering vectors are fundamentally non-identifiable due to large equivalence classes of behaviorally indistinguishable interventions. Empirically, we validate this across multiple models and semantic traits, showing orthogonal perturbations achieve near-equivalent efficacy with negligible effect sizes. However, identifiability is recoverable under structural assumptions including statistical independence, sparsity constraints, multi-environment validation or cross-layer consistency. These findings reveal fundamental interpretability limits and clarify structural assumptions required for reliable safety-critical control.
On the Identifiability of Steering Vectors in Large Language Models
Sohan Venkatesh  and Ashish Mahendran Kurapath Manipal Institute of Technology Bengaluru {sohan1, ashish}.mitblr2022@learner.manipal.edu
1 Introduction
Persona vector steering has emerged as a popular technique for controlling the behavior of large language models by adding learned directional vectors to intermediate activations. Empirically, such vectors can shift model outputs along interpretable dimensions such as politeness, political ideology or truthfulness, suggesting that representational alignment might afford fine-grained behavioral control without retraining ((Zou et al., 2023; Rimsky et al., 2024; Turner et al., 2023)). This line of work is closely connected to broader efforts in representation engineering and activation editing, where linear directions in activation space are used to modulate model behavior along semantic axes ((Elhage et al., 2022; Turner et al., 2024)).
Despite growing adoption in interpretability and alignment research, the theoretical foundations of persona steering remain poorly understood. Most existing methods implicitly assume that extracted steering vectors correspond to meaningful, uniquely determined latent factors—for example, "the politeness direction" or "the honesty direction"—and that these factors can be directly manipulated to achieve reliable control. However, classical results in latent variable modeling and causal inference show that such assumptions are often unjustified without additional structural constraints ((Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Schölkopf et al., 2021; Locatello et al., 2019)). Nonlinear ICA and causal representation learning further emphasize that recovering latent factors from high-dimensional observations is generically impossible without auxiliary information or strong inductive biases ((Hyvarinen and Morioka, 2017; Khemakhem et al., 2020; Ahuja et al., 2022)).
This raises a fundamental question for alignment and interpretability research: when does representational alignment genuinely afford reliable behavioral control and when does it merely exploit underdetermined or spurious correlations? Existing evidence from probing and representation analysis suggests that even seemingly meaningful linear directions can reflect artifacts of the measurement procedure rather than uniquely defined semantic variables ((Hewitt and Liang, 2019; Ravfogel et al., 2020; Elazar et al., 2021; Belinkov, 2022)). In the context of activation steering, these concerns become particularly acute, because the resulting vectors are often used not only for analysis but also for safety-relevant control interventions. Understanding identifiability is therefore critical for several reasons:
Alignment affordances.
If persona vectors are not identifiable, then representational alignment may provide only heuristic control rather than principled intervention. Characterizing when steering vectors are unique clarifies when alignment interventions can be trusted and when they should instead be viewed as exploiting one of many behaviorally equivalent directions.
Interpretability validity.
When multiple incompatible vectors produce identical observable behavior, claims that a specific vector "represents" a semantic concept become scientifically underdetermined. Identifiability theory distinguishes well-grounded interpretability claims from artifacts of measurement and projection ((Elazar et al., 2021; Marks and Tegmark, 2023)).
Robustness and safety.
Non-identifiable steering directions may rely on fragile correlations that fail under distribution shift, model updates or adversarial prompting. For safety-critical applications—where steering vectors may be used to enforce norms or prevent harmful behavior—understanding identifiability limits is essential to avoid brittle or misleading forms of control.
Methodological design.
Identifiability theory clarifies which experimental protocols provide meaningful evidence and which require additional structure to support reliable conclusions. In particular, it highlights when interventions on internal activations can be interpreted causally and when they merely reparameterize an equivalence class of representations ((Schölkopf et al., 2021; Ahuja et al., 2022)).
In this work, we provide, to our knowledge, the first formal identifiability analysis of persona vector steering. Our contributions are threefold. First, we prove that under standard observational regimes, persona vectors are generically non-identifiable due to null-space ambiguity in the model’s input-output Jacobian (Proposition 1). Specifically, under white-box single-layer access, infinitely many geometrically distinct steering directions induce identical observable behavior. Second, we characterize sufficient structural conditions under which identifiability can be recovered including statistical independence constraints, sparsity priors, multi-environment access and cross-layer consistency (Proposition 2). Third, we demonstrate empirically that contemporary steering operates in the non-identifiable regime across multiple models and semantic traits, with vectors perturbed by orthogonal components achieving near-equivalent efficacy. Crucially, our results are not purely negative. By explicitly characterizing both the limits and the affordances of persona vector steering, we provide principled guidance for when representational alignment can be trusted and which structural assumptions enable reliable control in safety-critical applications.
2 Related Work
Our work formalizes persona steering as a latent variable identification problem, bridging causal representation learning, activation editing in LLMs and mechanistic interpretability.
Causal and latent variable identifiability.
Classical results show that latent variable models are generically non-identifiable without structural assumptions (Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Kruskal, 1977). Recent work in causal representation learning extends these ideas to deep learning settings (Schölkopf et al., 2021; Ahuja et al., 2022; Locatello et al., 2019), establishing conditions under which latent factors can be recovered from high-dimensional observations. Nonlinear ICA methods (Khemakhem et al., 2020; Hyvarinen and Morioka, 2017) provide theoretical foundations for recovering latent variables using temporal or auxiliary information, which standard steering methods do not exploit.
Probing and representation learning.
Pimentel et al. (2020) and Ravfogel et al. (2020) demonstrate that probing classifiers can succeed for trivial reasons unrelated to the presence of target information. Elazar et al. (2021) show that removing information via projection is ill-defined without identifiability guarantees. Belinkov (2022) provides a comprehensive survey of probing methods and their limitations. The linear representation hypothesis (Park et al., 2023; Elhage et al., 2022; Marks and Tegmark, 2023) suggests that concepts correspond to directions in activation space but lacks formal identifiability guarantees.
Activation editing in LLMs.
Methods such as representation engineering (Zou et al., 2023), contrastive activation addition (Rimsky et al., 2024; Turner et al., 2024), activation patching (Meng et al., 2022) and causal tracing (Wang et al., 2022) manipulate model internals to control behavior. While empirically successful, these works generally do not address whether the resulting control directions are uniquely determined. Related work on linear mode connectivity (Entezari et al., 2021) and neural network reparameterization (Dinh et al., 2017) reveals symmetries in neural representations that motivate our null-space analysis.
Steering and control vectors.
Early work on steering vectors (Li et al., 2022; Turner et al., 2023; Tigges et al., 2023) develops techniques for extracting control directions from contrastive prompt pairs. Burns et al. (2022) investigate discovering latent knowledge in language models. These methods assume that extracted vectors correspond to uniquely determined semantic factors—an assumption we formally examine.
Mechanistic interpretability.
Work on mechanistic interpretability (Olsson et al., 2022; Elhage et al., 2021; Nanda et al., 2023) studies circuits and features in transformers, often assuming the existence of interpretable directions without formalizing identifiability constraints. Our theoretical analysis clarifies when such directions are well-defined.
Compressed sensing and sparse recovery.
The theoretical framework of compressed sensing (Candès and Wakin, 2008; Donoho, 2006) provides conditions under which sparse vectors can be uniquely recovered from linear measurements. These results directly inform the sparsity-based identifiability conditions in Proposition 2.
3 Problem Setup
3.1 Formal Model
Consider a pre-trained transformer language model $f_{\theta}$ with $L$ layers. For a given input prompt $x$ (tokenized as $x_{1},\ldots,x_{T}$), let $h_{\ell}(x)\in\mathbb{R}^{d}$ denote the hidden representation at layer $\ell$ and position $T$ (typically the final token position for autoregressive generation).
Latent persona variable.
We assume there exists an underlying latent variable $z\in\mathcal{Z}$ representing a semantic attribute or "persona" (e.g., formality, political stance, truthfulness).
Steering intervention.
A steering vector $v\in\mathbb{R}^{d}$ is applied as:
$\tilde{h}\ell(x)=h\ell(x)+\alpha v,$
where $\alpha\in\mathbb{R}$ is the steering strength. The modified representation $\tilde{h}_{\ell}$ is fed forward through subsequent layers to produce output logits $o(x,v,\alpha)$ over the vocabulary.
Generative model.
We posit that the true data-generating process involves:
$z\sim p(z),\quad h_{\ell}=g_{\ell}(x,z)+\epsilon,$
where $g_{\ell}:\mathcal{X}\times\mathcal{Z}\to\mathbb{R}^{d}$ encodes how persona $z$ modulates the representation for prompt $x$ and $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ is measurement noise. The goal of steering is to approximate the effect of varying $z$ by adding $v$.
3.2 Observational Regimes
We consider two data access regimes that determine what alignment interventions can afford:
  • Regime 1: Black-box input–output. The researcher observes only $(x,y)$ pairs, where $y$ is generated text. There is no access to internal representations. This is the weakest regime and corresponds to behavioral evaluation.
  • Regime 2: White-box single-layer access. The researcher can observe or manipulate activations $h_{\ell}(x)$ at a chosen layer $\ell$. This is the standard setting for most steering work and includes extracting vectors from contrastive prompt pairs.
Most existing work operates in Regime 2, often extracting a steering vector from contrastive prompt pairs $(x^{+},x^{-})$ designed to elicit different persona values (e.g., polite versus rude instructions):
$v\propto\mathbb{E}{x^{+}}[h\ell(x^{+})]-\mathbb{E}{x^{-}}[h\ell(x^{-})].$
Our analysis focuses primarily on Regime 2, as it represents the dominant paradigm in current steering research.
3.3 Linear Approximation and Nonlinear Case
Local linearization.
Near a reference distribution, we can approximate the effect of steering on output logits as:
$o(x,v,\alpha)\approx o(x,0,0)+\alpha J_{\ell}(x)v,$
where $J_{\ell}(x)=\frac{\partial o}{\partial h_{\ell}}\big|{h\ell(x)}\in\mathbb{R}^{V\times d}$ is the Jacobian and $V$ is the vocabulary size.
Nonlinear case.
In general, the mapping $h_{\ell}\mapsto o$ involves multiple nonlinear layers (attention, MLPs, layer norms). We denote this as:
$o=F_{\ell\to L}(h_{\ell}+\alpha v),$
where $F_{\ell\to L}$ is the composition of layers $\ell+1$ through $L$.
3.4 Assumptions
We make the following assumptions explicit:
A1 (Smoothness).
The functions $g_{\ell}$ and $F_{\ell\to L}$ are differentiable almost everywhere, enabling local linear approximation via Jacobians $J_{\ell}(x)=\frac{\partial o}{\partial h_{\ell}}(x)$.
A2 (Identifiable prompts).
The prompt distribution $p(x)$ has sufficient variability to probe different aspects of the latent persona $z$. Formally, the support of $p(x)$ is rich enough that different persona values induce distinguishable activation patterns $h_{\ell}(x)$.
A3 (Non-degeneracy).
The Jacobian $J_{\ell}(x)$ has rank at least $k\geq 1$ for typical $x\sim p(x)$, meaning steering can affect outputs. This excludes pathological cases where all perturbations to $h_{\ell}$ are ignored by subsequent layers.
We do not assume:
  • Statistical independence between $z$ and $x$ (confounding is allowed)
  • Linearity of $g_{\ell}$ in $z$
  • Uniqueness of the latent representation without additional structure
These assumptions are mild and satisfied by standard transformer architectures under typical operating conditions.
4 Definitions and Identifiability
4.1 Identifiability
Definition 1 (Parameter Identifiability).
A parameter $\theta$ in a statistical model $p(y\mid x;\theta)$ is identifiable if for any $\theta^{\prime}\neq\theta$, there exists a distribution over observations $(x,y)$ such that $p(y\mid x;\theta)\neq p(y\mid x;\theta^{\prime})$.
In our setting, the parameter is the steering vector $v\in\mathbb{R}^{d}$. We say $v$ is identifiable if no other vector $v^{\prime}\neq v$ (up to scaling) produces the same distribution over observable outputs across all prompts and steering strengths.
Definition 2 (Observational Equivalence).
Two steering vectors $v$ and $v^{\prime}$ are observationally equivalent in regime $\mathcal{R}$ if they produce identical distributions over all quantities observable in $\mathcal{R}$.
For Regime 2 (white-box single-layer access):
$v\sim_{\mathcal{R}}v^{\prime}\iff F_{\ell\to L}(h_{\ell}(x)+\alpha v)$ (1)
$=F_{\ell\to L}(h_{\ell}(x)+\alpha v^{\prime})\;\forall x,\alpha.$
4.2 Symmetries and Gauge Freedom
Scaling ambiguity.
For any $c\neq 0$, the vectors $v$ and $cv$ produce outputs that differ only by a rescaling of $\alpha$. This is unavoidable; we consider $v$ and $cv$ as the same direction.
Null space ambiguity.
If $v_{0}\in\ker(J_{\ell})$ (i.e., $J_{\ell}v_{0}=0$), then adding $v_{0}$ to any steering vector does not change the linearized output. Under linear approximation, $v$ and $v+v_{0}$ are observationally equivalent.
5 Main Results
We now state our main theoretical results, characterizing when observational conditions afford identifiable persona vectors and when they do not.
5.1 Proposition 1: Non-Identifiability Without Structural Constraints
Proposition 1.
Under Assumptions A1–A3, in Regime 2 (white-box single-layer access) without additional structural constraints, persona vectors are not identifiable. Specifically, for any steering vector $v\in\mathbb{R}^{d}$, there exist infinitely many vectors $v^{\prime}\not\propto v$ that are observationally equivalent.
Proof Sketch.
We establish non-identifiability via two complementary arguments: reparameterization symmetry and null-space ambiguity.
Reparameterization symmetry. Consider an invertible transformation $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and define the reparameterized representation $h^{\prime}\ell(x)=T(h\ell(x))$. If the subsequent layers can be equivalently expressed as $F_{\ell\to L}(h_{\ell})=F^{\prime}{\ell\to L}(T(h\ell))$, then using $(h_{\ell},v)$ or $(h^{\prime}_{\ell},v^{\prime})$ where $v^{\prime}=T(v)$ is observationally indistinguishable.
For overparameterized networks, many such reparameterizations exist that leave the input-output function invariant. Consider $T(h)=Ah+b$ where $A\in\mathbb{R}^{d\times d}$ is invertible. This transformation can be absorbed into the parameters of layer $\ell+1$ through weight reparameterization. While we cannot explicitly construct these transformations for a given frozen model without retraining, their existence follows from the network’s inherent symmetry structure. Let $A$ be any invertible matrix with $A\neq cI$ for any scalar $c$. Then $v^{\prime}=Av$ is not proportional to $v$, yet produces equivalent observations after appropriate reparameterization. Since there are infinitely many such $A$, there exist infinitely many equivalent $v^{\prime}$.
Null-space ambiguity (primary mechanism). Under the local linear approximation $o\approx o_{0}+\alpha J_{\ell}v$, any vector $v^{\prime}=v+v_{0}$ where $v_{0}\in\ker(J_{\ell})$ satisfies $J_{\ell}v^{\prime}=J_{\ell}v$ and thus produces identical linearized outputs. This establishes non-identifiability even without invoking global reparameterization symmetries.
Corollary 1.1.
Under the linear approximation $o\approx o_{0}+\alpha J_{\ell}v$, any vector $v^{\prime}=v+v_{0}$ where $v_{0}\in\ker(J_{\ell})$ is observationally equivalent to $v$.
Remark (Null-space dimensionality).
The null space $\ker(J_{\ell})$ is typically high-dimensional in practice. For a Jacobian $J_{\ell}\in\mathbb{R}^{V\times d}$ with vocabulary size $V$ and hidden dimension $d$, the maximum possible rank is $\min(V,d)$. In modern language models, $V\approx 50000$ and $d\approx 4000$, so max rank is $d$. However, the output distribution lies on a low-dimensional manifold, causing $\text{rank}(J_{\ell})\ll d$ in practice, consistent with observations that neural network representations concentrate on low-dimensional subspaces (Maennel et al., 2018; Li et al., 2018). This yields $\dim(\ker(J_{\ell}))=d-\text{rank}(J_{\ell})\gg 0$. This result establishes generic non-identifiability under local linear approximation.
5.2 Proposition 2: Identifiable Regimes Under Structural Assumptions
Proposition 2.
Persona vectors can be identified up to scaling and permutation, thus affording reliable alignment control, under the following sufficient structural conditions:
  • Statistical Independence (ICA). If the latent persona $z=(z_{1},\ldots,z_{k})$ has independent components and $h_{\ell}=Az+\epsilon$ where $A\in\mathbb{R}^{d\times k}$ is a mixing matrix, then under non-Gaussianity of $z_{i}$ and sufficient observations, AA (and hence the columns $v_{i}$) can be recovered up to permutation and scaling (Comon, 1994; Hyvärinen and Oja, 2000).
  • Sparsity Constraints. If the true persona vector $v$ is sparse (i.e., $|v|_{0}\leq s\ll d$) and we observe the effect of steering on multiple outputs, then under restricted isometry properties, $v$ can be uniquely recovered via $\ell_{1}$ minimization(Candes and Tao, 2005).
  • Multi-Environment or Interventional data. If we observe the same persona $z$ across multiple contexts (prompts, models or layers) where the spurious correlations change but the true signal remains stable, then techniques from causal representation learning allow identification of invariant factors. (Peters et al., 2017; Ahuja et al., 2022).
  • Cross-layer Consistency. If we assume persona vectors have consistent geometric relationships across multiple layers (e.g., $v_{\ell}\approx W_{\ell}v_{\ell-1}$ for known or estimable $W_{\ell}$), then the overdetermined system provides additional constraints that break symmetries.
Interpretation.
Proposition 2 serves as a natural theoretical extension of Proposition 1: by characterizing sufficient conditions under which identifiability could be recovered–such as statistical independence (ICA), sparsity priors, multi-environment access or cross-layer consistency. Proposition 2 clarifies precisely which structural assumptions are necessary to break the symmetries that cause non-identifiability.
6 Empirical Validation
We now validate that the non-identifiability characterized in Proposition 1 (Section 5.1) manifests in contemporary language models. Our empirical experiments test a behavioral consequence of the theoretical prediction: the existence of large equivalence classes of semantically indistinguishable steering directions, without directly estimating the Jacobian null space.
6.1 Experimental Setup
Models and Layers.
We evaluate two open-weight instruction-tuned language models to test generality across architectures and scales:
  • Qwen2.5-3B-Instruct: 24 layers, hidden dimension $d=2048$, layer $\ell=12$ (mid-network)
  • Llama-3.1-8B-Instruct: 32 layers, hidden dimension $d=4096$, layer $\ell=16$ (mid-network)
For both models, we focus on middle layers ($\ell=L/2$), following standard practice (Chen et al., 2025; Konen et al., 2024; Sun et al., ) in steering research. Middle layers balance semantic abstraction with steerability: early layers encode low-level features (tokens, syntax), while late layers specialize for next-token prediction.
Persona traits.
We test three semantic traits spanning distinct dimensions:
  • Formality: Formal versus informal style
  • Politeness: Polite versus rude social register
  • Humor: Humorous versus serious content
This selection ensures our findings are not specific to a single semantic dimension but reflect a general property of steering vector geometry.
Steering vector extraction.
For each trait, we construct 50 contrastive prompt pairs designed to elicit contrasting persona values. For example, formality pairs contrast "Write a professional and formal message about [topic]" versus "Write a casual and informal message about [topic]." For each prompt pair, we extract the hidden representation at layer $\ell$ for the final token position (Chen et al., 2025). The baseline steering vector is computed as the mean difference:
$v=\frac{1}{50}\sum_{i=1}^{50}[h_{\ell}(x^{+}i)-h\ell(x^{-}_{i})].$
Semantic probes.
We evaluate semantic equivalence using trait-specific scoring functions $\phi(o)$ that map generated text to real-valued scores. This provides a conservative test of observational equivalence: if vectors are distinguishable semantically, they are certainly non-identical in the full distributional sense.
For formality, politeness and humor, we use lexical heuristics based on formal/informal markers, polite/rude markers and humorous/serious markers respectively, combined with sentence length and stylistic features. These return scores in $[0,1]$ where 1 is maximally formal/polite/humorous.
Orthogonality test methodology.
To test Proposition 1’s prediction that $v$ and $v+v_{\perp}$ (where $v_{\perp}$ is orthogonal) produce observationally equivalent outputs, we implement the following procedure:
  1. 1.
    Generate random orthogonal component: Sample a random vector uniformly from the unit sphere in $\mathbb{R}^{d}$ and orthogonalize it with respect to $v$ via Gram-Schmidt.
  2. 2.
    Construct perturbed vector: Form $v^{\prime}=v+\alpha v_{\perp}$ where $\alpha$ is chosen such that $|v^{\prime}|\approx|v|$.
  3. 3.
    Generate steered outputs: For each of 100 held-out test prompts, generate text with steering vectors $v$ and $v^{\prime}$ at strength $\alpha=1.0$, producing 10 samples per prompt per vector.
  4. 4.
    Compute semantic equivalence: Measure Cohen’s $d$ effect size and Pearson correlation between semantic scores $\phi(o_{v})$ and $\phi(o_{v^{\prime}})$.
  5. 5.
    Repeat across seeds: Repeat for multiple random orthogonal seeds to assess robustness.
If $v$ and $v^{\prime}$ are observationally equivalent as predicted by Proposition 1, we expect Cohen’s $d<0.2$ (negligible effect) and high correlation between semantic scores.
Sample size design.
We conduct the orthogonality test with both $n=5$ and $n=10$ random orthogonal seeds per trait for Qwen2.5-3B and Llama-3.1-8B. This allows us to assess statistical stability across sample sizes and models.
Scale invariance test.
To verify that observational equivalence holds across different steering strengths, we additionally evaluate the formality trait at four $\alpha$ values: 0.0,0.5,1.0,2.0{0.0,0.5,1.0,2.0} for both models. For each $\alpha$, we measure semantic scores under steering with $v$ versus $v+v_{\perp}$ and plot response curves. If equivalence is scale-invariant, the curves should track closely across all $\alpha$ with overlapping confidence bands.
6.2 Orthogonal Perturbation Test Results
Table 1 presents our empirical findings across two models, three semantic traits and two sample sizes ($n=5$ and $n=10$ random orthogonal seeds). Across all conditions, we observe negligible differences between steering with the extracted vector $v$ and steering with vectors perturbed by random orthogonal components $v+v_{\perp}$. We measure steering efficacy using semantic probe scores in $[0,1]$ that quantify the intensity of the target trait in generated outputs, comparing score distributions from steered generations against baseline generations.
Table 1: Empirical validation of non-identifiability across models, traits and sample sizes. Cohen’s $d$ measures effect size between $v$ and $v+v_{\perp}$ (lower = more equivalent). All values show negligible differences ($d<0.2$), confirming observational equivalence.
Model Trait Seeds Cohen’s $d$ Correlation Perp-Only
Qwen2.5-3B-Instruct Formality $n=5$ $0.144\pm 0.095$ $0.210\pm 0.113$ $98.8\%$
$n=10$ $0.075\pm 0.058$ $0.285\pm 0.087$ $100.4\%$
Politeness $n=5$ $0.112\pm 0.077$ $0.319\pm 0.070$ $101.5\%$
$n=10$ $0.092\pm 0.069$ $0.414\pm 0.083$ $100.6\%$
Humor $n=5$ $0.103\pm 0.077$ $0.276\pm 0.177$ $99.7\%$
$n=10$ $072\pm 0.061 0.0$$044\pm 0.097$ $100.5\%$
Llama-3.1-8B-Instruct Formality $n=5$ $0.052\pm 0.029$ $0.164\pm 0.044$ $95.6\%$
$n=10$ $0.096\pm 0.068$ $0.192\pm 0.085$ $96.8\%$
Politeness $n=5$ $0.074\pm 0.041$ $0.324\pm 0.058$ $101.2\%$
$n=10$ $0.085\pm 0.043$ $0.347\pm 0.077$ $100.4\%$
Humor $n=5$ $159\pm 0.109 0.0$$032\pm 0.116$ $98.4\%$
$n=10$ $119\pm 0.119 0.0$$016\pm 0.104$ $95.9\%$
With $n=10$ seeds, the mean Cohen’s $d$ is 0.080 for Qwen2.5-3B and 0.100 for Llama-3.1-8B, both well below the threshold for small effects ($d=0.2$) and firmly in the negligible range. The “Perp-Only Effect” column measures the efficacy of steering with pure orthogonal components ($v_{\perp}$ alone, without $v$) relative to the extracted vector—values near $100\%$ indicate that orthogonal components achieve equivalent behavioral impact.
Statistical stability across sample sizes. The convergence between $n=5$ and $n=10$ results demonstrates robustness across both models. For Qwen2.5-3B, effect sizes change by <0.07<0.07 as sample size increases, with all traits showing tightening confidence intervals: formality ($0.144\to 0.075$), politeness ($0.112\to 0.092$), humor ($0.103\to 0.072$). For Llama-3.1-8B, we observe similar stability: formality ($0.052\to 0.096$), politeness ($0.074\to 0.085$), humor ($0.159\to 0.119$). This consistency demonstrates that the observed equivalence is not a sampling artifact but a stable property of the steering geometry.
Cross-model consistency. The close agreement between Qwen2.5-3B ($d=0.080$) and Llama-3.1-8B ($d=0.100$) demonstrates that observational equivalence is not model-specific. Despite differences in architecture, scale (3B vs. 8B parameters) and hidden dimension ($d=2048$ vs. $d=4096$), both models exhibit nearly identical patterns of non-identifiability. The consistency across traits within each model further confirms that the phenomenon is general rather than an artifact of specific semantic domains or model implementations.
Visualizing orthogonal component equivalence. Figure 1 illustrates the perp-only effect ratios across all traits and models using $n=10$ orthogonal seeds. The tight clustering around the perfect equivalence line (dashed red at $1.0$) demonstrates that pure orthogonal components achieve nearly identical steering efficacy to the extracted vectors.
(a) Qwen2.5-3B-Instruct
(b) Llama-3.1-8B-Instruct
Figure 1: Perp-only effect ratios ($v_{\perp}$ efficacy / $v$ efficacy) for $n=10$ orthogonal seeds. Values near 1.0 (dashed red line) indicate perfect equivalence.
[…]
6.3 Scale Invariance Analysis
To verify that observational equivalence holds across different steering strengths, we evaluate the formality trait at four steering magnitudes $\alpha\in\{0.0,\,0.5,\,1.0,\,2.0\}$ for both models. Figure 2 shows the response curves for the extracted vector $v$ and the observationally equivalent vector $v+v_{\perp}$.
(a) Qwen2.5-3B-Instruct
(b) Llama-3.1-8B-Instruct
Figure 2: Scale invariance of observational equivalence. Formality scores across steering strengths $\alpha\in\{0.0,\,0.5,\,1.0,\,2.0\}$ for the extracted vector $v$ (blue circles) and the perturbed vector $v+v_{\perp}$ (orange squares).
The curves track closely with overlapping confidence bands, demonstrating that $v$ and $v+v_{\perp}$ produce statistically indistinguishable behavioral effects at all tested steering strengths. For Qwen2.5-3B, mean differences remain below $0.022$ across all $\alpha$ values (less than $2.5\%$ deviation in formality scores); for Llama-3.1-8B, differences remain below $0.023$ ($2.5\%$ deviation). This confirms that non-identifiability is a structural property: any vector in the equivalence class $v+\ker(J)$ produces identical observations regardless of scaling.
7 Conclusion
Rather than viewing non-identifiability as a limitation, we frame it as a necessary foundation for principled steering research. Our characterization clarifies what behavioral observations can reveal about representations and provides a framework for designing interventions that are both empirically effective and theoretically grounded. The empirical validation demonstrates that these are not abstract concerns: non-identifiability affects real steering interventions in deployed models, with typical steering vectors containing substantial observationally irrelevant components. However, the constructive characterization in Proposition 2 shows that identifiability is achievable through principled design choices, enabling alignment methods that clearly specify their affordances and limitations.
8 Limitations
Our empirical validation covers two models at mid-network layers across three semantic traits (formality, politeness, humor). While our theoretical results apply generally, the empirical magnitude of non-identifiability, including null space dimensionality and the fraction of vector norm therein, may vary across model families, scales, architectures and layer positions. Our semantic evaluation uses lexical heuristics rather than full distributional metrics or human judgments, providing a conservative test of observational equivalence. The primary theoretical mechanism relies on local linear approximation, formally valid in a neighborhood of the reference distribution, though our empirical validation confirms large equivalence classes exist in practice.
Proposition 2 characterizes sufficient conditions for identifiability but comprehensive empirical evaluation of each regime across multiple models, tasks and experimental designs remains future work. The practical applicability of identifiable regimes involves trade-offs: ICA requires carefully curated contrastive prompts, sparsity regularization may reduce efficacy if true factors are not sparse and multi-environment validation requires substantial additional data. Understanding these trade-offs and extending our framework to other intervention modalities (activation patching, prompt based steering) are important directions for the representational alignment community.
9 Future Work
Systematic evaluation of identifiable regimes, comparing ICA-based extraction, sparse recovery, multi-environment training and cross-layer validation against standard contrastive methods, would clarify which structural assumptions are most effective in practice. Hybrid methods combining multiple identifiability conditions may achieve stronger identifiability than individual approaches, while layer-wise analysis could reveal whether certain network positions afford more identifiable steering and decomposing the Jacobian by architectural component could identify which elements contribute to non-identifiability.
The multi-environment identifiability condition connects to causal representation learning, where adapting invariant risk minimization to train steering vectors that maintain consistent effects across distributions could filter spurious correlations. Investigating adversarial robustness and out-of-distribution generalization would test whether identifiable methods exhibit superior robustness. Characterizing scaling laws for identifiability, including how null-space dimensionality evolves as models scale, would inform long-term strategy for controlling future systems.
Acknowledgments
We thank Vast.ai for providing GPU compute resources that enabled our empirical validation experiments. We also acknowledge the HuggingFace platform and maintainers of the open-source models (Qwen2.5-3B-Instruct, Llama-3.1-8B-Instruct) used in this work.
Appendix A Prompt Construction Details
A.1 Contrastive Prompt Pairs
We construct contrastive prompt pairs for each semantic trait using template-based generation. Figure 3 illustrates the general structure and Table 2 provides concrete examples.
Topic Pool 50 topics e.g., “weather”, “tech trends” Positive Template “formal”, “polite” Negative Template “informal”, “rude” 50 Contrastive Prompt Pairs
Figure 3: Contrastive prompt pair construction.
Trait Positive Prompt ($x^{+}$) Negative Prompt ($x^{-}$)
Formality Write a professional and formal message about technology trends. Write a casual and informal message about technology trends.
Politeness Write a polite and courteous response for a disagreement with someone. Write a rude and disrespectful response for a disagreement with someone.
Humor Write a humorous and funny response about an awkward moment. Write a serious and straightforward response about an awkward moment.
Table 2: Example contrastive prompt pairs for each semantic trait. Each pair consists of a positive prompt ($x^{+}$) designed to elicit high trait values and a negative prompt ($x^{-}$) designed to elicit low trait values.
A.2 Steering Extraction and Evaluation Pipeline
Figure 4 illustrates the pipeline used to extract and evaluate steering vectors. Contrastive prompt pairs are first used to compute a steering vector specific to a semantic trait. This vector is then applied to a separate set of held-out evaluation prompts to generate steered outputs.
Steering Extraction 50 contrastive pairs (trait-specific templates) Steering Vector $v\in\mathbb{R}^{d}$ Evaluation 100 held-out prompts (diverse, neutral) Steered Output $o(x,v,\alpha)$ extractapplygenerate Key: Test prompts are intentionally orthogonal to extraction prompts to ensure steering generalizes beyond training distribution.
Figure 4: Steering extraction and evaluation pipeline. Steering vectors are extracted from trait-specific contrastive pairs, then evaluated on diverse held-out prompts to test generalization.
Appendix B Detailed Proof of Proposition 1
B.1 Statement and Overview
Proposition 1.
Under Assumptions A1–A3, in Regime 2 (white-box single-layer access) without additional structural constraints, persona vectors are not identifiable. Specifically, for any steering vector $v\in\mathbb{R}^{d}$, there exist infinitely many vectors $v^{\prime}\not\propto v$ that are observationally equivalent.
Proof strategy.
We establish non-identifiability through two complementary mechanisms:
  • Null-space ambiguity (primary, constructive).
  • Reparameterization symmetry (existence-based).
B.2 Null-Space Ambiguity (Constructive Proof)
Setup. Consider the local linear approximation of the steering effect:
$o(x,v,\alpha)\approx o(x,0,0)+\alpha J_{\ell}(x)v$ (2)
where $J_{\ell}(x)=\frac{\partial o}{\partial h_{\ell}}\big|_{h_{\ell}(x)}\in\mathbb{R}^{V\times d}$ is the Jacobian.
Step 1: Null space characterization.
Define the null space of $J_{\ell}$ as:
$\mathcal{N}=\ker(J_{\ell})=\{v_{0}\in\mathbb{R}^{d}:J_{\ell}v_{0}=0\}$ (3)
By the rank-nullity theorem:
$\dim(\mathcal{N})=d-\operatorname{rank}(J_{\ell})$ (4)
Step 2: Rank Bound.
The Jacobian $J_{\ell}\in\mathbb{R}^{V\times d}$ has maximum possible rank:
$\text{rank}(J_{\ell})\leq\min(V,d)$ (5)
In modern language models:
  • Hidden dimension: $d\approx 4000$ (typical)
  • Vocabulary size: $V\approx 50000$ (typical)
  • Therefore: $\max\text{rank}(J_{\ell})=d$
Step 3: Effective rank is much lower.
Output distributions lie on a low-dimensional manifold. The effective rank satisfies:
$\text{rank}_{\epsilon}(J_{\ell})=\#\{\sigma_{i}:\sigma_{i}>\epsilon\cdot\sigma_{\max}\}\ll d$ (6)
where $\sigma_{i}$ are singular values of $J_{\ell}$ and $\epsilon$ is a threshold (e.g., $10^{-4}$).
Intuition: In overparameterized LLMs, the effective rank is expected to be strictly less than $d$ due to the low-dimensional structure of output distributions. Therefore $\dim(\mathcal{N})=d-\text{rank}(J_{\ell})$ is generically positive but this is not required for the proof—the argument establishes non-identifiability whenever $\dim(\mathcal{N})\geq 1$.
Step 4: Constructing equivalent vectors.
For any steering vector $v\in\mathbb{R}^{d}$ and any $v_{0}\in\mathcal{N}$, define
$v^{\prime}=v+v_{0}.$ (7)
Then for all $x$ and all $\alpha$:
$J_{\ell}(x)v^{\prime}$ $=J_{\ell}(x)(v+v_{0})$ (8)
$=J_{\ell}(x)v+J_{\ell}(x)v_{0}$
$=J_{\ell}(x)v.$
Therefore:
$o(x,v^{\prime},\alpha)$ $\approx o(x,0,0)+\alpha J_{\ell}(x)v^{\prime}$ (9)
$=o(x,0,0)+\alpha J_{\ell}(x)v$
$\approx o(x,v,\alpha).$
Step 5: Infinitely many distinct solutions.
Since $\dim(\mathcal{N})\geq 1$, the null space contains infinitely many directions. For any $v_{0}\in\mathcal{N}\setminus\{0\}$ and any $\beta\in\mathbb{R}$:
$v^{\prime}_{\beta}=v+\beta v_{0}$ (10)
generates infinitely many observationally equivalent vectors. Furthermore, if $\beta$ is chosen such that $v^{\prime}_{\beta}\not\propto v$ (which is always possible unless $v\propto v_{0}$), these vectors are geometrically distinct.
Step 6: Non-proportionality.
To ensure $v^{\prime}_{\beta}\not\propto v$, we need $v+\beta v_{0}\neq cv$ for any scalar $c$. This fails only if:
$\beta v_{0}=(c-1)v$ (11)
which requires $v\in\text{span}(v_{0})$. Since $v_{0}$ is an arbitrary element of a $\dim(\mathcal{N})$-dimensional space and $v$ is fixed, this occurs with probability zero. Therefore, for generic $v$ and generic $v_{0}\in\mathcal{N}$, we have $v^{\prime}_{\beta}\not\propto v$ for almost all $\beta$.
Conclusion (Null-space mechanism).
Under the linear approximation and for typical Jacobian structure, there exist infinitely many geometrically distinct steering vectors that are observationally equivalent.
B.3 Reparameterization Symmetry (Existence Proof)
Setup: Neural networks exhibit inherent symmetries arising from overparameterization. We show these symmetries induce non-identifiability even in the exact nonlinear case.
Step 1: Representation reparameterization.
Consider an invertible transformation $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$. Define the reparameterized representation:
$h^{\prime}_{\ell}(x)=T\bigl(h_{\ell}(x)\bigr).$ (12)
If the subsequent layers can be rewritten as:
$F_{\ell\to L}(h_{\ell})=F^{\prime}_{\ell\to L}(T(h_{\ell}))$ (13)
then $(h_{\ell},v)$ and $(h^{\prime}_{\ell},v^{\prime})$ where $v^{\prime}=DT(h_{\ell})\cdot v$ are observationally indistinguishable.
Step 2: Linear reparameterizations.
Consider linear transformations $T(h)=Ah$ where $A\in\mathbb{R}^{d\times d}$ is invertible. For layer $\ell+1$ with weight matrix $W_{\ell+1}\in\mathbb{R}^{d\times d}$ and bias $b_{\ell+1}$:
Original computation:
$h_{\ell+1}=\sigma(W_{\ell+1}h_{\ell}+b_{\ell+1})$ (14)
Reparameterized computation:
$h^{\prime}_{\ell+1}=\sigma(W^{\prime}_{\ell+1}h^{\prime}_{\ell}+b^{\prime}_{\ell+1})$ (15)
where $W^{\prime}_{\ell+1}=W_{\ell+1}A^{-1}$ and $b^{\prime}_{\ell+1}=b_{\ell+1}$.
Note: The bias remains invariant under linear reparameterization through the origin. For more general affine transformations $T(h)=Ah+c$ with translation $c\neq 0$, the bias would also transform as $b^{\prime}_{\ell+1}=b_{\ell+1}-W_{\ell+1}A^{-1}c$. We restrict to origin-preserving transformations for simplicity, as these suffice to establish non-identifiability.
Then:
$h^{\prime}_{\ell+1}$ $=\sigma\!\left(W_{\ell+1}A^{-1}Ah_{\ell}+b_{\ell+1}\right)$ (16)
$=\sigma\!\left(W_{\ell+1}h_{\ell}+b_{\ell+1}\right)$
$=h_{\ell+1}.$
Step 3: Steering under reparameterization.
Original steering:
$\tilde{h}_{\ell}=h_{\ell}+\alpha v$ (17)
$\tilde{h}_{\ell+1}=\sigma(W_{\ell+1}(h_{\ell}+\alpha v)+b_{\ell+1})$ (18)
Reparameterized steering with $v^{\prime}=Av$:
$\tilde{h}^{\prime}_{\ell}$ $=h^{\prime}_{\ell}+\alpha v^{\prime}$ (19)
$=Ah_{\ell}+\alpha Av$
$=A(h_{\ell}+\alpha v)$
$=A\tilde{h}_{\ell}$,
$\tilde{h}^{\prime}_{\ell+1}$ $=\sigma\!\left(W^{\prime}_{\ell+1}(h^{\prime}_{\ell}+\alpha v^{\prime})+b^{\prime}_{\ell+1}\right)$
$=\sigma\!\left(W_{\ell+1}A^{-1}A(h_{\ell}+\alpha v)+b_{\ell+1}\right)$
$=\sigma\!\left(W_{\ell+1}(h_{\ell}+\alpha v)+b_{\ell+1}\right)$
$=\tilde{h}_{\ell+1}.$
Step 4: Infinitely many reparameterizations.
For any invertible matrix $A$ with $A\neq cI$ (i.e., not a scalar multiple of identity), we obtain $v^{\prime}=Av\not\propto v$. The space of such matrices has dimension $d^{2}-1$ (excluding scalar multiples), providing infinitely many distinct reparameterizations.
Important note: While we cannot explicitly construct these reparameterizations for a frozen model without retraining (since this would require modifying $W_{\ell+1}$), their existence follows from the fundamental symmetry structure of neural networks. This establishes that identifiability cannot hold even in principle without additional constraints.
Practical implication: For frozen, deployed models, only the null-space mechanism is operationally relevant–the reparameterization symmetry is not a realizable operation. However, the reparameterization argument shows that even if we allowed weight modifications during training, identifiability would still fail without structural constraints. Thus, non-identifiability is both a practical limitation (null-space) and a fundamental theoretical barrier (gauge symmetry).
Conclusion (Reparameterization mechanism).
The inherent symmetries of neural network parameterizations induce infinitely many observationally equivalent steering vectors.
Appendix C Detailed Proof of Proposition 2
C.1 Statement and Overview
Proposition 2.
Persona vectors can be identified up to scaling and permutation, thus affording reliable alignment control, under the following sufficient structural conditions:
  • Statistical Independence (ICA). Latent components are statistically independent, allowing unique recovery via independent component analysis.
  • Sparsity constraints. Steering directions admit sparse representations that reduce null-space ambiguity.
  • Multi-environment or interventional data. Variation across environments or interventions breaks observational equivalences.
  • Cross-layer consistency. Valid semantic directions propagate coherently across layers, filtering spurious components.
We provide detailed proofs for each condition.
C.2 Proof of Condition: Statistical Independence (ICA)
Setup: Assume the latent persona is a vector $z=(z_{1},\ldots,z_{k})\in\mathbb{R}^{k}$ with independent components and the representation follows:
$h_{\ell}=Az+\epsilon$ (20)
where $A\in\mathbb{R}^{d\times k}$ is a mixing matrix and $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ is Gaussian noise.
Goal: Show that $A$ (and hence its columns, which are the steering vectors) can be recovered up to permutation and scaling.
Step 1: ICA identifiability theorem.
Let $x=As$ where $s\in\mathbb{R}^{k}$ has independent components and $A\in\mathbb{R}^{d\times k}$ is full column rank. Under the following sufficient conditions (Comon, 1994; Hyvärinen and Oja, 2000):
  • At most one component of $s$ is Gaussian
  • Components are statistically independent
  • Sufficient samples are observed
Then $A$ is identifiable up to column permutation and column scaling.
Important caveat: These are strong assumptions that may not hold exactly in practice for LLM personas. Statistical independence between semantic factors (e.g., formality and politeness) is an idealization; real personas may exhibit weak dependencies. The ICA framework provides a sufficient structural condition for identifiability, not a characterization of what typically holds in contemporary steering pipelines.
Step 2: Application to steering.
In our setting:
  • Observations: $h_{\ell}(x_{1}),\ldots,h_{\ell}(x_{N})$
  • Model: $h_{\ell}(x_{i})=Az(x_{i})+\epsilon_{i}$
  • Sources: $z(x_{1}),\ldots,z(x_{N})$ are realizations of independent persona components
Assumption verification:
  • Independence: We assume $z_{1},\ldots,z_{k}$ are statistically independent
  • Non-Gaussianity: At most one $z_{i}$ is Gaussian (e.g., politeness and formality are typically non-Gaussian in natural language)
  • Full rank: $A$ has full column rank (each persona has a non-zero effect on representations)
Step 3: Recovery via ICA algorithm.
Apply ICA algorithm (Hyvärinen and Oja, 2000) to observations $\{h_{\ell}(x_{i})\}_{i=1}^{N}$:
  • Whitening: Compute $\tilde{h}_{\ell}=\Sigma^{-1/2}(h_{\ell}-\mu)$ where $\mu=\mathbb{E}[h_{\ell}]$ and $\Sigma=\operatorname{Cov}(h_{\ell})$.
  • Independent component extraction: Find unmixing matrix $W$ that maximizes non-Gaussianity of $\hat{z}=W\tilde{h}_{\ell}$.
  • Recover mixing matrix: $\hat{A}=\Sigma^{1/2}W^{-1}$.
Step 4: Identifiability guarantee.
By the ICA theorem, $\hat{A}=APD$ where:
  • $P\in\mathbb{R}^{k\times k}$ is a permutation matrix
  • $D\in\mathbb{R}^{k\times k}$ is a diagonal scaling matrix
Therefore, the columns of $\hat{A}$ are the columns of $A$ up to permutation and scaling. Since steering vectors are defined as $v_{i}=Ae_{i}$ (the $i$-th column of $A$), we recover the true steering directions uniquely (up to unavoidable symmetries).
Step 5: Noise robustness.
With Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$, ICA remains consistent:
$h_{\ell}=Az+\epsilon$ (21)
Since $\epsilon$ is Gaussian and $z$ is non-Gaussian (by assumption), the ICA algorithm separates signal from noise:
  • Recovered sources: $\hat{z}=z+\tilde{\epsilon}$ where $\tilde{\epsilon}$ is small for $\sigma^{2}\ll\|Az\|^{2}$
  • Recovered mixing: $\hat{A}\approx APD$ with estimation error decreasing as $N\to\infty$
Conclusion (ICA): Under statistical independence, non-Gaussianity and full column rank—strong sufficient conditions that may not hold automatically in practice—persona vectors are identifiable up to permutation and scaling. These assumptions provide a theoretical pathway to identifiability but require careful experimental design to approximate in real steering settings.
C.3 Proof of Condition: Sparsity Constraints
Setup: Assume the true persona vector $v\in\mathbb{R}^{d}$ is sparse with $\|v\|_{0}=s\ll d$, meaning at most $s$ entries are non-zero. We observe the effect of steering:
$y=J_{\ell}v+\eta$ (22)
where $y\in\mathbb{R}^{m}$ are output measurements, $J_{\ell}\in\mathbb{R}^{m\times d}$ is the measurement matrix (Jacobian) and $\eta\sim\mathcal{N}(0,\sigma^{2}I)$ is noise.
Goal: Show that $v$ can be uniquely recovered via $\ell_{1}$ minimization.
Step 1: Compressed sensing setup.
The recovery problem is:
$\min_{v^{\prime}\in\mathbb{R}^{d}}\|v^{\prime}\|_{1}\quad\text{subject to}\quad\|J_{\ell}v^{\prime}-y\|_{2}\leq\epsilon$ (23)
where $\epsilon$ bounds the noise: $\epsilon\geq\|\eta\|_{2}$ with high probability.
Step 2: Restricted Isometry Property (RIP).
A matrix $J_{\ell}$ satisfies the RIP of order $s$ with constant $\delta_{s}$ if for all $s$-sparse vectors $v$:
$(1-\delta_{s})\|v\|_{2}^{2}\leq\|J_{\ell}v\|_{2}^{2}\leq(1+\delta_{s})\|v\|_{2}^{2}$ (24)
Theorem (Candes and Tao, 2005): If $J_{\ell}$ satisfies RIP with $\delta_{2s}<\sqrt{2}-1\approx 0.414$, then $\ell_{1}$ minimization recovers $v$ exactly (in the noiseless case) or approximately (with noise) with error:
$\|v-\hat{v}\|_{2}\leq C\epsilon$ (25)
for some constant $C$ depending on $\delta_{2s}$.
Important caveat: The RIP condition is a strong assumption. For Jacobians $J_{\ell}$ arising from neural network steering, RIP typically does not hold automatically and must be verified empirically or enforced through measurement design (diverse prompt selection). This condition is sufficient for identifiability but may not be satisfied in standard steering settings without careful experimental design.
Step 3: Recovery guarantee.
Solve:
$\hat{v}=\arg\min_{v^{\prime}}\|v^{\prime}\|_{1}\quad\text{s.t.}\quad\|J_{\ell}v^{\prime}-y\|_{2}\leq\epsilon$ (26)
By the compressed sensing theorem:
$\|v-\hat{v}\|_{2}\leq C_{1}\epsilon+C_{2}\frac{\|v-v_{s}\|_{1}}{\sqrt{s}}$ (27)
where $v_{s}$ is the best $s$-sparse approximation to $v$. If $v$ is exactly $s$-sparse, the second term vanishes and:
$\|v-\hat{v}\|_{2}\leq C_{1}\epsilon$ (28)
Step 4: Uniqueness.
Suppose two sparse vectors $v$ and $v^{\prime}$ both satisfy $\|J_{\ell}v-y\|_{2}\leq\epsilon$. Then:
$\|J_{\ell}(v-v^{\prime})\|_{2}\leq 2\epsilon$ (29)
If $v-v^{\prime}$ is $2s$-sparse and $J_{\ell}$ satisfies RIP, then by the RIP condition:
$(1-\delta_{2s})\|v-v^{\prime}\|_{2}^{2}\leq\|J_{\ell}(v-v^{\prime})\|_{2}^{2}\leq 4\epsilon^{2}$ (30)
Therefore:
$\|v-v^{\prime}\|_{2}\leq\frac{2\epsilon}{\sqrt{1-\delta_{2s}}}$ (31)
As $\epsilon\to 0$ (noiseless case), $v=v^{\prime}$, establishing uniqueness.
Conclusion (Sparsity): Under sparsity assumptions and RIP conditions—strong sufficient conditions that typically require careful measurement design—persona vectors are uniquely recoverable via $\ell_{1}$ minimization. RIP does not hold automatically for arbitrary Jacobians and must be verified or engineered through diverse prompt selection.
C.4 Proof of Condition: Multi-Environment Identification
Setup: Assume we observe the same persona $z$ across multiple environments $e\in\{1,\ldots,E\}$ where:
$h_{\ell}^{(e)}=g_{e}(x,z)+\epsilon^{(e)}$ (32)
The function $g_{e}$ may vary across environments (spurious correlations change) but the causal effect of $z$ on downstream outputs remains invariant.
Goal: Show that the invariant representation $z$ (and its associated direction) can be identified.
Step 1: Invariant causal mechanism.
Assume the causal graph:
[…]
where:
  • $z$: true persona (causal factor),
  • $x$: input prompt,
  • $e$: environment,
  • $h_{\ell}$: internal representation,
  • $o$: output.
The key assumption is invariance: the causal mechanism $h_{\ell}\to o$ is the same across environments, i.e., $F_{\ell\to L}$ does not depend on $e$.
Step 2: Invariant Risk Minimization (IRM).
The objective is to find representation $\Phi:\mathbb{R}^{d}\to\mathbb{R}^{k}$ and predictor $w:\mathbb{R}^{k}\to\mathbb{R}$ such that:
$\min_{\Phi,w}$ $\sum_{e=1}^{E}R^{e}(\Phi,w)$ (33)
s.t. $w\in\arg\min_{w^{\prime}}R^{e}(\Phi,w^{\prime})\quad\forall e$
where,
$R^{e}(\Phi,w)=\mathbb{E}_{(x,y)\sim\mathcal{P}^{e}}\left[\ell\big(w\cdot\Phi(h_{\ell}(x)),y\big)\right].$ (34)
Interpretation: Find a representation $\Phi(h_{\ell})=z$ such that the optimal predictor $w$ is the same across all environments. This filters out spurious correlations that vary with $e$.
Step 3: Identifiability under IRM.
Under the following idealized conditions:
  • Environments are sufficiently diverse: $\text{Cov}(x|e)$ varies across $e$
  • Causal mechanism is invariant: $p(o|z)$ is the same for all $e$
  • Sufficient environments: $E\geq k+1$ where $k=\dim(z)$
  • No unobserved confounders between environments and outcomes
The theorem (Ahuja et al., 2022; Von Kügelgen et al., 2021) states that the invariant risk minimization objective can recover $z$ up to an invertible transformation, though not necessarily uniquely.
Important caveat: The literature on invariant representation learning establishes conditions under which invariance constraints narrow the hypothesis class but full identifiability (unique recovery of $z$) requires additional assumptions beyond standard IRM formulations. The practical application of IRM to steering should be understood as identifying a stable, transferable representation rather than guaranteeing uniqueness. This is a sufficient condition in principle under strong assumptions, not a characterization of typical steering scenarios.
Step 4: Application to steering.
In practice:
  1. 1.
    Collect multi-environment data: Extract steering vectors from diverse prompt distributions, model checkpoints or instruction formats.
  2. 2.
    Learn invariant representation: Train $\Phi$ via IRM objective
  3. 3.
    Extract steering vectors: Compute $v_{i}=\nabla_{h_{\ell}}\Phi_{i}(h_{\ell})$ where $\Phi_{i}$ is the $i$-th component of $\Phi$
The resulting steering vectors $v_{i}$ capture invariant causal factors rather than spurious correlations.
Step 5: Theoretical guarantee.
Under mild conditions (sufficient diversity, invariance, identifiable causal structure), the IRM solution satisfies:
$\Phi(h_{\ell})=T(z)$ (35)
where $T$ is an invertible transformation. Since persona steering operates on $z$, the directions $v_{i}=\nabla_{h_{\ell}}\Phi_{i}$ recover the true causal factors.
Conclusion (Multi-environment): With diverse environments and invariant causal mechanisms—idealized sufficient conditions that narrow the hypothesis class—persona vectors corresponding to stable causal factors can be recovered up to invertible transformations. Full uniqueness requires additional assumptions beyond standard IRM formulations. This provides a principled approach to filtering spurious correlations but should be understood as identifying transferable representations rather than guaranteeing unique recovery.
C.5 Proof of Condition: Cross-Layer Consistency
Setup: Assume persona vectors exhibit consistent geometric relationships across layers:
$v_{\ell+1}=W_{\ell}v_{\ell}+\delta_{\ell}$ (36)
where $W_{\ell}\in\mathbb{R}^{d\times d}$ is the weight matrix connecting layers and $\|\delta_{\ell}\|$ is small.
Goal: Show that cross-layer constraints reduce the solution space and improve identifiability.
Step 1: Single-layer null space.
From proposition 1, at layer $\ell$:
$\mathcal{N}_{\ell}=\{v_{0}:J_{\ell}v_{0}=0\}$ (37)
with $\dim(\mathcal{N}_{\ell})=d-r_{\ell}$ where $r_{\ell}=\text{rank}(J_{\ell})$.
Step 2: Cross-layer propagation.
If $v_{\ell}\in\mathcal{N}_{\ell}$, does $v_{\ell+1}=W_{\ell}v_{\ell}\in\mathcal{N}_{\ell+1}$?
Generally, no. The null space changes across layers:
$v_{\ell}\in\mathcal{N}_{\ell}\not\Rightarrow W_{\ell}v_{\ell}\in\mathcal{N}_{\ell+1}$ (38)
Step 3: Intersection of constraints.
Consider steering vectors observed at multiple layers $\ell_{1},\ldots,\ell_{L}$. Each layer provides a constraint:
$J_{\ell_{i}}v_{\ell_{i}}=y_{i}$ (39)
If we additionally require consistency:
$v_{\ell_{i+1}}=W_{\ell_{i}}v_{\ell_{i}}+\delta_{\ell_{i}}$ (40)
Then the solution must satisfy:
$v_{\ell_{i}}\in\{v:J_{\ell_{i}}v=y_{i}\}\cap\{v:W_{\ell_{i}}v\approx v_{\ell_{i+1}}\}$ (41)
Step 4: Reduced null space (qualitative characterization).
The cross-layer constraints create an overdetermined system. The effective null space is:
$\mathcal{N}_{\text{eff}}=\bigcap_{i=1}^{L-1}\{v:W_{\ell_{i}}v\in\mathcal{N}_{\ell_{i+1}}\}$ (42)
Since each $W_{\ell_{i}}$ generically has full rank and maps null space vectors to non-null-space vectors, we expect:
$\dim(\mathcal{N}_{\text{eff}})\ll\dim(\mathcal{N}_{\ell})$
Intuition: Each additional layer imposes new independent constraints. If the propagation matrices $W_{\ell_{i}}$ are sufficiently "generic" (full rank with uncorrelated null space mappings), then each layer reduces the effective null-space dimension. For sufficiently many informative layers with uncorrelated constraint structures, the effective null space can shrink dramatically or even vanish.
Conclusion (Cross-layer): Cross-layer consistency constraints can substantially reduce null-space dimensionality by creating overdetermined systems. The extent of reduction depends on the specific geometric structure of propagation matrices and layer-wise Jacobians. While this approach does not guarantee complete identifiability in general, it provides a practical method for filtering spurious null-space components that lack geometric stability across layers.
Appendix D Intuitive Explanations
D.1 Geometric Intuition for Null-Space Ambiguity
Visual analogy:
Consider a simplified example where a 3D steering vector $v\in\mathbb{R}^{3}$ affects 2D outputs $o\in\mathbb{R}^{2}$ through a projection matrix $J\in\mathbb{R}^{2\times 3}$. By the rank-nullity theorem, the null space $\ker(J)$ is a 1D subspace since $\dim(\ker(J))=3-\text{rank}(J)=3-2=1$.
For concreteness, suppose $J=[I_{2}\mid 0]$ projects onto the first two coordinates. Then $\ker(J)=\text{span}\{(0,0,1)\}$ is the $v_{3}$ axis. The key insight generalizes: for any projection matrix $J$, directions in $\ker(J)$ are invisible to outputs. Figure 5 illustrates this geometric intuition.
[…]
Figure 5: Geometric intuition for null-space ambiguity. The output observes only the $(v_{1},v_{2})$ components (blue shaded region). The $v_{3}$ component lies in $\ker(J)$ and is invisible to the output. Adding any $\alpha v_{3}$ to $v$ leaves the output unchanged: $J(v+\alpha v_{3})=Jv$ for all $\alpha\in\mathbb{R}$.
Any vector of the form $v^{\prime}=v+\alpha v_{3}$ for $\alpha\in\mathbb{R}$ produces identical outputs:
$J(v+\alpha v_{3})$ $=Jv+\alpha J(0,0,1)^{\top}$ (43)
$=Jv+\alpha\cdot 0$
$=Jv$
Since $\alpha$ can take infinitely many values and $v^{\prime}\not\propto v$ for generic choices of $\alpha$, there exist infinitely many geometrically distinct steering vectors that are observationally equivalent.
D.2 Why More Data Doesn’t Help
Common misconception: "If we collect more steering examples, we can pin down the unique vector."
Why this fails:
Consider observing steering effects on $N$ prompts $\{x_{1},\ldots,x_{N}\}$. Each observation provides:
$o_{i}=J_{\ell}(x_{i})v+\eta_{i}$ (44)
Stacking these:
$\begin{bmatrix}o_{1}\\ o_{2}\\ \vdots\\ o_{N}\end{bmatrix}=\begin{bmatrix}J_{\ell}(x_{1})\\ J_{\ell}(x_{2})\\ \vdots\\ J_{\ell}(x_{N})\end{bmatrix}v+\begin{bmatrix}\eta_{1}\\ \eta_{2}\\ \vdots\\ \eta_{N}\end{bmatrix}$ (45)
The stacked Jacobian $J_{\text{stack}}\in\mathbb{R}^{(N\cdot V)\times d}$ has null space:
$\ker(J_{\text{stack}})=\bigcap_{i=1}^{N}\ker(J_{\ell}(x_{i}))$ (46)
Critical observation: If all $J_{\ell}(x_{i})$ share a common null space (e.g., all prompts probe similar aspects of the model), then:
$\ker(J_{\text{stack}})=\ker(J_{\ell}(x_{1}))\neq\{0\}$ (47)
This means adding more prompts does not reduce the null space when all prompts probe the same effective subspace—the ambiguity persists. Figure 6 illustrates this phenomenon.
[…]
Figure 6: Illustration of why prompt diversity does not resolve null-space ambiguity. Each prompt induces a Jacobian that measures the same effective subspace, leaving the null-space intersection unchanged.
D.3 Why Orthogonal Perturbations Preserve Semantics
Setup: We test steering with $v$ versus $v+v_{\perp}$, where $v_{\perp}\perp v$ and $\|v_{\perp}\|=1$. Empirically, adding a random orthogonal direction produces nearly equivalent semantic effects. Why?
Decomposing the perturbation.
Any perturbation $v_{\perp}$ can be decomposed into observable and invisible components:
$v_{\perp}=v_{\perp,\text{row}}+v_{\perp,\text{null}}$ (48)
where $v_{\perp,\text{row}}\in\operatorname{row}(J)$ (observable through outputs) and $v_{\perp,\text{null}}\in\ker(J)$ (invisible to outputs).
For a random $v_{\perp}\perp v$, the expected fraction in the null space is:
$\mathbb{E}[\|v_{\perp,\text{null}}\|^{2}]\approx\frac{\dim(\ker(J))}{d}\approx 0.20{-}0.25$ (49)
This means $\sim$20–25% of the perturbation is automatically invisible to model outputs.
Observable effect is diffuse and unstructured.
The actual output change from adding $v_{\perp}$ is:
$J(v+v_{\perp})=Jv+Jv_{\perp,\text{row}}$ (50)
While $Jv_{\perp,\text{row}}$ is not necessarily small in norm, it represents a random direction in the $\sim$3200-dimensional row space of $J$. In contrast, $Jv$ is a structured, semantically coherent steering effect (e.g., consistently shifting outputs toward “honesty” or “sycophancy”).
The key distinction is not magnitude but semantic coherence: $Jv$ produces aligned, directional changes across the vocabulary, while $Jv_{\perp,\text{row}}$ produces diffuse, incoherent perturbations that average out across tokens and do not systematically shift semantic meaning in any consistent direction.
Analogy: If $Jv$ is a strong wind blowing consistently north, then $Jv_{\perp,\text{row}}$ is turbulence—it may have comparable energy but lacks directional coherence, so the overall trajectory remains northward.
D.4 ICA Intuition: Why Independence Helps
Setup: Suppose representations encode two independent factors:
  • $z_{1}$: formality (independent)
  • $z_{2}$: politeness (independent)
And mix them:
$h=v_{1}z_{1}+v_{2}z_{2}$ (51)
Problem without independence.
If $z_{1}$ and $z_{2}$ are correlated (e.g., formal text tends to be polite), we can reparameterize:
$\tilde{z}_{1}=z_{1}+\alpha z_{2},\quad\tilde{z}_{2}=z_{2}$ (52)
$h=(v_{1}-\alpha v_{2})\tilde{z}_{1}+(v_{2}+\alpha v_{1})\tilde{z}_{2}$ (53)
giving infinitely many equivalent decompositions (different $\alpha$ values).
Solution with independence.
If $z_{1}\perp z_{2}$ (statistically independent), then $\tilde{z}_{1}=z_{1}+\alpha z_{2}$ is NOT independent of $\tilde{z}_{2}=z_{2}$ for $\alpha\neq 0$:
$I(\tilde{z}_{1};\tilde{z}_{2})=I(z_{1}+\alpha z_{2};z_{2})>0$ (54)
ICA finds the unique decomposition where sources are independent (refer Figure 7), breaking the symmetry.
$z_{1}$$z_{2}$Many rotations OK
(a) Correlated sources
$z_{1}$$z_{2}$Only THIS rotation
(b) Independent sources
Figure 7: Visual comparison of correlated versus independent sources.
D.5 Sparsity Intuition: Occam’s Razor
Setup: Suppose we observe:
$y=Jv+\eta$ (55)
and both $v_{1}$ and $v_{2}$ fit the data:
$\|Jv_{1}-y\|\approx\|Jv_{2}-y\|\approx\epsilon$ (56)
Question: Which is the “true” vector?
Sparsity principle.
Prefer the simplest explanation. If:
  • $v_{1}$ has 100 non-zero entries
  • $v_{2}$ has 10 non-zero entries
then $v_{2}$ is more likely to be the true sparse signal.
Why this works.
If the true vector is sparse, then dense solutions must include spurious noise components:
$v_{\text{dense}}=v_{\text{true}}+v_{\text{noise}}$
The $\ell_{1}$ penalty penalizes density, filtering out $v_{\text{noise}}$.
D.6 Multi-Environment Intuition: Finding What’s Stable
Setup: Observe “formality” steering in 3 environments:
  • E1 (academic): Formality correlates with technical jargon
  • E2 (business): Formality correlates with professional tone
  • E3 (legal): Formality correlates with precise language
Spurious correlation problem.
Single-environment extraction might learn:
$v_{1}$ $=v_{\text{formality}}+v_{\text{jargon}}$
$v_{2}$ $=v_{\text{formality}}+v_{\text{professional}}$
$v_{3}$ $=v_{\text{formality}}+v_{\text{precise}}$
Each includes the true factor plus environment-specific confounds.
Invariance solution.
Find the component present in all environments:
$v_{\text{invariant}}=\text{stable component of }\{v_{1},v_{2},v_{3}\}$ (57)
The spurious parts ($v_{\text{jargon}},v_{\text{professional}},v_{\text{precise}}$) vary across environments and get filtered out.
E1: formality + jargonE2: formality + professionalE3: formality + preciseIntersectionformality (stable core)
Figure 8: Multi-environment filtering extracts the stable core. Each environment contains the true semantic factor (formality) plus environment-specific confounds. The intersection across environments isolates the invariant component.
D.7 Cross-Layer Intuition: Consistency Check
Setup: Extract steering vectors at layers 8, 12, 16, 20.
Consistency hypothesis.
True semantic factors should propagate coherently:
$v_{12}\approx W_{8}v_{8},\quad v_{16}\approx W_{12}v_{12},\quad\text{etc.}$
Inconsistency indicates null-space artifacts.
If $v_{8}$ includes a large null-space component $v_{\perp,8}$:
$v_{8}=v_{\text{true},8}+v_{\perp,8}$ (58)
Then after propagation:
$W_{8}v_{8}=W_{8}v_{\text{true},8}+W_{8}v_{\perp,8}$ (59)
Since $v_{\perp,8}\in\ker(J_{8})$ but generically $W_{8}v_{\perp,8}\notin\ker(J_{12})$, the null-space component becomes observable at layer 12, causing inconsistency.
[…]
Figure 9: Cross-layer consistency check. True semantic factors propagate coherently, while null-space artifacts become inconsistent.
Filtering strategy.
Project onto the subspace of vectors that are consistent across layers:
$v_{\text{consistent}}=\arg\min_{v}\sum_{\ell}\|W_{\ell-1}v_{\ell-1}-v_{\ell}\|^{2}$ (60)
This filters out layer-specific artifacts (See Figure 9), retaining only the stable semantic direction.
Appendix E Mathematical Derivations
E.1 Dimension of Observational Equivalence Class
Setup: For steering vector $v\in\mathbb{R}^{d}$ and Jacobian $J\in\mathbb{R}^{V\times d}$ with rank $r$, consider:
$v]=\{v^{\prime}\in\mathbb{R}^{d}:Jv^{\prime}=Jv\}$ (61)
Question: How many degrees of freedom exist in the equivalence class of observationally equivalent steering vectors?
Analysis: The equivalence class is an affine subspace:
$v]=v+\ker(J)$ (62)
where $\ker(J)$ is a (dr)(d-r)-dimensional linear subspace.
Parameterization: Let $\{u_{1},\ldots,u_{d-r}\}$ be an orthonormal basis for $\ker(J)$. Then:
$v]=\left\{v+\sum_{i=1}^{d-r}\alpha_{i}u_{i}:\alpha_{i}\in\mathbb{R}\right\}$ (63)
The equivalence class has (dr)(d-r) degrees of freedom.
Scaling ambiguity: Additionally, $v$ and $cv$ produce equivalent outputs (up to rescaling $\alpha$). The equivalence class modulo scaling is a (dr)(d-r)-dimensional projective space.
Illustrative example. Suppose a model has hidden size $d=4096$ and an effective Jacobian rank $r\approx 3100$. Then the equivalence class dimension is $d-r\approx 996$.
Interpretation: For every identifiable parameter (row-space direction), there are $996/3100\approx 0.32$ unidentifiable parameters (null-space directions). The under-determination ratio is approximately $1:3$.
E.2 Fisher Information and Cramér-Rao Bound
Setup: Consider the statistical model:
$o(x;v,\alpha)\sim p(o|x,v,\alpha)$ (64)
where $v$ is the parameter to estimate.
Question: Can we achieve better identifiability by collecting more samples?
Fisher Information Matrix: For the linear Gaussian model $o=Jv+\eta$ with $\eta\sim\mathcal{N}(0,\sigma^{2}I)$:
$\mathcal{I}(v)$ $=\mathbb{E}_{o}\Bigg[\left(\frac{\partial\log p(o\mid x,v,\alpha)}{\partial v}\right)$ (65)
$\qquad\quad\left(\frac{\partial\log p(o\mid x,v,\alpha)}{\partial v}\right)^{\!\top}\Bigg]$
$=\frac{1}{\sigma^{2}}J^{\top}J.$
Cramér-Rao Lower Bound: The covariance of any unbiased estimator $\hat{v}$ satisfies:
$\text{Cov}(\hat{v})\succeq\mathcal{I}(v)^{-1}=\sigma^{2}(J^{\top}J)^{+}$ (66)
where $(J^{\top}J)^{+}$ is the Moore-Penrose pseudo-inverse.
Implications: For null-space directions $u_{i}\in\ker(J)$, the Fisher information is degenerate:
$u_{i}^{\top}(J^{\top}J)u_{i}=0$ (67)
Therefore, the variance of any unbiased estimator along null-space directions is unbounded:
$\text{Var}(u_{i}^{\top}\hat{v})\geq u_{i}^{\top}\mathcal{I}(v)^{-1}u_{i}=\infty$ (68)
Conclusion: The Cramér-Rao bound is infinite for null-space components, confirming that no finite amount of data can resolve the ambiguity. Non-identifiability is fundamental, not a small-sample problem. More data cannot help because the information geometry has infinite uncertainty in null-space directions.
E.3 Proof that ICA Breaks Gauge Symmetry
Setup: Consider two decompositions:
$h=A_{1}z_{1}=A_{2}z_{2}$ (69)
where both $z_{1}$ and $z_{2}$ have independent components.
Question: When are these equivalent ($A_{1}=A_{2}$ up to permutation/scaling)?
Analysis: If both decompositions are valid:
$A_{1}z_{1}=A_{2}z_{2}\implies z_{1}=A_{1}^{-1}A_{2}z_{2}=Gz_{2}$ (70)
where $G=A_{1}^{-1}A_{2}$.
ICA Constraint: If both $z_{1}$ and $z_{2}$ have independent components:
$I(z_{1,i};z_{1,j})$ $=0$ $\text{for }i\neq j,$ (71)
$I\big((Gz_{2})_{i};(Gz_{2})_{j}\big)$ $=0$ $\text{for }i\neq j.$
Key Theorem (Comon, 1994): If $z_{2}$ has independent components and $Gz_{2}$ also has independent components, then $G$ must be a generalized permutation matrix:
$G=PD$ (72)
where $P$ is a permutation matrix and $D$ is a diagonal scaling matrix.
Implication:
$A_{1}z_{1}$ $=A_{2}z_{2},$ (73)
$A_{1}Gz_{2}$ $=A_{2}z_{2},$
$A_{1}PD$ $=A_{2}.$
Therefore $A_{2}=A_{1}PD$, meaning $A_{2}$ is $A_{1}$ with permuted and scaled columns.
Conclusion: ICA constraints (statistical independence) force $G$ to be a permutation-scaling matrix, breaking arbitrary gauge transformations and establishing identifiability up to unavoidable symmetries. Independence is the structural assumption that eliminates the infinite equivalence class.
Appendix F Disclosure of the Use of AI
We used ChatGPT to assist with refining sections of the manuscript for language and clarity and used Claude Code and GitHub Copilot for coding assistance during implementation. All research ideas, technical content, proofs, experiments and final writing decisions are entirely our own and we take full responsibility for the work presented.

### v2

@@
-On the Identifiability of Steering Vectors in Large Language Models
+On the Non-Identifiability of Steering Vectors in Large Language Models
 Abstract
-Activation steering methods, such as persona vectors, are widely used to control large language model behavior and increasingly interpreted as revealing meaningful internal representations.
-This interpretation implicitly assumes steering directions are identifiable and uniquely recoverable from input–output behavior.
-We formalize steering as an intervention on internal representations and prove that, under realistic modeling and data conditions, steering vectors are fundamentally non-identifiable due to large equivalence classes of behaviorally indistinguishable interventions.
-Empirically, we validate this across multiple models and semantic traits, showing orthogonal perturbations achieve near-equivalent efficacy with negligible effect sizes.
-However, identifiability is recoverable under structural assumptions including statistical independence, sparsity constraints, multi-environment validation or cross-layer consistency.
-These findings reveal fundamental interpretability limits and clarify structural assumptions required for reliable safety-critical control.
-On the Identifiability of Steering Vectors in Large Language Models
-Sohan Venkatesh  and Ashish Mahendran Kurapath Manipal Institute of Technology Bengaluru {sohan1, ashish}.mitblr2022@learner.manipal.edu
+Activation steering methods are widely used to control large language model (LLM) behavior and are often interpreted as revealing meaningful internal representations.
+This interpretation assumes steering directions are identifiable and uniquely recoverable from input–output behavior.
+We show that, under white-box single-layer access, steering vectors are fundamentally non-identifiable due to large equivalence classes of behaviorally indistinguishable interventions.
+Empirically, we show that orthogonal perturbations achieve near-equivalent efficacy with negligible effect sizes across multiple models and traits.
+Critically, we show that the non-identifiability is a robust geometric property that persists across diverse prompt distributions.
+These findings reveal fundamental interpretability limits and highlight the need for structural constraints beyond behavioral testing to enable reliable alignment interventions.
+[^1] Code: https://github.com/sohv/non-identifiability
+On the Non-Identifiability of Steering Vectors in Large Language Models
+Sohan Venkateshthanks: Correspondence: soh.venkatesh@gmail.com     Ashish Mahendran Kurapath Manipal Institute of Technology Bengaluru
 1 Introduction
-Persona vector steering has emerged as a popular technique for controlling the behavior of large language models by adding learned directional vectors to intermediate activations.
-Empirically, such vectors can shift model outputs along interpretable dimensions such as politeness, political ideology or truthfulness, suggesting that representational alignment might afford fine-grained behavioral control without retraining ((Zou et al., 2023; Rimsky et al., 2024; Turner et al., 2023)).
-This line of work is closely connected to broader efforts in representation engineering and activation editing, where linear directions in activation space are used to modulate model behavior along semantic axes ((Elhage et al., 2022; Turner et al., 2024)).
+Persona vector steering (Chen et al., 2025) has emerged as a popular technique for controlling the behavior of large language models by adding learned directional vectors to intermediate activations.
+This line of work is closely connected to broader efforts in representation engineering (Zou et al., 2023) and activation editing (Rimsky et al., 2024; Turner et al., 2023), where linear directions in activation space are used to modulate model behavior along semantic axes (Elhage et al., 2022; Turner et al., 2024).
 Despite growing adoption in interpretability and alignment research, the theoretical foundations of persona steering remain poorly understood.
-Most existing methods implicitly assume that extracted steering vectors correspond to meaningful, uniquely determined latent factors—for example, "the politeness direction" or "the honesty direction"—and that these factors can be directly manipulated to achieve reliable control.
-However, classical results in latent variable modeling and causal inference show that such assumptions are often unjustified without additional structural constraints ((Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Schölkopf et al., 2021; Locatello et al., 2019)).
-Nonlinear ICA and causal representation learning further emphasize that recovering latent factors from high-dimensional observations is generically impossible without auxiliary information or strong inductive biases ((Hyvarinen and Morioka, 2017; Khemakhem et al., 2020; Ahuja et al., 2022)).
-This raises a fundamental question for alignment and interpretability research: when does representational alignment genuinely afford reliable behavioral control and when does it merely exploit underdetermined or spurious correlations?
-Existing evidence from probing and representation analysis suggests that even seemingly meaningful linear directions can reflect artifacts of the measurement procedure rather than uniquely defined semantic variables ((Hewitt and Liang, 2019; Ravfogel et al., 2020; Elazar et al., 2021; Belinkov, 2022)).
-In the context of activation steering, these concerns become particularly acute, because the resulting vectors are often used not only for analysis but also for safety-relevant control interventions.
-Understanding identifiability is therefore critical for several reasons:
-Alignment affordances.
-If persona vectors are not identifiable, then representational alignment may provide only heuristic control rather than principled intervention.
-Characterizing when steering vectors are unique clarifies when alignment interventions can be trusted and when they should instead be viewed as exploiting one of many behaviorally equivalent directions.
-Interpretability validity.
-When multiple incompatible vectors produce identical observable behavior, claims that a specific vector "represents" a semantic concept become scientifically underdetermined.
-Identifiability theory distinguishes well-grounded interpretability claims from artifacts of measurement and projection ((Elazar et al., 2021; Marks and Tegmark, 2023)).
-Robustness and safety.
-Non-identifiable steering directions may rely on fragile correlations that fail under distribution shift, model updates or adversarial prompting.
-For safety-critical applications—where steering vectors may be used to enforce norms or prevent harmful behavior—understanding identifiability limits is essential to avoid brittle or misleading forms of control.
-Methodological design.
-Identifiability theory clarifies which experimental protocols provide meaningful evidence and which require additional structure to support reliable conclusions.
-In particular, it highlights when interventions on internal activations can be interpreted causally and when they merely reparameterize an equivalence class of representations ((Schölkopf et al., 2021; Ahuja et al., 2022)).
-In this work, we provide, to our knowledge, the first formal identifiability analysis of persona vector steering.
-Our contributions are threefold.
-First, we prove that under standard observational regimes, persona vectors are generically non-identifiable due to null-space ambiguity in the model’s input-output Jacobian (Proposition 1).
-Specifically, under white-box single-layer access, infinitely many geometrically distinct steering directions induce identical observable behavior.
-Second, we characterize sufficient structural conditions under which identifiability can be recovered including statistical independence constraints, sparsity priors, multi-environment access and cross-layer consistency (Proposition 2).
-Third, we demonstrate empirically that contemporary steering operates in the non-identifiable regime across multiple models and semantic traits, with vectors perturbed by orthogonal components achieving near-equivalent efficacy.
-Crucially, our results are not purely negative.
-By explicitly characterizing both the limits and the affordances of persona vector steering, we provide principled guidance for when representational alignment can be trusted and which structural assumptions enable reliable control in safety-critical applications.
+Most existing methods implicitly assume that extracted steering vectors correspond to meaningful, uniquely determined latent factors and that these factors can be directly manipulated to achieve reliable control.
+However, classical results in latent variable modeling and causal inference show that such assumptions are often unjustified without additional structural constraints (Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Schölkopf et al., 2021; Locatello et al., 2019).
+This raises a fundamental question: when does representational alignment afford reliable behavioral control?
+Understanding identifiability is critical for several reasons.
+First, it determines when persona vectors provide principled causal interventions versus merely heuristic control that exploits one of many behaviorally equivalent directions.
+Second, it clarifies when claims that a specific vector “represents” a semantic concept are scientifically well-grounded rather than artifacts of measurement and projection (Elazar et al., 2021; Marks and Tegmark, 2023).
+Third, it reveals when steering directions may rely on fragile correlations that fail under distribution shift, model updates or adversarial prompting.
+Finally, identifiability theory clarifies which experimental protocols provide meaningful evidence for causal interpretation of interventions on internal activations (Schölkopf et al., 2021; Ahuja et al., 2022).
+Figure 1: Geometric illustration of steering vector non-identifiability.
+Two geometrically distinct vectors, the original steering vector $v$ (left) and a perturbed version $v+v_{\perp}$ where $v_{\perp}\in\mathrm{ker}(J)$ (right), produce identical observable outputs.
+The null-space component (red, unobservable region) can be arbitrarily modified without changing outputs, demonstrating the existence of large equivalence classes of behaviorally indistinguishable steering directions.
+We make three contributions.
+First, we provide a formal identifiability analysis of persona vector steering, proving that under white-box single-layer access, persona vectors are fundamentally non-identifiable: infinitely many geometrically distinct directions induce identical observable behavior via null-space ambiguity.
+Second, we empirically demonstrate that extracted steering vectors contain substantial non-identifiable components, with orthogonal perturbations achieving 95–100% of the original steering efficacy across multiple models and traits.
+Third, we validate through multi-environment experiments that this equivalence persists under distribution shift, confirming non-identifiability is a robust geometric property rather than an artifact of limited prompt diversity.
+These results clarify when alignment methods can be trusted and which structural assumptions are required for reliable control.
 2 Related Work
-Our work formalizes persona steering as a latent variable identification problem, bridging causal representation learning, activation editing in LLMs and mechanistic interpretability.
+Our work formalizes persona steering  [^2] We use persona vectors and steering vectors interchangeably to refer to directional vectors added to model activations to modulate semantic or behavioral attributes (Chen et al., 2025).
+as a latent variable identification problem, bridging causal representation learning, activation editing in LLMs and mechanistic interpretability.
 Causal and latent variable identifiability.
-Classical results show that latent variable models are generically non-identifiable without structural assumptions (Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Kruskal, 1977).
+Classical results show that latent variable models are non-identifiable without structural assumptions (Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Kruskal, 1977).
 Recent work in causal representation learning extends these ideas to deep learning settings (Schölkopf et al., 2021; Ahuja et al., 2022; Locatello et al., 2019), establishing conditions under which latent factors can be recovered from high-dimensional observations.
-Nonlinear ICA methods (Khemakhem et al., 2020; Hyvarinen and Morioka, 2017) provide theoretical foundations for recovering latent variables using temporal or auxiliary information, which standard steering methods do not exploit.
 Probing and representation learning.
-Pimentel et al.
-(2020) and Ravfogel et al.
-(2020) demonstrate that probing classifiers can succeed for trivial reasons unrelated to the presence of target information.
-Elazar et al.
-(2021) show that removing information via projection is ill-defined without identifiability guarantees.
-Belinkov (2022) provides a comprehensive survey of probing methods and their limitations.
-The linear representation hypothesis (Park et al., 2023; Elhage et al., 2022; Marks and Tegmark, 2023) suggests that concepts correspond to directions in activation space but lacks formal identifiability guarantees.
+Work on probing classifiers (Pimentel et al., 2020; Ravfogel et al., 2020; Elazar et al., 2021) shows that linear directions can succeed for reasons unrelated to target information.
+The linear representation hypothesis (Park et al., 2023; Elhage et al., 2022) suggests concepts correspond to directions in activation space but lacks identifiability guarantees.
 Activation editing in LLMs.
-Methods such as representation engineering (Zou et al., 2023), contrastive activation addition (Rimsky et al., 2024; Turner et al., 2024), activation patching (Meng et al., 2022) and causal tracing (Wang et al., 2022) manipulate model internals to control behavior.
-While empirically successful, these works generally do not address whether the resulting control directions are uniquely determined.
-Related work on linear mode connectivity (Entezari et al., 2021) and neural network reparameterization (Dinh et al., 2017) reveals symmetries in neural representations that motivate our null-space analysis.
-Steering and control vectors.
-Early work on steering vectors (Li et al., 2022; Turner et al., 2023; Tigges et al., 2023) develops techniques for extracting control directions from contrastive prompt pairs.
-Burns et al.
-(2022) investigate discovering latent knowledge in language models.
-These methods assume that extracted vectors correspond to uniquely determined semantic factors—an assumption we formally examine.
+Methods such as representation engineering (Zou et al., 2023), contrastive activation addition (Rimsky et al., 2024; Turner et al., 2024) and activation patching (Meng et al., 2022) manipulate model internals but do not address whether control directions are uniquely determined.
+Work on neural network symmetries (Entezari et al., 2021; Dinh et al., 2017) motivates our null-space analysis.
+Robustness and invariant representations.
+Work on distribution shift and domain adaptation (Rabanser et al., 2019; Lipton et al., 2018) establishes that learned representations often exploit spurious correlations.
+Methods for learning invariant representations (Arjovsky et al., 2019; Krueger et al., 2021) aim to extract features stable across domains.
+Recent work on multi-environment causal representation learning (Ahuja et al., 2022; Von Kügelgen et al., 2021; Rojas-Carulla et al., 2018) proves that diverse training environments enable identification of causal factors under structural assumptions.
 Mechanistic interpretability.
@@
 Our theoretical analysis clarifies when such directions are well-defined.
-Compressed sensing and sparse recovery.
-The theoretical framework of compressed sensing (Candès and Wakin, 2008; Donoho, 2006) provides conditions under which sparse vectors can be uniquely recovered from linear measurements.
-These results directly inform the sparsity-based identifiability conditions in Proposition 2.
 3 Problem Setup
@@
 A steering vector $v\in\mathbb{R}^{d}$ is applied as:
-$\tilde{h}\ell(x)=h\ell(x)+\alpha v,$
+$\tilde{h}\ell(x)=h\ell(x)+\alpha v$ (1)
 where $\alpha\in\mathbb{R}$ is the steering strength.
@@
 We posit that the true data-generating process involves:
-$z\sim p(z),\quad h_{\ell}=g_{\ell}(x,z)+\epsilon,$
+$z\sim p(z),\quad h_{\ell}=g_{\ell}(x,z)+\epsilon$ (2)
 where $g_{\ell}:\mathcal{X}\times\mathcal{Z}\to\mathbb{R}^{d}$ encodes how persona $z$ modulates the representation for prompt $x$ and $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ is measurement noise.
@@
 Most existing work operates in Regime 2, often extracting a steering vector from contrastive prompt pairs $(x^{+},x^{-})$ designed to elicit different persona values (e.g., polite versus rude instructions):
-$v\propto\mathbb{E}{x^{+}}[h\ell(x^{+})]-\mathbb{E}{x^{-}}[h\ell(x^{-})].$
+$v\propto\mathbb{E}{x^{+}}[h\ell(x^{+})]-\mathbb{E}{x^{-}}[h\ell(x^{-})]$ (3)
 Our analysis focuses primarily on Regime 2, as it represents the dominant paradigm in current steering research.
@@
 Near a reference distribution, we can approximate the effect of steering on output logits as:
-$o(x,v,\alpha)\approx o(x,0,0)+\alpha J_{\ell}(x)v,$
-where $J_{\ell}(x)=\frac{\partial o}{\partial h_{\ell}}\big|{h\ell(x)}\in\mathbb{R}^{V\times d}$ is the Jacobian and $V$ is the vocabulary size.
+$o(x,v,\alpha)\approx o(x,0,0)+\alpha J_{\ell}(x)v$ (4)
+where $J_{\ell}(x)=\left.\frac{\partial o}{\partial h_{\ell}}\right|_{h_{\ell}(x)}\in\mathbb{R}^{V\times d}$ denotes the Jacobian of the output and $V$ is the vocabulary size.
 Nonlinear case.
 In general, the mapping $h_{\ell}\mapsto o$ involves multiple nonlinear layers (attention, MLPs, layer norms). We denote this as:
-$o=F_{\ell\to L}(h_{\ell}+\alpha v),$
+$o=F_{\ell\to L}(h_{\ell}+\alpha v)$ (5)
 where $F_{\ell\to L}$ is the composition of layers $\ell+1$ through $L$.
@@
 5 Main Results
-We now state our main theoretical results, characterizing when observational conditions afford identifiable persona vectors and when they do not.
-5.1 Proposition 1: Non-Identifiability Without Structural Constraints
+We now state our main theoretical result, proving that persona vectors are fundamentally non-identifiable under standard observational regimes.
 Proposition 1.
@@
 Proof Sketch.
-We establish non-identifiability via two complementary arguments: reparameterization symmetry and null-space ambiguity.
+We establish non-identifiability via two complementary arguments: null-space ambiguity and reparameterization symmetry.
+Null-space ambiguity (primary mechanism).
+Under the local linear approximation $o\approx o_{0}+\alpha J_{\ell}v$, any vector $v^{\prime}=v+v_{0}$ where $v_{0}\in\ker(J_{\ell})$ satisfies $J_{\ell}v^{\prime}=J_{\ell}v$, and thus produces identical linearized outputs.
+This establishes non-identifiability even without invoking global reparameterization symmetries.
 Reparameterization symmetry.
-Consider an invertible transformation $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and define the reparameterized representation $h^{\prime}\ell(x)=T(h\ell(x))$.
-If the subsequent layers can be equivalently expressed as $F_{\ell\to L}(h_{\ell})=F^{\prime}{\ell\to L}(T(h\ell))$, then using $(h_{\ell},v)$ or $(h^{\prime}_{\ell},v^{\prime})$ where $v^{\prime}=T(v)$ is observationally indistinguishable.
-For overparameterized networks, many such reparameterizations exist that leave the input-output function invariant.
-Consider $T(h)=Ah+b$ where $A\in\mathbb{R}^{d\times d}$ is invertible.
-This transformation can be absorbed into the parameters of layer $\ell+1$ through weight reparameterization.
-While we cannot explicitly construct these transformations for a given frozen model without retraining, their existence follows from the network’s inherent symmetry structure.
-Let $A$ be any invertible matrix with $A\neq cI$ for any scalar $c$.
-Then $v^{\prime}=Av$ is not proportional to $v$, yet produces equivalent observations after appropriate reparameterization.
-Since there are infinitely many such $A$, there exist infinitely many equivalent $v^{\prime}$.
-Null-space ambiguity (primary mechanism).
-Under the local linear approximation $o\approx o_{0}+\alpha J_{\ell}v$, any vector $v^{\prime}=v+v_{0}$ where $v_{0}\in\ker(J_{\ell})$ satisfies $J_{\ell}v^{\prime}=J_{\ell}v$ and thus produces identical linearized outputs.
-This establishes non-identifiability even without invoking global reparameterization symmetries.
-Corollary 1.1.
-Under the linear approximation $o\approx o_{0}+\alpha J_{\ell}v$, any vector $v^{\prime}=v+v_{0}$ where $v_{0}\in\ker(J_{\ell})$ is observationally equivalent to $v$.
+Neural networks exhibit inherent parameter symmetries arising from overparameterization.
+Consider an invertible linear transformation $T(h)=Ah$ where $A\in\mathbb{R}^{d\times d}$.
+For layer $\ell+1$ with weight matrix $W_{\ell+1}$ and bias $b_{\ell+1}$, the computation $h_{\ell+1}=\sigma(W_{\ell+1}h_{\ell}+b_{\ell+1})$ can be equivalently expressed through reparameterized weights $W^{\prime}_{\ell+1}=W_{\ell+1}A^{-1}$:
+$h^{\prime}_{\ell+1}$ $=\sigma\!\left(W^{\prime}_{\ell+1}Ah_{\ell}+b_{\ell+1}\right)$ (7)
+$=\sigma\!\left(W_{\ell+1}h_{\ell}+b_{\ell+1}\right)$
+$=h_{\ell+1}$
+Under this reparameterization, steering with $v^{\prime}=Av$ on the transformed representation $h^{\prime}_{\ell}=Ah_{\ell}$ produces identical outputs to steering with $v$ on $h_{\ell}$.
+For any invertible matrix $A\neq cI$ (not a scalar multiple of identity), we obtain $v^{\prime}\not\propto v$, yet these vectors are observationally equivalent.
+Since the space of such matrices has dimension $d^{2}-1$, infinitely many distinct reparameterizations exist.
+Practical vs.
+theoretical relevance: The reparameterization argument establishes non-identifiability as a fundamental property of the model class, independent of any specific instantiation.
+While these symmetries cannot be exploited in frozen models without modifying weights, their existence proves that identifiability fails even in principle.
+For analyzing deployed models, the null-space mechanism provides direct, operationally relevant non-identifiability.
+Corollary 1.1 (Null-space equivalence).
+Under the linear approximation $o\approx o_{0}+\alpha J_{\ell}v$, any vector $v^{\prime}=v+v_{0}$ where $v_{0}\in\ker(J_{\ell})$ is observationally equivalent to $v$, since $J_{\ell}v^{\prime}=J_{\ell}v+J_{\ell}v_{0}=J_{\ell}v$.
 Remark (Null-space dimensionality).
@@
 For a Jacobian $J_{\ell}\in\mathbb{R}^{V\times d}$ with vocabulary size $V$ and hidden dimension $d$, the maximum possible rank is $\min(V,d)$.
-In modern language models, $V\approx 50000$ and $d\approx 4000$, so max rank is $d$.
-However, the output distribution lies on a low-dimensional manifold, causing $\text{rank}(J_{\ell})\ll d$ in practice, consistent with observations that neural network representations concentrate on low-dimensional subspaces (Maennel et al., 2018; Li et al., 2018).
-This yields $\dim(\ker(J_{\ell}))=d-\text{rank}(J_{\ell})\gg 0$.
-This result establishes generic non-identifiability under local linear approximation.
-5.2 Proposition 2: Identifiable Regimes Under Structural Assumptions
-Proposition 2.
-Persona vectors can be identified up to scaling and permutation, thus affording reliable alignment control, under the following sufficient structural conditions:
-  • Statistical Independence (ICA).
-If the latent persona $z=(z_{1},\ldots,z_{k})$ has independent components and $h_{\ell}=Az+\epsilon$ where $A\in\mathbb{R}^{d\times k}$ is a mixing matrix, then under non-Gaussianity of $z_{i}$ and sufficient observations, AA (and hence the columns $v_{i}$) can be recovered up to permutation and scaling (Comon, 1994; Hyvärinen and Oja, 2000).
-  • Sparsity Constraints.
-If the true persona vector $v$ is sparse (i.e., $|v|_{0}\leq s\ll d$) and we observe the effect of steering on multiple outputs, then under restricted isometry properties, $v$ can be uniquely recovered via $\ell_{1}$ minimization(Candes and Tao, 2005).
-  • Multi-Environment or Interventional data.
-If we observe the same persona $z$ across multiple contexts (prompts, models or layers) where the spurious correlations change but the true signal remains stable, then techniques from causal representation learning allow identification of invariant factors.
-(Peters et al., 2017; Ahuja et al., 2022).
-  • Cross-layer Consistency.
-If we assume persona vectors have consistent geometric relationships across multiple layers (e.g., $v_{\ell}\approx W_{\ell}v_{\ell-1}$ for known or estimable $W_{\ell}$), then the overdetermined system provides additional constraints that break symmetries.
-Interpretation.
-Proposition 2 serves as a natural theoretical extension of Proposition 1: by characterizing sufficient conditions under which identifiability could be recovered–such as statistical independence (ICA), sparsity priors, multi-environment access or cross-layer consistency.
-Proposition 2 clarifies precisely which structural assumptions are necessary to break the symmetries that cause non-identifiability.
+In modern language models, $V\approx 50{,}000$ and $d\approx 4{,}000$, so the maximal rank is $d$.
+However, empirical evidence shows that (1) neural network representations exhibit low intrinsic dimensionality (Li et al., 2018; Maennel et al., 2018), and (2) output sensitivity concentrates in a small number of Hessian directions (Ghorbani et al., 2019).
+These findings suggest the Jacobian $J_{\ell}$ has effective rank substantially smaller than $d$, implying a high-dimensional null space:
+$\dim(\ker(J_{\ell}))=d-\mathrm{rank}(J_{\ell})$ (8)
 6 Empirical Validation
-We now validate that the non-identifiability characterized in Proposition 1 (Section 5.1) manifests in contemporary language models.
+We now validate that the non-identifiability characterized in Proposition 1 (Section 5) manifests in contemporary language models.
 Our empirical experiments test a behavioral consequence of the theoretical prediction: the existence of large equivalence classes of semantically indistinguishable steering directions, without directly estimating the Jacobian null space.
@@
 Models and Layers.
-We evaluate two open-weight instruction-tuned language models to test generality across architectures and scales:
-  • Qwen2.5-3B-Instruct: 24 layers, hidden dimension $d=2048$, layer $\ell=12$ (mid-network)
-  • Llama-3.1-8B-Instruct: 32 layers, hidden dimension $d=4096$, layer $\ell=16$ (mid-network)
-For both models, we focus on middle layers ($\ell=L/2$), following standard practice (Chen et al., 2025; Konen et al., 2024; Sun et al., ) in steering research.
-Middle layers balance semantic abstraction with steerability: early layers encode low-level features (tokens, syntax), while late layers specialize for next-token prediction.
+We evaluate two open-weight instruction-tuned language models to test generality across architectures and scales: Qwen2.5-3B-Instruct (layer $\ell=12$) and Llama-3.1-8B-Instruct (layer $\ell=16$).
+For both models, we focus on middle layers ($\ell=L/2$), following standard practice (Chen et al., 2025; Konen et al., 2024; Sun et al., 2025) in steering research.
 Persona traits.
-We test three semantic traits spanning distinct dimensions:
-  • Formality: Formal versus informal style
-  • Politeness: Polite versus rude social register
-  • Humor: Humorous versus serious content
+We test three semantic traits spanning distinct dimensions: Formality, Politeness and Humor.
 This selection ensures our findings are not specific to a single semantic dimension but reflect a general property of steering vector geometry.
@@
 The baseline steering vector is computed as the mean difference:
-$v=\frac{1}{50}\sum_{i=1}^{50}[h_{\ell}(x^{+}i)-h\ell(x^{-}_{i})].$
+$v=\frac{1}{50}\sum_{i=1}^{50}[h_{\ell}(x^{+}i)-h\ell(x^{-}_{i})]$ (9)
 Semantic probes.
 We evaluate semantic equivalence using trait-specific scoring functions $\phi(o)$ that map generated text to real-valued scores.
-This provides a conservative test of observational equivalence: if vectors are distinguishable semantically, they are certainly non-identical in the full distributional sense.
 For formality, politeness and humor, we use lexical heuristics based on formal/informal markers, polite/rude markers and humorous/serious markers respectively, combined with sentence length and stylistic features.
@@
 6.2 Orthogonal Perturbation Test Results
-Table 1 presents our empirical findings across two models, three semantic traits and two sample sizes ($n=5$ and $n=10$ random orthogonal seeds).
-Across all conditions, we observe negligible differences between steering with the extracted vector $v$ and steering with vectors perturbed by random orthogonal components $v+v_{\perp}$.
-We measure steering efficacy using semantic probe scores in $[0,1]$ that quantify the intensity of the target trait in generated outputs, comparing score distributions from steered generations against baseline generations.
 Table 1: Empirical validation of non-identifiability across models, traits and sample sizes.
@@
 $n=10$ $119\pm 0.119 0.0$$016\pm 0.104$ $95.9\%$
+Table 1 presents our empirical findings across two models, three semantic traits and two sample sizes ($n=5$ and $n=10$ random orthogonal seeds).
+Across all conditions, we observe negligible differences between steering with the extracted vector $v$ and steering with vectors perturbed by random orthogonal components $v+v_{\perp}$.
+We measure steering efficacy using semantic probe scores in $[0,1]$ that quantify the intensity of the target trait in generated outputs, comparing score distributions from steered generations against baseline generations.
 With $n=10$ seeds, the mean Cohen’s $d$ is 0.080 for Qwen2.5-3B and 0.100 for Llama-3.1-8B, both well below the threshold for small effects ($d=0.2$) and firmly in the negligible range.
@@
 For Qwen2.5-3B, mean differences remain below $0.022$ across all $\alpha$ values (less than $2.5\%$ deviation in formality scores); for Llama-3.1-8B, differences remain below $0.023$ ($2.5\%$ deviation).
-This confirms that non-identifiability is a structural property: any vector in the equivalence class $v+\ker(J)$ produces identical observations regardless of scaling.
+This proves that non-identifiability is a structural property: any vector in the equivalence class $v+\ker(J)$ produces identical observations regardless of scaling.
+6.4 Robustness of Non-Identifiability Under Distribution Shift
+We test whether equivalence classes persist across distributional shifts or whether prompt diversity alone induces identifiability.
+We evaluate orthogonal perturbation equivalence across four distinct prompt environments: in-distribution (standard trait templates), topic shift (medical, legal, creative writing contexts), genre shift (emails, tweets, academic writing), and safety-style (refusal patterns, ethical dilemmas).
+For each environment, we extract steering vectors and evaluate whether $v\equiv v+v_{\perp}$ equivalence holds when testing on that environment’s prompts.
+Table 2 and Figure 4 present Cohen’s $d$ effect sizes across all trait-environment combinations for both models.
+Across 24 conditions, we observe effect sizes ranging from $d=0.21$ to $d=0.63$ (mean $d=0.35$).
+Effect sizes are predominantly small ($d<0.5$ for 88% of conditions; mean $d=0.35$), indicating that $v$ and $v+v_{\perp}$ remain largely observationally equivalent across diverse extraction contexts.
+Table 2: Multi-environment validation. Cohen’s $d$ effect sizes across extraction contexts. ID = in-distribution; Top = topic shift; Gen = genre shift; Saf = safety-style prompts.
+Model Trait ID Top Gen Saf
+Qwen-3B Formality 0.44 0.25 0.29 0.21
+Politeness 0.22 0.24 0.52 0.41
+Humor 0.42 0.51 0.24 0.51
+Llama-8B Formality 0.30 0.54 0.30 0.21
+Politeness 0.37 0.27 0.51 0.63
+Humor 0.29 0.56 0.30 0.40
+Figure 4: Multi-environment validation results.
+Environment-dependent effect sizes and cross-environment matrices show small-to-moderate Cohen’s $d$ values across traits and prompt contexts, confirming robustness of observational equivalence under distributional shift.
+Formality shows most consistency (mean $d=0.32$), with particularly strong equivalence in safety-style contexts ($d=0.21$ for both models).
+Politeness exhibits higher variability ($d$ reaching 0.52–0.63 in genre-shift and safety contexts) but remains within the small-effect range.
+Humor demonstrates intermediate stability (mean $d=0.39$), with genre-shift producing notably lower effects ($d=0.24$ for Qwen, $d=0.30$ for Llama).
+Figure 4 visualizes these patterns through environment-dependent effect sizes (left panels) and cross-environment transfer matrices (right panels), showing that equivalence remains stable even when vectors are extracted in one environment and tested in another.
+Implications.
+The persistence of small effect sizes across diverse prompt distributions demonstrates that non-identifiability is not an artifact of in-distribution prompts but a fundamental geometric property.
+Critically, this shows that naive distributional diversity without explicit invariance constraints or structural assumptions is insufficient to recover identifiable representations.
 7 Conclusion
-Rather than viewing non-identifiability as a limitation, we frame it as a necessary foundation for principled steering research.
-Our characterization clarifies what behavioral observations can reveal about representations and provides a framework for designing interventions that are both empirically effective and theoretically grounded.
-The empirical validation demonstrates that these are not abstract concerns: non-identifiability affects real steering interventions in deployed models, with typical steering vectors containing substantial observationally irrelevant components.
-However, the constructive characterization in Proposition 2 shows that identifiability is achievable through principled design choices, enabling alignment methods that clearly specify their affordances and limitations.
+We have established that steering vectors are fundamentally non-identifiable under standard observational regimes, with infinitely many distinct directions producing equivalent behavioral effects.
+This non-identifiability persists across models, traits, and prompt distributions, revealing a fundamental tension: methods optimized for behavioral control may not yield interpretable representations.
+Our work clarifies when steering provides reliable heuristic modification versus principled causal understanding, highlighting that interpretability claims require validation beyond input-output equivalence alone.
 8 Limitations
 Our empirical validation covers two models at mid-network layers across three semantic traits (formality, politeness, humor).
-While our theoretical results apply generally, the empirical magnitude of non-identifiability, including null space dimensionality and the fraction of vector norm therein, may vary across model families, scales, architectures and layer positions.
+While our theoretical results apply generally, the empirical magnitude of non-identifiability may vary across model families, scales, architectures and layer positions.
 Our semantic evaluation uses lexical heuristics rather than full distributional metrics or human judgments, providing a conservative test of observational equivalence.
-The primary theoretical mechanism relies on local linear approximation, formally valid in a neighborhood of the reference distribution, though our empirical validation confirms large equivalence classes exist in practice.
-Proposition 2 characterizes sufficient conditions for identifiability but comprehensive empirical evaluation of each regime across multiple models, tasks and experimental designs remains future work.
-The practical applicability of identifiable regimes involves trade-offs: ICA requires carefully curated contrastive prompts, sparsity regularization may reduce efficacy if true factors are not sparse and multi-environment validation requires substantial additional data.
-Understanding these trade-offs and extending our framework to other intervention modalities (activation patching, prompt based steering) are important directions for the representational alignment community.
+Our multi-environment experiments demonstrate that equivalence persists across diverse prompt distributions, but we do not test whether explicit structural interventions such as independence constraints (ICA), sparsity regularization, or invariance objectives (IRM) can recover identifiable representations.
+While theoretical frameworks suggest these approaches might enable identifiability, empirical validation across multiple models and experimental designs remains an important direction for future work.
+Additionally, we do not directly measure null-space dimensionality or test whether null-space versus row-space components exhibit differential robustness under distribution shift.
 9 Future Work
-Systematic evaluation of identifiable regimes, comparing ICA-based extraction, sparse recovery, multi-environment training and cross-layer validation against standard contrastive methods, would clarify which structural assumptions are most effective in practice.
-Hybrid methods combining multiple identifiability conditions may achieve stronger identifiability than individual approaches, while layer-wise analysis could reveal whether certain network positions afford more identifiable steering and decomposing the Jacobian by architectural component could identify which elements contribute to non-identifiability.
-The multi-environment identifiability condition connects to causal representation learning, where adapting invariant risk minimization to train steering vectors that maintain consistent effects across distributions could filter spurious correlations.
-Investigating adversarial robustness and out-of-distribution generalization would test whether identifiable methods exhibit superior robustness.
-Characterizing scaling laws for identifiability, including how null-space dimensionality evolves as models scale, would inform long-term strategy for controlling future systems.
-Acknowledgments
+Our results establish the necessity of validation beyond input-output behavior.
+Important directions include testing whether structural constraints can recover identifiable representations, analyzing how non-identifiability varies across network layers and architectural components, and characterizing how null-space dimensionality scales with model size.
+Investigating whether alternative steering methods exhibit superior robustness under adversarial attacks or distribution shift would clarify the practical consequences of non-identifiability.
+Human evaluation would provide stronger validation than our lexical heuristics, and direct measurement of null-space dimensionality would quantify our theoretical predictions.
+10 Acknowledgments
 We thank Vast.ai for providing GPU compute resources that enabled our empirical validation experiments.
 We also acknowledge the HuggingFace platform and maintainers of the open-source models (Qwen2.5-3B-Instruct, Llama-3.1-8B-Instruct) used in this work.
-Appendix A Prompt Construction Details
-A.1 Contrastive Prompt Pairs
-We construct contrastive prompt pairs for each semantic trait using template-based generation. Figure 3 illustrates the general structure and Table 2 provides concrete examples.
-Topic Pool 50 topics e.g., “weather”, “tech trends” Positive Template “formal”, “polite” Negative Template “informal”, “rude” 50 Contrastive Prompt Pairs
-Figure 3: Contrastive prompt pair construction.
-Trait Positive Prompt ($x^{+}$) Negative Prompt ($x^{-}$)
-Formality Write a professional and formal message about technology trends. Write a casual and informal message about technology trends.
-Politeness Write a polite and courteous response for a disagreement with someone. Write a rude and disrespectful response for a disagreement with someone.
-Humor Write a humorous and funny response about an awkward moment. Write a serious and straightforward response about an awkward moment.
+Appendix A Logit-Level Equivalence Test
+To complement our semantic evaluations, we perform a direct logit-level equivalence test.
+For each model and trait, we compare the next-token logit vectors produced by steering with $v$ and with $v+v_{\perp}$, where $v_{\perp}$ is a norm-matched direction orthogonal to $v$.
+We measure the $\ell_{2}$ distance between logit vectors and normalize by the corresponding distance induced by a norm-matched random direction:
+$R=\frac{\|\ell(v)-\ell(v+v_{\perp})\|_{2}}{\|\ell(v)-\ell(v_{\text{rand}})\|_{2}}$ (10)
+Values $R<1$ indicate that orthogonal perturbations induce smaller logit deviations than random directions.
+In all tested models and traits, this condition holds.
+Ratios range from $0.49$ (Llama, formality) to $0.90$ (Qwen, humor), indicating that orthogonal perturbations systematically induce smaller logit shifts than random directions.
+The next-token predictions remain highly stable: token agreement between $v$ and $v+v_{\perp}$ ranges from $81\%$ to $96\%$, and top-10 token overlap ranges from $86\%$ to $96\%$.
+These results provide direct evidence that equivalence is not an artifact of semantic measurement.
+While orthogonal perturbations induce some logit deviation, they systematically preserve output distributions more faithfully than random directions, consistent with Proposition 1’s prediction of high-dimensional equivalence classes.
+Table 3 reports mean $\ell_{2}$ logit distances, normalized ratios, token agreement, and top-$k$ overlap for all models and traits.
+Table 3: Logit-level equivalence statistics.
+Model Trait $R$ (v_perp / random) Token Agreement Top-10 Overlap
+Qwen-3B Humor 0.90 81% 86%
+Qwen-3B Formality 0.76 94% 96%
+Qwen-3B Politeness 0.75 93% 96%
+Llama-8B Humor 0.64 96% 90%
+Llama-8B Formality 0.49 93% 95%
+Llama-8B Politeness 0.65 91% 94%
+Appendix B Prompt Construction Details
+B.1 Contrastive Prompt Pairs
+We construct contrastive prompt pairs for each semantic trait using template-based generation. Table 4 provides concrete examples for all three traits used in the empirical validation.
+Trait Positive ($x^{+}$) Negative ($x^{-}$)
+Formality Draft a professional email summarizing quarterly performance. Text a friend about how the quarter went.
+Politeness Respond diplomatically to criticism from a colleague. Reply bluntly to someone who criticized you.
+Humor Tell a lighthearted story about an embarrassing mistake. Provide a factual account of an embarrassing mistake.
 Table 4: Example contrastive prompt pairs for each semantic trait.
-Each pair consists of a positive prompt ($x^{+}$) designed to elicit high trait values and a negative prompt ($x^{-}$) designed to elicit low trait values.
-A.2 Steering Extraction and Evaluation Pipeline
+B.2 Steering Extraction and Evaluation Pipeline
+Figure 5: Steering extraction and evaluation pipeline.
 Figure 5 illustrates the pipeline used to extract and evaluate steering vectors.
@@
 This vector is then applied to a separate set of held-out evaluation prompts to generate steered outputs.
-Steering Extraction 50 contrastive pairs (trait-specific templates) Steering Vector $v\in\mathbb{R}^{d}$ Evaluation 100 held-out prompts (diverse, neutral) Steered Output $o(x,v,\alpha)$ extractapplygenerate Key: Test prompts are intentionally orthogonal to extraction prompts to ensure steering generalizes beyond training distribution.
-Figure 4: Steering extraction and evaluation pipeline. Steering vectors are extracted from trait-specific contrastive pairs, then evaluated on diverse held-out prompts to test generalization.
-Appendix B Detailed Proof of Proposition 1
-B.1 Statement and Overview
-Proposition 1.
-Under Assumptions A1–A3, in Regime 2 (white-box single-layer access) without additional structural constraints, persona vectors are not identifiable.
-Specifically, for any steering vector $v\in\mathbb{R}^{d}$, there exist infinitely many vectors $v^{\prime}\not\propto v$ that are observationally equivalent.
-Proof strategy.
-We establish non-identifiability through two complementary mechanisms:
-  • Null-space ambiguity (primary, constructive).
-  • Reparameterization symmetry (existence-based).
-B.2 Null-Space Ambiguity (Constructive Proof)
+Appendix C Detailed Proof of Proposition 1
+We provide proofs for establishing non-identifiability via two complementary arguments: null-space ambiguity and reparameterization symmetry.
+C.1 Null-Space Ambiguity (Constructive Proof)
 Setup. Consider the local linear approximation of the steering effect:
@@
 Then for all $x$ and all $\alpha$:
-$J_{\ell}(x)v^{\prime}$ $=J_{\ell}(x)(v+v_{0})$ (8)
-$=J_{\ell}(x)v+J_{\ell}(x)v_{0}$
-$=J_{\ell}(x)v.$
+$J_{\ell}(x)v^{\prime}=J_{\ell}(x)(v+v_{0})=J_{\ell}(x)v+J_{\ell}(x)v_{0}=J_{\ell}(x)v.$ (17)
 Therefore:
-$o(x,v^{\prime},\alpha)$ $\approx o(x,0,0)+\alpha J_{\ell}(x)v^{\prime}$ (9)
-$=o(x,0,0)+\alpha J_{\ell}(x)v$
-$\approx o(x,v,\alpha).$
+$o(x,v^{\prime},\alpha)\approx o(x,0,0)+\alpha J_{\ell}(x)v^{\prime}=o(x,0,0)+\alpha J_{\ell}(x)v\approx o(x,v,\alpha).$ (18)
 Step 5: Infinitely many distinct solutions.
@@
 Under the linear approximation and for typical Jacobian structure, there exist infinitely many geometrically distinct steering vectors that are observationally equivalent.
-B.3 Reparameterization Symmetry (Existence Proof)
+C.2 Reparameterization Symmetry (Existence Proof)
 Setup: Neural networks exhibit inherent symmetries arising from overparameterization. We show these symmetries induce non-identifiability even in the exact nonlinear case.
@@
 Then:
-$h^{\prime}_{\ell+1}$ $=\sigma\!\left(W_{\ell+1}A^{-1}Ah_{\ell}+b_{\ell+1}\right)$ (16)
-$=\sigma\!\left(W_{\ell+1}h_{\ell}+b_{\ell+1}\right)$
-$=h_{\ell+1}.$
+$h^{\prime}_{\ell+1}=\sigma\!\left(W_{\ell+1}A^{-1}Ah_{\ell}+b_{\ell+1}\right)=\sigma\!\left(W_{\ell+1}h_{\ell}+b_{\ell+1}\right)=h_{\ell+1}.$ (25)
 Step 3: Steering under reparameterization.
@@
 Reparameterized steering with $v^{\prime}=Av$:
-$\tilde{h}^{\prime}_{\ell}$ $=h^{\prime}_{\ell}+\alpha v^{\prime}$ (19)
-$=Ah_{\ell}+\alpha Av$
-$=A(h_{\ell}+\alpha v)$
-$=A\tilde{h}_{\ell}$,
-$\tilde{h}^{\prime}_{\ell+1}$ $=\sigma\!\left(W^{\prime}_{\ell+1}(h^{\prime}_{\ell}+\alpha v^{\prime})+b^{\prime}_{\ell+1}\right)$
+$\tilde{h}^{\prime}_{\ell}=h^{\prime}_{\ell}+\alpha v^{\prime}=Ah_{\ell}+\alpha Av=A(h_{\ell}+\alpha v)=A\tilde{h}_{\ell}$. (28)
+$\tilde{h}^{\prime}_{\ell+1}$ $=\sigma\!\left(W^{\prime}_{\ell+1}(h^{\prime}_{\ell}+\alpha v^{\prime})+b^{\prime}_{\ell+1}\right)$ (29)
 $=\sigma\!\left(W_{\ell+1}A^{-1}A(h_{\ell}+\alpha v)+b_{\ell+1}\right)$
@@
 The inherent symmetries of neural network parameterizations induce infinitely many observationally equivalent steering vectors.
-Appendix C Detailed Proof of Proposition 2
-C.1 Statement and Overview
-Proposition 2.
-Persona vectors can be identified up to scaling and permutation, thus affording reliable alignment control, under the following sufficient structural conditions:
-  • Statistical Independence (ICA). Latent components are statistically independent, allowing unique recovery via independent component analysis.
-  • Sparsity constraints. Steering directions admit sparse representations that reduce null-space ambiguity.
-  • Multi-environment or interventional data. Variation across environments or interventions breaks observational equivalences.
-  • Cross-layer consistency. Valid semantic directions propagate coherently across layers, filtering spurious components.
-We provide detailed proofs for each condition.
-C.2 Proof of Condition: Statistical Independence (ICA)
-Setup: Assume the latent persona is a vector $z=(z_{1},\ldots,z_{k})\in\mathbb{R}^{k}$ with independent components and the representation follows:
-$h_{\ell}=Az+\epsilon$ (20)
-where $A\in\mathbb{R}^{d\times k}$ is a mixing matrix and $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$ is Gaussian noise.
-Goal: Show that $A$ (and hence its columns, which are the steering vectors) can be recovered up to permutation and scaling.
-Step 1: ICA identifiability theorem.
-Let $x=As$ where $s\in\mathbb{R}^{k}$ has independent components and $A\in\mathbb{R}^{d\times k}$ is full column rank. Under the following sufficient conditions (Comon, 1994; Hyvärinen and Oja, 2000):
-  • At most one component of $s$ is Gaussian
-  • Components are statistically independent
-  • Sufficient samples are observed
-Then $A$ is identifiable up to column permutation and column scaling.
-Important caveat: These are strong assumptions that may not hold exactly in practice for LLM personas.
-Statistical independence between semantic factors (e.g., formality and politeness) is an idealization; real personas may exhibit weak dependencies.
-The ICA framework provides a sufficient structural condition for identifiability, not a characterization of what typically holds in contemporary steering pipelines.
-Step 2: Application to steering.
-In our setting:
-  • Observations: $h_{\ell}(x_{1}),\ldots,h_{\ell}(x_{N})$
-  • Model: $h_{\ell}(x_{i})=Az(x_{i})+\epsilon_{i}$
-  • Sources: $z(x_{1}),\ldots,z(x_{N})$ are realizations of independent persona components
-Assumption verification:
-  • Independence: We assume $z_{1},\ldots,z_{k}$ are statistically independent
-  • Non-Gaussianity: At most one $z_{i}$ is Gaussian (e.g., politeness and formality are typically non-Gaussian in natural language)
-  • Full rank: $A$ has full column rank (each persona has a non-zero effect on representations)
-Step 3: Recovery via ICA algorithm.
-Apply ICA algorithm (Hyvärinen and Oja, 2000) to observations $\{h_{\ell}(x_{i})\}_{i=1}^{N}$:
-  • Whitening: Compute $\tilde{h}_{\ell}=\Sigma^{-1/2}(h_{\ell}-\mu)$ where $\mu=\mathbb{E}[h_{\ell}]$ and $\Sigma=\operatorname{Cov}(h_{\ell})$.
-  • Independent component extraction: Find unmixing matrix $W$ that maximizes non-Gaussianity of $\hat{z}=W\tilde{h}_{\ell}$.
-  • Recover mixing matrix: $\hat{A}=\Sigma^{1/2}W^{-1}$.
-Step 4: Identifiability guarantee.
-By the ICA theorem, $\hat{A}=APD$ where:
-  • $P\in\mathbb{R}^{k\times k}$ is a permutation matrix
-  • $D\in\mathbb{R}^{k\times k}$ is a diagonal scaling matrix
-Therefore, the columns of $\hat{A}$ are the columns of $A$ up to permutation and scaling.
-Since steering vectors are defined as $v_{i}=Ae_{i}$ (the $i$-th column of $A$), we recover the true steering directions uniquely (up to unavoidable symmetries).
-Step 5: Noise robustness.
-With Gaussian noise $\epsilon\sim\mathcal{N}(0,\sigma^{2}I)$, ICA remains consistent:
-$h_{\ell}=Az+\epsilon$ (21)
-Since $\epsilon$ is Gaussian and $z$ is non-Gaussian (by assumption), the ICA algorithm separates signal from noise:
-  • Recovered sources: $\hat{z}=z+\tilde{\epsilon}$ where $\tilde{\epsilon}$ is small for $\sigma^{2}\ll\|Az\|^{2}$
-  • Recovered mixing: $\hat{A}\approx APD$ with estimation error decreasing as $N\to\infty$
-Conclusion (ICA): Under statistical independence, non-Gaussianity and full column rank—strong sufficient conditions that may not hold automatically in practice—persona vectors are identifiable up to permutation and scaling.
-These assumptions provide a theoretical pathway to identifiability but require careful experimental design to approximate in real steering settings.
-C.3 Proof of Condition: Sparsity Constraints
-Setup: Assume the true persona vector $v\in\mathbb{R}^{d}$ is sparse with $\|v\|_{0}=s\ll d$, meaning at most $s$ entries are non-zero. We observe the effect of steering:
-$y=J_{\ell}v+\eta$ (22)
-where $y\in\mathbb{R}^{m}$ are output measurements, $J_{\ell}\in\mathbb{R}^{m\times d}$ is the measurement matrix (Jacobian) and $\eta\sim\mathcal{N}(0,\sigma^{2}I)$ is noise.
-Goal: Show that $v$ can be uniquely recovered via $\ell_{1}$ minimization.
-Step 1: Compressed sensing setup.
-The recovery problem is:
-$\min_{v^{\prime}\in\mathbb{R}^{d}}\|v^{\prime}\|_{1}\quad\text{subject to}\quad\|J_{\ell}v^{\prime}-y\|_{2}\leq\epsilon$ (23)
-where $\epsilon$ bounds the noise: $\epsilon\geq\|\eta\|_{2}$ with high probability.
-Step 2: Restricted Isometry Property (RIP).
-A matrix $J_{\ell}$ satisfies the RIP of order $s$ with constant $\delta_{s}$ if for all $s$-sparse vectors $v$:
-$(1-\delta_{s})\|v\|_{2}^{2}\leq\|J_{\ell}v\|_{2}^{2}\leq(1+\delta_{s})\|v\|_{2}^{2}$ (24)
-Theorem (Candes and Tao, 2005): If $J_{\ell}$ satisfies RIP with $\delta_{2s}<\sqrt{2}-1\approx 0.414$, then $\ell_{1}$ minimization recovers $v$ exactly (in the noiseless case) or approximately (with noise) with error:
-$\|v-\hat{v}\|_{2}\leq C\epsilon$ (25)
-for some constant $C$ depending on $\delta_{2s}$.
-Important caveat: The RIP condition is a strong assumption.
-For Jacobians $J_{\ell}$ arising from neural network steering, RIP typically does not hold automatically and must be verified empirically or enforced through measurement design (diverse prompt selection).
-This condition is sufficient for identifiability but may not be satisfied in standard steering settings without careful experimental design.
-Step 3: Recovery guarantee.
-Solve:
-$\hat{v}=\arg\min_{v^{\prime}}\|v^{\prime}\|_{1}\quad\text{s.t.}\quad\|J_{\ell}v^{\prime}-y\|_{2}\leq\epsilon$ (26)
-By the compressed sensing theorem:
-$\|v-\hat{v}\|_{2}\leq C_{1}\epsilon+C_{2}\frac{\|v-v_{s}\|_{1}}{\sqrt{s}}$ (27)
-where $v_{s}$ is the best $s$-sparse approximation to $v$. If $v$ is exactly $s$-sparse, the second term vanishes and:
-$\|v-\hat{v}\|_{2}\leq C_{1}\epsilon$ (28)
-Step 4: Uniqueness.
-Suppose two sparse vectors $v$ and $v^{\prime}$ both satisfy $\|J_{\ell}v-y\|_{2}\leq\epsilon$. Then:
-$\|J_{\ell}(v-v^{\prime})\|_{2}\leq 2\epsilon$ (29)
-If $v-v^{\prime}$ is $2s$-sparse and $J_{\ell}$ satisfies RIP, then by the RIP condition:
-$(1-\delta_{2s})\|v-v^{\prime}\|_{2}^{2}\leq\|J_{\ell}(v-v^{\prime})\|_{2}^{2}\leq 4\epsilon^{2}$ (30)
-Therefore:
-$\|v-v^{\prime}\|_{2}\leq\frac{2\epsilon}{\sqrt{1-\delta_{2s}}}$ (31)
-As $\epsilon\to 0$ (noiseless case), $v=v^{\prime}$, establishing uniqueness.
-Conclusion (Sparsity): Under sparsity assumptions and RIP conditions—strong sufficient conditions that typically require careful measurement design—persona vectors are uniquely recoverable via $\ell_{1}$ minimization.
-RIP does not hold automatically for arbitrary Jacobians and must be verified or engineered through diverse prompt selection.
-C.4 Proof of Condition: Multi-Environment Identification
-Setup: Assume we observe the same persona $z$ across multiple environments $e\in\{1,\ldots,E\}$ where:
-$h_{\ell}^{(e)}=g_{e}(x,z)+\epsilon^{(e)}$ (32)
-The function $g_{e}$ may vary across environments (spurious correlations change) but the causal effect of $z$ on downstream outputs remains invariant.
-Goal: Show that the invariant representation $z$ (and its associated direction) can be identified.
-Step 1: Invariant causal mechanism.
-Assume the causal graph:
-[…]
-where:
-  • $z$: true persona (causal factor),
-  • $x$: input prompt,
-  • $e$: environment,
-  • $h_{\ell}$: internal representation,
-  • $o$: output.
-The key assumption is invariance: the causal mechanism $h_{\ell}\to o$ is the same across environments, i.e., $F_{\ell\to L}$ does not depend on $e$.
-Step 2: Invariant Risk Minimization (IRM).
-The objective is to find representation $\Phi:\mathbb{R}^{d}\to\mathbb{R}^{k}$ and predictor $w:\mathbb{R}^{k}\to\mathbb{R}$ such that:
-$\min_{\Phi,w}$ $\sum_{e=1}^{E}R^{e}(\Phi,w)$ (33)
-s.t. $w\in\arg\min_{w^{\prime}}R^{e}(\Phi,w^{\prime})\quad\forall e$
-where,
-$R^{e}(\Phi,w)=\mathbb{E}_{(x,y)\sim\mathcal{P}^{e}}\left[\ell\big(w\cdot\Phi(h_{\ell}(x)),y\big)\right].$ (34)
-Interpretation: Find a representation $\Phi(h_{\ell})=z$ such that the optimal predictor $w$ is the same across all environments. This filters out spurious correlations that vary with $e$.
-Step 3: Identifiability under IRM.
-Under the following idealized conditions:
-  • Environments are sufficiently diverse: $\text{Cov}(x|e)$ varies across $e$
-  • Causal mechanism is invariant: $p(o|z)$ is the same for all $e$
-  • Sufficient environments: $E\geq k+1$ where $k=\dim(z)$
-  • No unobserved confounders between environments and outcomes
-The theorem (Ahuja et al., 2022; Von Kügelgen et al., 2021) states that the invariant risk minimization objective can recover $z$ up to an invertible transformation, though not necessarily uniquely.
-Important caveat: The literature on invariant representation learning establishes conditions under which invariance constraints narrow the hypothesis class but full identifiability (unique recovery of $z$) requires additional assumptions beyond standard IRM formulations.
-The practical application of IRM to steering should be understood as identifying a stable, transferable representation rather than guaranteeing uniqueness.
-This is a sufficient condition in principle under strong assumptions, not a characterization of typical steering scenarios.
-Step 4: Application to steering.
-In practice:
-  1. 1.
-    Collect multi-environment data: Extract steering vectors from diverse prompt distributions, model checkpoints or instruction formats.
-  2. 2.
-    Learn invariant representation: Train $\Phi$ via IRM objective
-  3. 3.
-    Extract steering vectors: Compute $v_{i}=\nabla_{h_{\ell}}\Phi_{i}(h_{\ell})$ where $\Phi_{i}$ is the $i$-th component of $\Phi$
-The resulting steering vectors $v_{i}$ capture invariant causal factors rather than spurious correlations.
-Step 5: Theoretical guarantee.
-Under mild conditions (sufficient diversity, invariance, identifiable causal structure), the IRM solution satisfies:
-$\Phi(h_{\ell})=T(z)$ (35)
-where $T$ is an invertible transformation. Since persona steering operates on $z$, the directions $v_{i}=\nabla_{h_{\ell}}\Phi_{i}$ recover the true causal factors.
-Conclusion (Multi-environment): With diverse environments and invariant causal mechanisms—idealized sufficient conditions that narrow the hypothesis class—persona vectors corresponding to stable causal factors can be recovered up to invertible transformations.
-Full uniqueness requires additional assumptions beyond standard IRM formulations.
-This provides a principled approach to filtering spurious correlations but should be understood as identifying transferable representations rather than guaranteeing unique recovery.
-C.5 Proof of Condition: Cross-Layer Consistency
-Setup: Assume persona vectors exhibit consistent geometric relationships across layers:
-$v_{\ell+1}=W_{\ell}v_{\ell}+\delta_{\ell}$ (36)
-where $W_{\ell}\in\mathbb{R}^{d\times d}$ is the weight matrix connecting layers and $\|\delta_{\ell}\|$ is small.
-Goal: Show that cross-layer constraints reduce the solution space and improve identifiability.
-Step 1: Single-layer null space.
-From proposition 1, at layer $\ell$:
-$\mathcal{N}_{\ell}=\{v_{0}:J_{\ell}v_{0}=0\}$ (37)
-with $\dim(\mathcal{N}_{\ell})=d-r_{\ell}$ where $r_{\ell}=\text{rank}(J_{\ell})$.
-Step 2: Cross-layer propagation.
-If $v_{\ell}\in\mathcal{N}_{\ell}$, does $v_{\ell+1}=W_{\ell}v_{\ell}\in\mathcal{N}_{\ell+1}$?
-Generally, no. The null space changes across layers:
-$v_{\ell}\in\mathcal{N}_{\ell}\not\Rightarrow W_{\ell}v_{\ell}\in\mathcal{N}_{\ell+1}$ (38)
-Step 3: Intersection of constraints.
-Consider steering vectors observed at multiple layers $\ell_{1},\ldots,\ell_{L}$. Each layer provides a constraint:
-$J_{\ell_{i}}v_{\ell_{i}}=y_{i}$ (39)
-If we additionally require consistency:
-$v_{\ell_{i+1}}=W_{\ell_{i}}v_{\ell_{i}}+\delta_{\ell_{i}}$ (40)
-Then the solution must satisfy:
-$v_{\ell_{i}}\in\{v:J_{\ell_{i}}v=y_{i}\}\cap\{v:W_{\ell_{i}}v\approx v_{\ell_{i+1}}\}$ (41)
-Step 4: Reduced null space (qualitative characterization).
-The cross-layer constraints create an overdetermined system. The effective null space is:
-$\mathcal{N}_{\text{eff}}=\bigcap_{i=1}^{L-1}\{v:W_{\ell_{i}}v\in\mathcal{N}_{\ell_{i+1}}\}$ (42)
-Since each $W_{\ell_{i}}$ generically has full rank and maps null space vectors to non-null-space vectors, we expect:
-$\dim(\mathcal{N}_{\text{eff}})\ll\dim(\mathcal{N}_{\ell})$
-Intuition: Each additional layer imposes new independent constraints.
-If the propagation matrices $W_{\ell_{i}}$ are sufficiently "generic" (full rank with uncorrelated null space mappings), then each layer reduces the effective null-space dimension.
-For sufficiently many informative layers with uncorrelated constraint structures, the effective null space can shrink dramatically or even vanish.
-Conclusion (Cross-layer): Cross-layer consistency constraints can substantially reduce null-space dimensionality by creating overdetermined systems.
-The extent of reduction depends on the specific geometric structure of propagation matrices and layer-wise Jacobians.
-While this approach does not guarantee complete identifiability in general, it provides a practical method for filtering spurious null-space components that lack geometric stability across layers.
 Appendix D Intuitive Explanations
 D.1 Geometric Intuition for Null-Space Ambiguity
+[…]
+Figure 6: Geometric intuition for null-space ambiguity.
+The output observes only the $(v_{1},v_{2})$ components (blue shaded region).
+The $v_{3}$ component lies in $\ker(J)$ and is invisible to the output.
+Adding any $\alpha v_{3}$ to $v$ leaves the output unchanged: $J(v+\alpha v_{3})=Jv$ for all $\alpha\in\mathbb{R}$.
 Visual analogy:
@@
 Figure 6 illustrates this geometric intuition.
-[…]
-Figure 5: Geometric intuition for null-space ambiguity.
-The output observes only the $(v_{1},v_{2})$ components (blue shaded region).
-The $v_{3}$ component lies in $\ker(J)$ and is invisible to the output.
-Adding any $\alpha v_{3}$ to $v$ leaves the output unchanged: $J(v+\alpha v_{3})=Jv$ for all $\alpha\in\mathbb{R}$.
 Any vector of the form $v^{\prime}=v+\alpha v_{3}$ for $\alpha\in\mathbb{R}$ produces identical outputs:
-$J(v+\alpha v_{3})$ $=Jv+\alpha J(0,0,1)^{\top}$ (43)
-$=Jv+\alpha\cdot 0$
-$=Jv$
+$J(v+\alpha v_{3})=Jv+\alpha J(0,0,1)^{\top}=Jv+\alpha\cdot 0=Jv.$ (30)
 Since $\alpha$ can take infinitely many values and $v^{\prime}\not\propto v$ for generic choices of $\alpha$, there exist infinitely many geometrically distinct steering vectors that are observationally equivalent.
@@
 $\ker(J_{\text{stack}})=\bigcap_{i=1}^{N}\ker(J_{\ell}(x_{i}))$ (33)
+[…]
+Figure 7: Illustration of why prompt diversity does not resolve null-space ambiguity.
+Each prompt induces a Jacobian that measures the same effective subspace, leaving the null-space intersection unchanged.
 Critical observation: If all $J_{\ell}(x_{i})$ share a common null space (e.g., all prompts probe similar aspects of the model), then:
@@
 This means adding more prompts does not reduce the null space when all prompts probe the same effective subspace—the ambiguity persists. Figure 7 illustrates this phenomenon.
-[…]
-Figure 6: Illustration of why prompt diversity does not resolve null-space ambiguity.
-Each prompt induces a Jacobian that measures the same effective subspace, leaving the null-space intersection unchanged.
 D.3 Why Orthogonal Perturbations Preserve Semantics
@@
 Analogy: If $Jv$ is a strong wind blowing consistently north, then $Jv_{\perp,\text{row}}$ is turbulence—it may have comparable energy but lacks directional coherence, so the overall trajectory remains northward.
-D.4 ICA Intuition: Why Independence Helps
-Setup: Suppose representations encode two independent factors:
-  • $z_{1}$: formality (independent)
-  • $z_{2}$: politeness (independent)
-And mix them:
-$h=v_{1}z_{1}+v_{2}z_{2}$ (51)
-Problem without independence.
-If $z_{1}$ and $z_{2}$ are correlated (e.g., formal text tends to be polite), we can reparameterize:
-$\tilde{z}_{1}=z_{1}+\alpha z_{2},\quad\tilde{z}_{2}=z_{2}$ (52)
-$h=(v_{1}-\alpha v_{2})\tilde{z}_{1}+(v_{2}+\alpha v_{1})\tilde{z}_{2}$ (53)
-giving infinitely many equivalent decompositions (different $\alpha$ values).
-Solution with independence.
-If $z_{1}\perp z_{2}$ (statistically independent), then $\tilde{z}_{1}=z_{1}+\alpha z_{2}$ is NOT independent of $\tilde{z}_{2}=z_{2}$ for $\alpha\neq 0$:
-$I(\tilde{z}_{1};\tilde{z}_{2})=I(z_{1}+\alpha z_{2};z_{2})>0$ (54)
-ICA finds the unique decomposition where sources are independent (refer Figure 7), breaking the symmetry.
-$z_{1}$$z_{2}$Many rotations OK
-(a) Correlated sources
-$z_{1}$$z_{2}$Only THIS rotation
-(b) Independent sources
-Figure 7: Visual comparison of correlated versus independent sources.
-D.5 Sparsity Intuition: Occam’s Razor
-Setup: Suppose we observe:
-$y=Jv+\eta$ (55)
-and both $v_{1}$ and $v_{2}$ fit the data:
-$\|Jv_{1}-y\|\approx\|Jv_{2}-y\|\approx\epsilon$ (56)
-Question: Which is the “true” vector?
-Sparsity principle.
-Prefer the simplest explanation. If:
-  • $v_{1}$ has 100 non-zero entries
-  • $v_{2}$ has 10 non-zero entries
-then $v_{2}$ is more likely to be the true sparse signal.
-Why this works.
-If the true vector is sparse, then dense solutions must include spurious noise components:
-$v_{\text{dense}}=v_{\text{true}}+v_{\text{noise}}$
-The $\ell_{1}$ penalty penalizes density, filtering out $v_{\text{noise}}$.
-D.6 Multi-Environment Intuition: Finding What’s Stable
-Setup: Observe “formality” steering in 3 environments:
-  • E1 (academic): Formality correlates with technical jargon
-  • E2 (business): Formality correlates with professional tone
-  • E3 (legal): Formality correlates with precise language
-Spurious correlation problem.
-Single-environment extraction might learn:
-$v_{1}$ $=v_{\text{formality}}+v_{\text{jargon}}$
-$v_{2}$ $=v_{\text{formality}}+v_{\text{professional}}$
-$v_{3}$ $=v_{\text{formality}}+v_{\text{precise}}$
-Each includes the true factor plus environment-specific confounds.
-Invariance solution.
-Find the component present in all environments:
-$v_{\text{invariant}}=\text{stable component of }\{v_{1},v_{2},v_{3}\}$ (57)
-The spurious parts ($v_{\text{jargon}},v_{\text{professional}},v_{\text{precise}}$) vary across environments and get filtered out.
-E1: formality + jargonE2: formality + professionalE3: formality + preciseIntersectionformality (stable core)
-Figure 8: Multi-environment filtering extracts the stable core.
-Each environment contains the true semantic factor (formality) plus environment-specific confounds.
-The intersection across environments isolates the invariant component.
-D.7 Cross-Layer Intuition: Consistency Check
-Setup: Extract steering vectors at layers 8, 12, 16, 20.
-Consistency hypothesis.
-True semantic factors should propagate coherently:
-$v_{12}\approx W_{8}v_{8},\quad v_{16}\approx W_{12}v_{12},\quad\text{etc.}$
-Inconsistency indicates null-space artifacts.
-If $v_{8}$ includes a large null-space component $v_{\perp,8}$:
-$v_{8}=v_{\text{true},8}+v_{\perp,8}$ (58)
-Then after propagation:
-$W_{8}v_{8}=W_{8}v_{\text{true},8}+W_{8}v_{\perp,8}$ (59)
-Since $v_{\perp,8}\in\ker(J_{8})$ but generically $W_{8}v_{\perp,8}\notin\ker(J_{12})$, the null-space component becomes observable at layer 12, causing inconsistency.
-[…]
-Figure 9: Cross-layer consistency check. True semantic factors propagate coherently, while null-space artifacts become inconsistent.
-Filtering strategy.
-Project onto the subspace of vectors that are consistent across layers:
-$v_{\text{consistent}}=\arg\min_{v}\sum_{\ell}\|W_{\ell-1}v_{\ell-1}-v_{\ell}\|^{2}$ (60)
-This filters out layer-specific artifacts (See Figure 9), retaining only the stable semantic direction.
 Appendix E Mathematical Derivations
@@
 Fisher Information Matrix: For the linear Gaussian model $o=Jv+\eta$ with $\eta\sim\mathcal{N}(0,\sigma^{2}I)$:
-$\mathcal{I}(v)$ $=\mathbb{E}_{o}\Bigg[\left(\frac{\partial\log p(o\mid x,v,\alpha)}{\partial v}\right)$ (65)
-$\qquad\quad\left(\frac{\partial\log p(o\mid x,v,\alpha)}{\partial v}\right)^{\!\top}\Bigg]$
-$=\frac{1}{\sigma^{2}}J^{\top}J.$
+$\mathcal{I}(v)=\mathbb{E}_{o}\left[\left(\frac{\partial\log p(o\mid x,v,\alpha)}{\partial v}\right)\left(\frac{\partial\log p(o\mid x,v,\alpha)}{\partial v}\right)^{\!\top}\right]=\frac{1}{\sigma^{2}}J^{\top}J.$ (42)
 Cramér-Rao Lower Bound: The covariance of any unbiased estimator $\hat{v}$ satisfies:
@@
 More data cannot help because the information geometry has infinite uncertainty in null-space directions.
-E.3 Proof that ICA Breaks Gauge Symmetry
-Setup: Consider two decompositions:
-$h=A_{1}z_{1}=A_{2}z_{2}$ (69)
-where both $z_{1}$ and $z_{2}$ have independent components.
-Question: When are these equivalent ($A_{1}=A_{2}$ up to permutation/scaling)?
-Analysis: If both decompositions are valid:
-$A_{1}z_{1}=A_{2}z_{2}\implies z_{1}=A_{1}^{-1}A_{2}z_{2}=Gz_{2}$ (70)
-where $G=A_{1}^{-1}A_{2}$.
-ICA Constraint: If both $z_{1}$ and $z_{2}$ have independent components:
-$I(z_{1,i};z_{1,j})$ $=0$ $\text{for }i\neq j,$ (71)
-$I\big((Gz_{2})_{i};(Gz_{2})_{j}\big)$ $=0$ $\text{for }i\neq j.$
-Key Theorem (Comon, 1994): If $z_{2}$ has independent components and $Gz_{2}$ also has independent components, then $G$ must be a generalized permutation matrix:
-$G=PD$ (72)
-where $P$ is a permutation matrix and $D$ is a diagonal scaling matrix.
-Implication:
-$A_{1}z_{1}$ $=A_{2}z_{2},$ (73)
-$A_{1}Gz_{2}$ $=A_{2}z_{2},$
-$A_{1}PD$ $=A_{2}.$
-Therefore $A_{2}=A_{1}PD$, meaning $A_{2}$ is $A_{1}$ with permuted and scaled columns.
-Conclusion: ICA constraints (statistical independence) force $G$ to be a permutation-scaling matrix, breaking arbitrary gauge transformations and establishing identifiability up to unavoidable symmetries.
-Independence is the structural assumption that eliminates the infinite equivalence class.
 Appendix F Disclosure of the Use of AI

### v3

@@
 Activation steering methods are widely used to control large language model (LLM) behavior and are often interpreted as revealing meaningful internal representations.
-This interpretation assumes steering directions are identifiable and uniquely recoverable from input–output behavior.
+This interpretation assumes that steering directions are identifiable and uniquely recoverable from input–output behavior.
 We show that, under white-box single-layer access, steering vectors are fundamentally non-identifiable due to large equivalence classes of behaviorally indistinguishable interventions.
-Empirically, we show that orthogonal perturbations achieve near-equivalent efficacy with negligible effect sizes across multiple models and traits.
-Critically, we show that the non-identifiability is a robust geometric property that persists across diverse prompt distributions.
+Empirically, we find that orthogonal perturbations achieve near-equivalent efficacy with negligible effect sizes across multiple models and traits.
+Critically, we show that non-identifiability is a robust geometric property that persists across diverse prompt distributions.
 These findings reveal fundamental interpretability limits and highlight the need for structural constraints beyond behavioral testing to enable reliable alignment interventions.
-[^1] Code: https://github.com/sohv/non-identifiability
 On the Non-Identifiability of Steering Vectors in Large Language Models
-Sohan Venkateshthanks: Correspondence: soh.venkatesh@gmail.com     Ashish Mahendran Kurapath Manipal Institute of Technology Bengaluru
+Sohan Venkatesh1thanks: Correspondence: soh.venkatesh@gmail.com        Ashish Mahendran Kurapath1 1Manipal Institute of Technology Bengaluru
 1 Introduction
@@
 Despite growing adoption in interpretability and alignment research, the theoretical foundations of persona steering remain poorly understood.
-Most existing methods implicitly assume that extracted steering vectors correspond to meaningful, uniquely determined latent factors and that these factors can be directly manipulated to achieve reliable control.
+Most existing methods implicitly assume that extracted steering vectors correspond to meaningful and uniquely determined latent factors and that these factors can be directly manipulated to achieve reliable control.
 However, classical results in latent variable modeling and causal inference show that such assumptions are often unjustified without additional structural constraints (Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Schölkopf et al., 2021; Locatello et al., 2019).
@@
 Finally, identifiability theory clarifies which experimental protocols provide meaningful evidence for causal interpretation of interventions on internal activations (Schölkopf et al., 2021; Ahuja et al., 2022).
-Figure 1: Geometric illustration of steering vector non-identifiability.
-Two geometrically distinct vectors, the original steering vector $v$ (left) and a perturbed version $v+v_{\perp}$ where $v_{\perp}\in\mathrm{ker}(J)$ (right), produce identical observable outputs.
-The null-space component (red, unobservable region) can be arbitrarily modified without changing outputs, demonstrating the existence of large equivalence classes of behaviorally indistinguishable steering directions.
+Figure 1: Steering vector non-identifiability: The original vector $v$ and a perturbed vector $v+v_{\perp}$ with $v_{\perp}\in\mathrm{ker}(J)$ produce identical outputs since null-space components do not affect JvJv, the first-order change in output logits induced by steering.
 We make three contributions.
@@
 Second, we empirically demonstrate that extracted steering vectors contain substantial non-identifiable components, with orthogonal perturbations achieving 95–100% of the original steering efficacy across multiple models and traits.
-Third, we validate through multi-environment experiments that this equivalence persists under distribution shift, confirming non-identifiability is a robust geometric property rather than an artifact of limited prompt diversity.
+Third, we validate through multi-environment experiments that this equivalence persists under distribution shift, confirming non-identifiability is a robust geometric property rather than just an artifact of limited prompt diversity.
 These results clarify when alignment methods can be trusted and which structural assumptions are required for reliable control.
@@
 as a latent variable identification problem, bridging causal representation learning, activation editing in LLMs and mechanistic interpretability.
-Causal and latent variable identifiability.
+Causal and latent variable identifiability
 Classical results show that latent variable models are non-identifiable without structural assumptions (Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Kruskal, 1977).
 Recent work in causal representation learning extends these ideas to deep learning settings (Schölkopf et al., 2021; Ahuja et al., 2022; Locatello et al., 2019), establishing conditions under which latent factors can be recovered from high-dimensional observations.
-Probing and representation learning.
+Probing and representation learning
 Work on probing classifiers (Pimentel et al., 2020; Ravfogel et al., 2020; Elazar et al., 2021) shows that linear directions can succeed for reasons unrelated to target information.
 The linear representation hypothesis (Park et al., 2023; Elhage et al., 2022) suggests concepts correspond to directions in activation space but lacks identifiability guarantees.
-Activation editing in LLMs.
+Activation editing in LLMs
 Methods such as representation engineering (Zou et al., 2023), contrastive activation addition (Rimsky et al., 2024; Turner et al., 2024) and activation patching (Meng et al., 2022) manipulate model internals but do not address whether control directions are uniquely determined.
 Work on neural network symmetries (Entezari et al., 2021; Dinh et al., 2017) motivates our null-space analysis.
-Robustness and invariant representations.
+Robustness and invariant representations
 Work on distribution shift and domain adaptation (Rabanser et al., 2019; Lipton et al., 2018) establishes that learned representations often exploit spurious correlations.
@@
 Recent work on multi-environment causal representation learning (Ahuja et al., 2022; Von Kügelgen et al., 2021; Rojas-Carulla et al., 2018) proves that diverse training environments enable identification of causal factors under structural assumptions.
-Mechanistic interpretability.
+Mechanistic interpretability
 Work on mechanistic interpretability (Olsson et al., 2022; Elhage et al., 2021; Nanda et al., 2023) studies circuits and features in transformers, often assuming the existence of interpretable directions without formalizing identifiability constraints.
@@
 For a given input prompt $x$ (tokenized as $x_{1},\ldots,x_{T}$), let $h_{\ell}(x)\in\mathbb{R}^{d}$ denote the hidden representation at layer $\ell$ and position $T$ (typically the final token position for autoregressive generation).
-Latent persona variable.
+Latent persona variable
 We assume there exists an underlying latent variable $z\in\mathcal{Z}$ representing a semantic attribute or "persona" (e.g., formality, political stance, truthfulness).
-Steering intervention.
+Steering intervention
 A steering vector $v\in\mathbb{R}^{d}$ is applied as:
@@
 The modified representation $\tilde{h}_{\ell}$ is fed forward through subsequent layers to produce output logits $o(x,v,\alpha)$ over the vocabulary.
-Generative model.
+Generative model
 We posit that the true data-generating process involves:
@@
 We consider two data access regimes that determine what alignment interventions can afford:
-  • Regime 1: Black-box input–output.
-The researcher observes only $(x,y)$ pairs, where $y$ is generated text.
+  • Regime 1: Black-box input–output The researcher observes only $(x,y)$ pairs, where $y$ is generated text.
 There is no access to internal representations.
 This is the weakest regime and corresponds to behavioral evaluation.
-  • Regime 2: White-box single-layer access.
-The researcher can observe or manipulate activations $h_{\ell}(x)$ at a chosen layer $\ell$.
+  • Regime 2: White-box single-layer access The researcher can observe or manipulate activations $h_{\ell}(x)$ at a chosen layer $\ell$.
 This is the standard setting for most steering work and includes extracting vectors from contrastive prompt pairs.
@@
 3.3 Linear Approximation and Nonlinear Case
-Local linearization.
+Local linearization
 Near a reference distribution, we can approximate the effect of steering on output logits as:
@@
 where $J_{\ell}(x)=\left.\frac{\partial o}{\partial h_{\ell}}\right|_{h_{\ell}(x)}\in\mathbb{R}^{V\times d}$ denotes the Jacobian of the output and $V$ is the vocabulary size.
-Nonlinear case.
+Nonlinear case
 In general, the mapping $h_{\ell}\mapsto o$ involves multiple nonlinear layers (attention, MLPs, layer norms). We denote this as:
@@
 3.4 Assumptions
-We make the following assumptions explicit:
-A1 (Smoothness).
-The functions $g_{\ell}$ and $F_{\ell\to L}$ are differentiable almost everywhere, enabling local linear approximation via Jacobians $J_{\ell}(x)=\frac{\partial o}{\partial h_{\ell}}(x)$.
-A2 (Identifiable prompts).
-The prompt distribution $p(x)$ has sufficient variability to probe different aspects of the latent persona $z$.
+We make the following assumptions explicit.
+A1 (Smoothness) The functions $g_{\ell}$ and $F_{\ell\to L}$ are differentiable almost everywhere, enabling local linear approximation via Jacobians $J_{\ell}(x)=\frac{\partial o}{\partial h_{\ell}}(x)$.
+A2 (Identifiable prompts) The prompt distribution $p(x)$ has sufficient variability to probe different aspects of the latent persona $z$.
 Formally, the support of $p(x)$ is rich enough that different persona values induce distinguishable activation patterns $h_{\ell}(x)$.
-A3 (Non-degeneracy).
-The Jacobian $J_{\ell}(x)$ has rank at least $k\geq 1$ for typical $x\sim p(x)$, meaning steering can affect outputs.
+A3 (Non-degeneracy) The Jacobian $J_{\ell}(x)$ has rank at least $k\geq 1$ for typical $x\sim p(x)$, meaning steering can affect outputs.
 This excludes pathological cases where all perturbations to $h_{\ell}$ are ignored by subsequent layers.
-We do not assume:
-  • Statistical independence between $z$ and $x$ (confounding is allowed)
-  • Linearity of $g_{\ell}$ in $z$
-  • Uniqueness of the latent representation without additional structure
-These assumptions are mild and satisfied by standard transformer architectures under typical operating conditions.
+We do not assume: (i) statistical independence between $z$ and $x$ (confounding is allowed), (ii) linearity of $g_{\ell}$ in $z$ or (iii) uniqueness of the latent representation without additional structure.
+These assumptions are mild and generally satisfied by standard transformer architectures under typical operating conditions.
 4 Definitions and Identifiability
 4.1 Identifiability
-Definition 1 (Parameter Identifiability).
+Definition 1 (Parameter Identifiability)
 A parameter $\theta$ in a statistical model $p(y\mid x;\theta)$ is identifiable if for any $\theta^{\prime}\neq\theta$, there exists a distribution over observations $(x,y)$ such that $p(y\mid x;\theta)\neq p(y\mid x;\theta^{\prime})$.
@@
 We say $v$ is identifiable if no other vector $v^{\prime}\neq v$ (up to scaling) produces the same distribution over observable outputs across all prompts and steering strengths.
-Definition 2 (Observational Equivalence).
+Definition 2 (Observational Equivalence)
 Two steering vectors $v$ and $v^{\prime}$ are observationally equivalent in regime $\mathcal{R}$ if they produce identical distributions over all quantities observable in $\mathcal{R}$.
@@
 $v\sim_{\mathcal{R}}v^{\prime}$ $\iff F_{\ell\to L}(h_{\ell}(x)+\alpha v)$ (6)
-$=F_{\ell\to L}(h_{\ell}(x)+\alpha v^{\prime})\;\forall x,\alpha.$
+$\quad=F_{\ell\to L}(h_{\ell}(x)+\alpha v^{\prime})\;\forall x,\alpha.$
 4.2 Symmetries and Gauge Freedom
-Scaling ambiguity.
+Scaling ambiguity
 For any $c\neq 0$, the vectors $v$ and $cv$ produce outputs that differ only by a rescaling of $\alpha$. This is unavoidable; we consider $v$ and $cv$ as the same direction.
-Null space ambiguity.
+Null space ambiguity
 If $v_{0}\in\ker(J_{\ell})$ (i.e., $J_{\ell}v_{0}=0$), then adding $v_{0}$ to any steering vector does not change the linearized output.
@@
 We now state our main theoretical result, proving that persona vectors are fundamentally non-identifiable under standard observational regimes.
-Proposition 1.
-Under Assumptions A1–A3, in Regime 2 (white-box single-layer access) without additional structural constraints, persona vectors are not identifiable.
-Specifically, for any steering vector $v\in\mathbb{R}^{d}$, there exist infinitely many vectors $v^{\prime}\not\propto v$ that are observationally equivalent.
-Proof Sketch.
-We establish non-identifiability via two complementary arguments: null-space ambiguity and reparameterization symmetry.
-Null-space ambiguity (primary mechanism).
-Under the local linear approximation $o\approx o_{0}+\alpha J_{\ell}v$, any vector $v^{\prime}=v+v_{0}$ where $v_{0}\in\ker(J_{\ell})$ satisfies $J_{\ell}v^{\prime}=J_{\ell}v$, and thus produces identical linearized outputs.
-This establishes non-identifiability even without invoking global reparameterization symmetries.
-Reparameterization symmetry.
-Neural networks exhibit inherent parameter symmetries arising from overparameterization.
-Consider an invertible linear transformation $T(h)=Ah$ where $A\in\mathbb{R}^{d\times d}$.
-For layer $\ell+1$ with weight matrix $W_{\ell+1}$ and bias $b_{\ell+1}$, the computation $h_{\ell+1}=\sigma(W_{\ell+1}h_{\ell}+b_{\ell+1})$ can be equivalently expressed through reparameterized weights $W^{\prime}_{\ell+1}=W_{\ell+1}A^{-1}$:
-$h^{\prime}_{\ell+1}$ $=\sigma\!\left(W^{\prime}_{\ell+1}Ah_{\ell}+b_{\ell+1}\right)$ (7)
-$=\sigma\!\left(W_{\ell+1}h_{\ell}+b_{\ell+1}\right)$
-$=h_{\ell+1}$
-Under this reparameterization, steering with $v^{\prime}=Av$ on the transformed representation $h^{\prime}_{\ell}=Ah_{\ell}$ produces identical outputs to steering with $v$ on $h_{\ell}$.
-For any invertible matrix $A\neq cI$ (not a scalar multiple of identity), we obtain $v^{\prime}\not\propto v$, yet these vectors are observationally equivalent.
-Since the space of such matrices has dimension $d^{2}-1$, infinitely many distinct reparameterizations exist.
-Practical vs.
-theoretical relevance: The reparameterization argument establishes non-identifiability as a fundamental property of the model class, independent of any specific instantiation.
-While these symmetries cannot be exploited in frozen models without modifying weights, their existence proves that identifiability fails even in principle.
-For analyzing deployed models, the null-space mechanism provides direct, operationally relevant non-identifiability.
-Corollary 1.1 (Null-space equivalence).
-Under the linear approximation $o\approx o_{0}+\alpha J_{\ell}v$, any vector $v^{\prime}=v+v_{0}$ where $v_{0}\in\ker(J_{\ell})$ is observationally equivalent to $v$, since $J_{\ell}v^{\prime}=J_{\ell}v+J_{\ell}v_{0}=J_{\ell}v$.
-Remark (Null-space dimensionality).
-The null space $\ker(J_{\ell})$ is typically high-dimensional in practice.
-For a Jacobian $J_{\ell}\in\mathbb{R}^{V\times d}$ with vocabulary size $V$ and hidden dimension $d$, the maximum possible rank is $\min(V,d)$.
-In modern language models, $V\approx 50{,}000$ and $d\approx 4{,}000$, so the maximal rank is $d$.
-However, empirical evidence shows that (1) neural network representations exhibit low intrinsic dimensionality (Li et al., 2018; Maennel et al., 2018), and (2) output sensitivity concentrates in a small number of Hessian directions (Ghorbani et al., 2019).
-These findings suggest the Jacobian $J_{\ell}$ has effective rank substantially smaller than $d$, implying a high-dimensional null space:
-$\dim(\ker(J_{\ell}))=d-\mathrm{rank}(J_{\ell})$ (8)
+Proposition 1
+Under Assumptions A1–A3 within the local linear approximation, in Regime 2 (white-box single-layer access) without additional structural constraints, persona vectors are not identifiable.
+Specifically, for any steering vector $v\in\mathbb{R}^{d}$, there exist infinitely many vectors $v^{\prime}\not\propto v$ that are observationally equivalent under linearization.
+Proof Sketch
+We establish non-identifiability via the null-space ambiguity argument. A related weight-space symmetry result is given in (Appendix B.2).
+Null-space ambiguity (primary mechanism). Under the local linear approximation (Equation 4), the steered output is
+$o(x,v,\alpha)\approx o_{0}(x)+\alpha J_{\ell}(x)v.$ (7)
+For any $v_{0}\in\ker(J_{\ell})$, i.e., $J_{\ell}v_{0}=0$, define $v^{\prime}=v+v_{0}$. Then
+$o(x,v^{\prime},\alpha)$ $\approx o_{0}(x)+\alpha J_{\ell}(x)(v+v_{0})$ (8)
+$=o_{0}(x)+\alpha J_{\ell}(x)v\approx o(x,v,\alpha),$
+Since $\dim(\ker(J_{\ell}))=d-\mathrm{rank}(J_{\ell})$, any case where $\mathrm{rank}(J_{\ell})<d$ implies a non-trivial null space.
+In overparameterized language models the effective rank of $J_{\ell}$ is often smaller than $d$ (Oymak et al., 2019), yielding $\dim(\ker(J_{\ell}))\geq 1$ in practice.
+Hence infinitely many such $v_{0}$ exist, producing observationally equivalent vectors $v^{\prime}\not\propto v$.
+Remark (Scope of linearization)
+Proposition 1 is proved under the local linear approximation (Equation 4), which holds exactly for piecewise-linear activations (ReLU MLPs on activation-region interiors) and as a first-order approximation for smooth nonlinearities (GeLU, attention softmax).
+For small steering strengths $\alpha$, the nonlinear residual is $O(\alpha^{2})$ and empirically negligible.
+Extending the result to the exact nonlinear regime $o=F_{\ell\to L}(h_{\ell}+\alpha v)$ would require characterizing the Hessian structure of $F_{\ell\to L}$ and is left for future work.
+Corollary 1.1 (Null-space equivalence)
+Under the linear approximation $o\approx o_{0}+\alpha J_{\ell}v$, any vector $v^{\prime}=v+v_{0}$ where $v_{0}\in\ker(J_{\ell})$ is observationally equivalent to $v$ since $J_{\ell}v^{\prime}=J_{\ell}v+J_{\ell}v_{0}=J_{\ell}v$.
 6 Empirical Validation
@@
 Our empirical experiments test a behavioral consequence of the theoretical prediction: the existence of large equivalence classes of semantically indistinguishable steering directions, without directly estimating the Jacobian null space.
+A complementary logit-level analysis in Appendix A further examines output distribution preservation beyond the semantic probes used in the main experiments.
 6.1 Experimental Setup
-Models and Layers.
-We evaluate two open-weight instruction-tuned language models to test generality across architectures and scales: Qwen2.5-3B-Instruct (layer $\ell=12$) and Llama-3.1-8B-Instruct (layer $\ell=16$).
+Models and Layers
+We evaluate two open-weight instruction-tuned language models to test generality across architectures and model scales: Qwen2.5-3B-Instruct (layer $\ell=12$) and Llama-3.1-8B-Instruct (layer $\ell=16$) (Yang et al., 2024; Meta AI, 2024).
 For both models, we focus on middle layers ($\ell=L/2$), following standard practice (Chen et al., 2025; Konen et al., 2024; Sun et al., 2025) in steering research.
-Persona traits.
+Persona traits
 We test three semantic traits spanning distinct dimensions: Formality, Politeness and Humor.
 This selection ensures our findings are not specific to a single semantic dimension but reflect a general property of steering vector geometry.
-Steering vector extraction.
+Steering vector extraction
 For each trait, we construct 50 contrastive prompt pairs designed to elicit contrasting persona values.
@@
 $v=\frac{1}{50}\sum_{i=1}^{50}[h_{\ell}(x^{+}i)-h\ell(x^{-}_{i})]$ (9)
-Semantic probes.
+Semantic probes
 We evaluate semantic equivalence using trait-specific scoring functions $\phi(o)$ that map generated text to real-valued scores.
@@
 These return scores in $[0,1]$ where 1 is maximally formal/polite/humorous.
-Orthogonality test methodology.
+Direct distributional evidence is provided by the logit-level analysis in Appendix A.
+Orthogonality test methodology
 To test Proposition 1’s prediction that $v$ and $v+v_{\perp}$ (where $v_{\perp}$ is orthogonal) produce observationally equivalent outputs, we implement the following procedure:
@@
 If $v$ and $v^{\prime}$ are observationally equivalent as predicted by Proposition 1, we expect Cohen’s $d<0.2$ (negligible effect) and high correlation between semantic scores.
-Sample size design.
+Sample size design
 We conduct the orthogonality test with both $n=5$ and $n=10$ random orthogonal seeds per trait for Qwen2.5-3B and Llama-3.1-8B.
 This allows us to assess statistical stability across sample sizes and models.
-Scale invariance test.
+Scale invariance test
 To verify that observational equivalence holds across different steering strengths, we additionally evaluate the formality trait at four $\alpha$ values: 0.0,0.5,1.0,2.0{0.0,0.5,1.0,2.0} for both models.
@@
 6.2 Orthogonal Perturbation Test Results
-Table 1: Empirical validation of non-identifiability across models, traits and sample sizes.
-Cohen’s $d$ measures effect size between $v$ and $v+v_{\perp}$ (lower = more equivalent).
-All values show negligible differences ($d<0.2$), confirming observational equivalence.
+Table 1 presents our empirical findings across two models, three semantic traits and two sample sizes ($n=5$ and $n=10$ random orthogonal seeds).
+Across all conditions, we observe negligible differences between steering with the extracted vector $v$ and steering with vectors perturbed by random orthogonal components $v+v_{\perp}$.
+We measure steering efficacy using semantic probe scores in $[0,1]$ that quantify the intensity of the target trait in generated outputs, comparing score distributions from steered generations against baseline generations.
+Table 1: Empirical validation of non-identifiability across models, traits, and sample sizes.
+Cohen’s $d$ measures the effect size between $v$ and $v+v_{\perp}$ (smaller = more equivalent).
+All results show negligible differences ($d<0.2$), supporting observational equivalence.
 Model Trait Seeds Cohen’s $d$ Correlation Perp-Only
@@
 $n=10$ $119\pm 0.119 0.0$$016\pm 0.104$ $95.9\%$
-Table 1 presents our empirical findings across two models, three semantic traits and two sample sizes ($n=5$ and $n=10$ random orthogonal seeds).
-Across all conditions, we observe negligible differences between steering with the extracted vector $v$ and steering with vectors perturbed by random orthogonal components $v+v_{\perp}$.
-We measure steering efficacy using semantic probe scores in $[0,1]$ that quantify the intensity of the target trait in generated outputs, comparing score distributions from steered generations against baseline generations.
 With $n=10$ seeds, the mean Cohen’s $d$ is 0.080 for Qwen2.5-3B and 0.100 for Llama-3.1-8B, both well below the threshold for small effects ($d=0.2$) and firmly in the negligible range.
+Near-zero correlations (e.g., $0.044$, 0.032-0.032) reflect probe ceiling effects rather than independence; Cohen’s $d$ is the more reliable metric in these cases.
 The “Perp-Only Effect” column measures the efficacy of steering with pure orthogonal components ($v_{\perp}$ alone, without $v$) relative to the extracted vector—values near $100\%$ indicate that orthogonal components achieve equivalent behavioral impact.
-Statistical stability across sample sizes.
-The convergence between $n=5$ and $n=10$ results demonstrates robustness across both models.
+Values marginally above $100\%$ reflect probe ceiling effects at the upper boundary of the scoring range.
+Statistical stability across sample sizes The convergence between $n=5$ and $n=10$ results demonstrates robustness across both models.
 For Qwen2.5-3B, effect sizes change by <0.07<0.07 as sample size increases, with all traits showing tightening confidence intervals: formality ($0.144\to 0.075$), politeness ($0.112\to 0.092$), humor ($0.103\to 0.072$).
@@
 This consistency demonstrates that the observed equivalence is not a sampling artifact but a stable property of the steering geometry.
-Cross-model consistency.
-The close agreement between Qwen2.5-3B ($d=0.080$) and Llama-3.1-8B ($d=0.100$) demonstrates that observational equivalence is not model-specific.
+Cross-model consistency The close agreement between Qwen2.5-3B ($d=0.080$) and Llama-3.1-8B ($d=0.100$) demonstrates that observational equivalence is not model-specific.
 Despite differences in architecture, scale (3B vs.
@@
 The consistency across traits within each model further confirms that the phenomenon is general rather than an artifact of specific semantic domains or model implementations.
-Visualizing orthogonal component equivalence.
-Figure 2 illustrates the perp-only effect ratios across all traits and models using $n=10$ orthogonal seeds.
+Visualizing orthogonal component equivalence Figure 2 illustrates the perp-only effect ratios across all traits and models using $n=10$ orthogonal seeds.
 The tight clustering around the perfect equivalence line (dashed red at $1.0$) demonstrates that pure orthogonal components achieve nearly identical steering efficacy to the extracted vectors.
@@
 (b) Llama-3.1-8B-Instruct
-Figure 2: Perp-only effect ratios ($v_{\perp}$ efficacy / $v$ efficacy) for $n=10$ orthogonal seeds. Values near 1.0 (dashed red line) indicate perfect equivalence.
+Figure 2: Perp-only effect ratios ($v_{\perp}$ efficacy / $v$ efficacy) for $n=10$ orthogonal seeds. Values near 1.0 indicate equivalence.
 […]
@@
 We test whether equivalence classes persist across distributional shifts or whether prompt diversity alone induces identifiability.
-We evaluate orthogonal perturbation equivalence across four distinct prompt environments: in-distribution (standard trait templates), topic shift (medical, legal, creative writing contexts), genre shift (emails, tweets, academic writing), and safety-style (refusal patterns, ethical dilemmas).
+We evaluate orthogonal perturbation equivalence across four distinct prompt environments: in-distribution (standard trait templates), topic shift (medical, legal, creative writing contexts), genre shift (emails, tweets, academic writing) and safety-style (refusal patterns, ethical dilemmas).
 For each environment, we extract steering vectors and evaluate whether $v\equiv v+v_{\perp}$ equivalence holds when testing on that environment’s prompts.
-Table 2 and Figure 4 present Cohen’s $d$ effect sizes across all trait-environment combinations for both models.
+Table 2 presents Cohen’s $d$ effect sizes across all trait-environment combinations for both models.
 Across 24 conditions, we observe effect sizes ranging from $d=0.21$ to $d=0.63$ (mean $d=0.35$).
 Effect sizes are predominantly small ($d<0.5$ for 88% of conditions; mean $d=0.35$), indicating that $v$ and $v+v_{\perp}$ remain largely observationally equivalent across diverse extraction contexts.
-Table 2: Multi-environment validation. Cohen’s $d$ effect sizes across extraction contexts. ID = in-distribution; Top = topic shift; Gen = genre shift; Saf = safety-style prompts.
-Model Trait ID Top Gen Saf
+Table 2: Multi-environment validation. Cohen’s $d$ effect sizes across prompt contexts (in-dis: in-distribution, top: topic shift, gen: genre shift, saf: safety-style).
+Model Trait In-dis Top Gen Saf
 Qwen-3B Formality 0.44 0.25 0.29 0.21
@@
 Humor 0.29 0.56 0.30 0.40
-Figure 4: Multi-environment validation results.
-Environment-dependent effect sizes and cross-environment matrices show small-to-moderate Cohen’s $d$ values across traits and prompt contexts, confirming robustness of observational equivalence under distributional shift.
 Formality shows most consistency (mean $d=0.32$), with particularly strong equivalence in safety-style contexts ($d=0.21$ for both models).
@@
 Humor demonstrates intermediate stability (mean $d=0.39$), with genre-shift producing notably lower effects ($d=0.24$ for Qwen, $d=0.30$ for Llama).
-Figure 4 visualizes these patterns through environment-dependent effect sizes (left panels) and cross-environment transfer matrices (right panels), showing that equivalence remains stable even when vectors are extracted in one environment and tested in another.
-Implications.
-The persistence of small effect sizes across diverse prompt distributions demonstrates that non-identifiability is not an artifact of in-distribution prompts but a fundamental geometric property.
-Critically, this shows that naive distributional diversity without explicit invariance constraints or structural assumptions is insufficient to recover identifiable representations.
-7 Conclusion
-We have established that steering vectors are fundamentally non-identifiable under standard observational regimes, with infinitely many distinct directions producing equivalent behavioral effects.
-This non-identifiability persists across models, traits, and prompt distributions, revealing a fundamental tension: methods optimized for behavioral control may not yield interpretable representations.
-Our work clarifies when steering provides reliable heuristic modification versus principled causal understanding, highlighting that interpretability claims require validation beyond input-output equivalence alone.
+Figure 4 visualizes these patterns through environment-dependent effect sizes, showing that equivalence remains stable even when vectors are extracted in one environment and tested in another.
+Figure 4: Environment-dependent effect sizes across prompt contexts.
+Implications
+The persistence of small effect sizes across diverse prompt distributions indicates that non-identifiability is not an artifact of in-distribution prompts but a structural geometric property.
+Cases where $d$ approaches 0.50.50.60.6 (e.g., politeness under genre or safety-style shifts) suggest that certain trait–context combinations impose partial constraints without eliminating null-space ambiguity since $\ker(J)$ is determined by model weights rather than the prompt distribution.
+Although effect sizes increase under distribution shift (reaching $d=0.63$ in some settings), distributional diversity alone does not eliminate observational equivalence.
 7 Limitations
@@
 While our theoretical results apply generally, the empirical magnitude of non-identifiability may vary across model families, scales, architectures and layer positions.
-Our semantic evaluation uses lexical heuristics rather than full distributional metrics or human judgments, providing a conservative test of observational equivalence.
-Our multi-environment experiments demonstrate that equivalence persists across diverse prompt distributions, but we do not test whether explicit structural interventions such as independence constraints (ICA), sparsity regularization, or invariance objectives (IRM) can recover identifiable representations.
-While theoretical frameworks suggest these approaches might enable identifiability, empirical validation across multiple models and experimental designs remains an important direction for future work.
+Our semantic evaluation uses lexical heuristics rather than full distributional metrics or human evaluation, providing a conservative test of observational equivalence.
+Our multi-environment experiments demonstrate that equivalence persists across diverse prompt distributions but we do not test whether explicit structural interventions such as independence constraints (ICA), sparsity regularization or invariance objectives (IRM) can recover identifiable representations.
+Empirical validation of these approaches remains an important direction for future work.
 Additionally, we do not directly measure null-space dimensionality or test whether null-space versus row-space components exhibit differential robustness under distribution shift.
@@
 Our results establish the necessity of validation beyond input-output behavior.
-Important directions include testing whether structural constraints can recover identifiable representations, analyzing how non-identifiability varies across network layers and architectural components, and characterizing how null-space dimensionality scales with model size.
+Important directions include testing whether structural constraints can recover identifiable representations, analyzing how non-identifiability varies across network layers and architectural components and characterizing how null-space dimensionality scales with model size.
 Investigating whether alternative steering methods exhibit superior robustness under adversarial attacks or distribution shift would clarify the practical consequences of non-identifiability.
-Human evaluation would provide stronger validation than our lexical heuristics, and direct measurement of null-space dimensionality would quantify our theoretical predictions.
-10 Acknowledgments
+Human evaluation would provide stronger validation than our lexical heuristics and direct measurement of null-space dimensionality would quantify our theoretical predictions.
+9 Conclusion
+We have established that steering vectors are fundamentally non-identifiable under the local linear approximation, with infinitely many distinct directions producing equivalent behavioral effects.
+This non-identifiability persists across models, traits and prompt distributions, revealing a fundamental tension: methods optimized for behavioral control may fail to yield interpretable representations.
+Our work clarifies when steering provides reliable heuristic modification versus principled causal understanding, highlighting that interpretability claims require validation beyond input-output equivalence alone.
+More broadly, when internal interventions admit large equivalence classes under the model’s local geometry, consistent output shifts do not imply that a discovered direction corresponds to a canonical or causally privileged feature.
+Stronger interpretability claims therefore require structural constraints or additional assumptions that break these symmetries and distinguish meaningful factors from geometrically arbitrary ones.
+Acknowledgments
 We thank Vast.ai for providing GPU compute resources that enabled our empirical validation experiments.
-We also acknowledge the HuggingFace platform and maintainers of the open-source models (Qwen2.5-3B-Instruct, Llama-3.1-8B-Instruct) used in this work.
+We also acknowledge the HuggingFace platform and maintainers of the open-source models used in this work.
+Generative AI tools were used for language editing and code assistance; all technical content was reviewed and verified by the authors.
 Appendix A Logit-Level Equivalence Test
@@
 Ratios range from $0.49$ (Llama, formality) to $0.90$ (Qwen, humor), indicating that orthogonal perturbations systematically induce smaller logit shifts than random directions.
-The next-token predictions remain highly stable: token agreement between $v$ and $v+v_{\perp}$ ranges from $81\%$ to $96\%$, and top-10 token overlap ranges from $86\%$ to $96\%$.
+The next-token predictions remain highly stable: token agreement between $v$ and $v+v_{\perp}$ ranges from $81\%$ to $96\%$ and top-10 token overlap ranges from $86\%$ to $96\%$.
 These results provide direct evidence that equivalence is not an artifact of semantic measurement.
 While orthogonal perturbations induce some logit deviation, they systematically preserve output distributions more faithfully than random directions, consistent with Proposition 1’s prediction of high-dimensional equivalence classes.
-Table 3 reports mean $\ell_{2}$ logit distances, normalized ratios, token agreement, and top-$k$ overlap for all models and traits.
-Table 3: Logit-level equivalence statistics.
-Model Trait $R$ (v_perp / random) Token Agreement Top-10 Overlap
+Table 3 reports mean $\ell_{2}$ logit distances, normalized ratios, token agreement and top-$k$ overlap for all models and traits.
+Table 3: Logit-level equivalence statistics. $R$ denotes the ratio $v_{\perp}$ / random. Agreement% refers to token agreement, and Overlap% indicates top-10 overlap.
+Model Trait $R$ Agreement% Overlap%
 Qwen-3B Humor 0.90 81% 86%
-Qwen-3B Formality 0.76 94% 96%
-Qwen-3B Politeness 0.75 93% 96%
+Formality 0.76 94% 96%
+Politeness 0.75 93% 96%
 Llama-8B Humor 0.64 96% 90%
-Llama-8B Formality 0.49 93% 95%
-Llama-8B Politeness 0.65 91% 94%
-Appendix B Prompt Construction Details
-B.1 Contrastive Prompt Pairs
-We construct contrastive prompt pairs for each semantic trait using template-based generation. Table 4 provides concrete examples for all three traits used in the empirical validation.
-Trait Positive ($x^{+}$) Negative ($x^{-}$)
-Formality Draft a professional email summarizing quarterly performance. Text a friend about how the quarter went.
-Politeness Respond diplomatically to criticism from a colleague. Reply bluntly to someone who criticized you.
-Humor Tell a lighthearted story about an embarrassing mistake. Provide a factual account of an embarrassing mistake.
-Table 4: Example contrastive prompt pairs for each semantic trait.
-B.2 Steering Extraction and Evaluation Pipeline
-Figure 5: Steering extraction and evaluation pipeline.
-Figure 5 illustrates the pipeline used to extract and evaluate steering vectors.
-Contrastive prompt pairs are first used to compute a steering vector specific to a semantic trait.
-This vector is then applied to a separate set of held-out evaluation prompts to generate steered outputs.
-Appendix C Detailed Proof of Proposition 1
-We provide proofs for establishing non-identifiability via two complementary arguments: null-space ambiguity and reparameterization symmetry.
-C.1 Null-Space Ambiguity (Constructive Proof)
-Setup. Consider the local linear approximation of the steering effect:
+Formality 0.49 93% 95%
+Politeness 0.65 91% 94%
+Appendix B Detailed Proof of Proposition 1
+B.1 Null-Space Ambiguity (Constructive Proof)
+Setup Consider the local linear approximation of the steering effect:
 $o(x,v,\alpha)\approx o(x,0,0)+\alpha J_{\ell}(x)v$ (11)
 where $J_{\ell}(x)=\frac{\partial o}{\partial h_{\ell}}\big|_{h_{\ell}(x)}\in\mathbb{R}^{V\times d}$ is the Jacobian.
-Step 1: Null space characterization.
+Step 1: Null space characterization
 Define the null space of $J_{\ell}$ as:
@@
 $\dim(\mathcal{N})=d-\operatorname{rank}(J_{\ell})$ (13)
-Step 2: Rank Bound.
+Step 2: Rank Bound
 The Jacobian $J_{\ell}\in\mathbb{R}^{V\times d}$ has maximum possible rank:
 $\text{rank}(J_{\ell})\leq\min(V,d)$ (14)
-In modern language models:
-  • Hidden dimension: $d\approx 4000$ (typical)
-  • Vocabulary size: $V\approx 50000$ (typical)
-  • Therefore: $\max\text{rank}(J_{\ell})=d$
-Step 3: Effective rank is much lower.
+In modern language models, the hidden dimension is typically $d\approx 4000$ while the vocabulary size is $V\approx 50000$. Therefore, the Jacobian satisfies $\max\mathrm{rank}(J_{\ell})=d$.
+Step 3: Effective rank is much lower
 Output distributions lie on a low-dimensional manifold. The effective rank satisfies:
@@
 where $\sigma_{i}$ are singular values of $J_{\ell}$ and $\epsilon$ is a threshold (e.g., $10^{-4}$).
-Intuition: In overparameterized LLMs, the effective rank is expected to be strictly less than $d$ due to the low-dimensional structure of output distributions.
-Therefore $\dim(\mathcal{N})=d-\text{rank}(J_{\ell})$ is generically positive but this is not required for the proof—the argument establishes non-identifiability whenever $\dim(\mathcal{N})\geq 1$.
-Step 4: Constructing equivalent vectors.
+Intuition: Empirical evidence suggests that in overparameterized LLMs the effective rank of $J_{\ell}$ may be substantially lower than $d$, consistent with low dimensional structure observed in related weight space studies (Li et al., 2018; Maennel et al., 2018) and concentration of output sensitivity in a limited number of directions (Ghorbani et al., 2019).
+Consequently, $\dim(\mathcal{N})=d-\text{rank}(J_{\ell})$ is often positive in practice.
+However, this empirical observation is not required for the proof; the argument establishes non identifiability whenever $\dim(\mathcal{N})\geq 1$.
+Step 4: Constructing equivalent vectors
 For any steering vector $v\in\mathbb{R}^{d}$ and any $v_{0}\in\mathcal{N}$, define
@@
 Then for all $x$ and all $\alpha$:
-$J_{\ell}(x)v^{\prime}=J_{\ell}(x)(v+v_{0})=J_{\ell}(x)v+J_{\ell}(x)v_{0}=J_{\ell}(x)v.$ (17)
+$J_{\ell}(x)v^{\prime}$ $=J_{\ell}(x)(v+v_{0})$ (17)
+$=J_{\ell}(x)v+J_{\ell}(x)v_{0}$
+$=J_{\ell}(x)v.$
 Therefore:
-$o(x,v^{\prime},\alpha)\approx o(x,0,0)+\alpha J_{\ell}(x)v^{\prime}=o(x,0,0)+\alpha J_{\ell}(x)v\approx o(x,v,\alpha).$ (18)
-Step 5: Infinitely many distinct solutions.
+$o(x,v^{\prime},\alpha)$ $\approx o(x,0,0)+\alpha J_{\ell}(x)v^{\prime}$ (18)
+$=o(x,0,0)+\alpha J_{\ell}(x)v$
+$\approx o(x,v,\alpha).$
+Step 5: Infinitely many distinct solutions
 Since $\dim(\mathcal{N})\geq 1$, the null space contains infinitely many directions. For any $v_{0}\in\mathcal{N}\setminus\{0\}$ and any $\beta\in\mathbb{R}$:
@@
 Furthermore, if $\beta$ is chosen such that $v^{\prime}_{\beta}\not\propto v$ (which is always possible unless $v\propto v_{0}$), these vectors are geometrically distinct.
-Step 6: Non-proportionality.
+Step 6: Non-proportionality
 To ensure $v^{\prime}_{\beta}\not\propto v$, we need $v+\beta v_{0}\neq cv$ for any scalar $c$. This fails only if:
@@
 Therefore, for generic $v$ and generic $v_{0}\in\mathcal{N}$, we have $v^{\prime}_{\beta}\not\propto v$ for almost all $\beta$.
-Conclusion (Null-space mechanism).
-Under the linear approximation and for typical Jacobian structure, there exist infinitely many geometrically distinct steering vectors that are observationally equivalent.
-C.2 Reparameterization Symmetry (Existence Proof)
-Setup: Neural networks exhibit inherent symmetries arising from overparameterization. We show these symmetries induce non-identifiability even in the exact nonlinear case.
-Step 1: Representation reparameterization.
+Conclusion (Null-space mechanism)
+Under the local linear approximation and for typical Jacobian structure, there exist infinitely many geometrically distinct steering vectors that are observationally equivalent.
+B.2 Weight Space Gauge Symmetry (Supplementary Observation)
+While Proposition 1 fixes model weights and asks whether v is recoverable from behavior, this section offers a complementary perspective: even relaxing the weight-fixing assumption, non-identifiability persists via linear reparameterization symmetries.
+Setup: We record a supplementary observation about gauge symmetry in neural network parameterization.
+This is distinct from the null-space non-identifiability of Proposition 1: rather than fixing model weights and asking whether $v$ is recoverable from behavioral observations, we ask whether different $(W,v)$ and $(W^{\prime},v^{\prime})$ pairs can produce identical input–output behavior.
+The answer is yes, via the linear reparameterization below.
+This symmetry is not realizable as a steering operation on a frozen model as it requires modifying $W_{\ell+1}$.
+Step 1: Representation reparameterization
 Consider an invertible transformation $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$. Define the reparameterized representation:
@@
 then $(h_{\ell},v)$ and $(h^{\prime}_{\ell},v^{\prime})$ where $v^{\prime}=DT(h_{\ell})\cdot v$ are observationally indistinguishable.
-Step 2: Linear reparameterizations.
+Step 2: Linear reparameterizations
 Consider linear transformations $T(h)=Ah$ where $A\in\mathbb{R}^{d\times d}$ is invertible. For layer $\ell+1$ with weight matrix $W_{\ell+1}\in\mathbb{R}^{d\times d}$ and bias $b_{\ell+1}$:
@@
 Then:
-$h^{\prime}_{\ell+1}=\sigma\!\left(W_{\ell+1}A^{-1}Ah_{\ell}+b_{\ell+1}\right)=\sigma\!\left(W_{\ell+1}h_{\ell}+b_{\ell+1}\right)=h_{\ell+1}.$ (25)
-Step 3: Steering under reparameterization.
+$h^{\prime}_{\ell+1}$ $=\sigma\!\left(W_{\ell+1}A^{-1}Ah_{\ell}+b_{\ell+1}\right)$ (25)
+$=\sigma\!\left(W_{\ell+1}h_{\ell}+b_{\ell+1}\right)$
+$=h_{\ell+1}.$
+Step 3: Steering under reparameterization
 Original steering:
@@
 Reparameterized steering with $v^{\prime}=Av$:
-$\tilde{h}^{\prime}_{\ell}=h^{\prime}_{\ell}+\alpha v^{\prime}=Ah_{\ell}+\alpha Av=A(h_{\ell}+\alpha v)=A\tilde{h}_{\ell}$. (28)
+$\tilde{h}^{\prime}_{\ell}$ $=h^{\prime}_{\ell}+\alpha v^{\prime}$ (28)
+$=Ah_{\ell}+\alpha Av$
+$=A(h_{\ell}+\alpha v)$
+$=A\tilde{h}_{\ell}$.
 $\tilde{h}^{\prime}_{\ell+1}$ $=\sigma\!\left(W^{\prime}_{\ell+1}(h^{\prime}_{\ell}+\alpha v^{\prime})+b^{\prime}_{\ell+1}\right)$ (29)
@@
 $=\tilde{h}_{\ell+1}.$
-Step 4: Infinitely many reparameterizations.
+Step 4: Infinitely many reparameterizations
 For any invertible matrix $A$ with $A\neq cI$ (i.e., not a scalar multiple of identity), we obtain $v^{\prime}=Av\not\propto v$.
 The space of such matrices has dimension $d^{2}-1$ (excluding scalar multiples), providing infinitely many distinct reparameterizations.
-Important note: While we cannot explicitly construct these reparameterizations for a frozen model without retraining (since this would require modifying $W_{\ell+1}$), their existence follows from the fundamental symmetry structure of neural networks.
-This establishes that identifiability cannot hold even in principle without additional constraints.
-Practical implication: For frozen, deployed models, only the null-space mechanism is operationally relevant–the reparameterization symmetry is not a realizable operation.
-However, the reparameterization argument shows that even if we allowed weight modifications during training, identifiability would still fail without structural constraints.
-Thus, non-identifiability is both a practical limitation (null-space) and a fundamental theoretical barrier (gauge symmetry).
-Conclusion (Reparameterization mechanism).
-The inherent symmetries of neural network parameterizations induce infinitely many observationally equivalent steering vectors.
-Appendix D Intuitive Explanations
-D.1 Geometric Intuition for Null-Space Ambiguity
-[…]
-Figure 6: Geometric intuition for null-space ambiguity.
-The output observes only the $(v_{1},v_{2})$ components (blue shaded region).
-The $v_{3}$ component lies in $\ker(J)$ and is invisible to the output.
-Adding any $\alpha v_{3}$ to $v$ leaves the output unchanged: $J(v+\alpha v_{3})=Jv$ for all $\alpha\in\mathbb{R}$.
+Conclusion (Weight-space gauge symmetry)
+The symmetries of neural network parameterization imply the existence of infinitely many $(W,v)$ and $(W^{\prime},v^{\prime})$ pairs that yield identical input–output behavior.
+Appendix C Intuitive Explanations
+C.1 Geometric Intuition for Null-Space Ambiguity
 Visual analogy:
@@
 By the rank-nullity theorem, the null space $\ker(J)$ is a 1D subspace since $\dim(\ker(J))=3-\text{rank}(J)=3-2=1$.
+[…]
+Figure 5: Geometric intuition for null-space ambiguity.
+The output depends only on the $(v_{1},v_{2})$ projection (row space of $J$), while $v_{3}\in\ker(J)$ is invisible; hence $J(v+\alpha v_{3})=Jv$.
 For concreteness, suppose $J=[I_{2}\mid 0]$ projects onto the first two coordinates.
@@
 Any vector of the form $v^{\prime}=v+\alpha v_{3}$ for $\alpha\in\mathbb{R}$ produces identical outputs:
-$J(v+\alpha v_{3})=Jv+\alpha J(0,0,1)^{\top}=Jv+\alpha\cdot 0=Jv.$ (30)
+$J(v+\alpha v_{3})$ $=Jv+\alpha Jv_{3}$ (30)
+$=Jv+\alpha\cdot 0\quad(v_{3}\in\ker(J))$
+$=Jv.$
 Since $\alpha$ can take infinitely many values and $v^{\prime}\not\propto v$ for generic choices of $\alpha$, there exist infinitely many geometrically distinct steering vectors that are observationally equivalent.
-D.2 Why More Data Doesn’t Help
+C.2 Why More Data Doesn’t Help
 Common misconception: "If we collect more steering examples, we can pin down the unique vector."
 Why this fails:
-Consider observing steering effects on $N$ prompts $\{x_{1},\ldots,x_{N}\}$. Each observation provides:
+Consider observing steering effects on $N$ prompts $\{x_{1},\ldots,x_{N}\}$.
+More fundamentally, any $v_{0}\in\ker(J_{\ell})$ satisfies $J_{\ell}v_{0}=0$ by definition since this is a property of the model weights and not the prompt distribution.
+Stacking observations across prompts therefore cannot remove directions from $\ker(J_{\ell})$.
+Each observation provides:
 $o_{i}=J_{\ell}(x_{i})v+\eta_{i}$ (31)
@@
 This means adding more prompts does not reduce the null space when all prompts probe the same effective subspace—the ambiguity persists. Figure 6 illustrates this phenomenon.
-D.3 Why Orthogonal Perturbations Preserve Semantics
+C.3 Why Orthogonal Perturbations Preserve Semantics
 Setup: We test steering with $v$ versus $v+v_{\perp}$, where $v_{\perp}\perp v$ and $\|v_{\perp}\|=1$.
@@
 Why?
-Decomposing the perturbation.
+Decomposing the perturbation
 Any perturbation $v_{\perp}$ can be decomposed into observable and invisible components:
@@
 This means $\sim$20–25% of the perturbation is automatically invisible to model outputs.
-Observable effect is diffuse and unstructured.
+Observable effect is diffuse and unstructured
 The actual output change from adding $v_{\perp}$ is:
@@
 Analogy: If $Jv$ is a strong wind blowing consistently north, then $Jv_{\perp,\text{row}}$ is turbulence—it may have comparable energy but lacks directional coherence, so the overall trajectory remains northward.
+Appendix D Prompt Construction Details
+We construct contrastive prompt pairs for each semantic trait using template-based generation. Table 4 provides concrete examples for all three traits used in the empirical validation.
+Table 4: Example contrastive prompt pairs for each semantic trait.
+Trait Positive ($x^{+}$) Negative ($x^{-}$)
+Formality Draft a professional email summarizing quarterly performance. Text a friend about how the quarter went.
+Politeness Respond diplomatically to criticism from a colleague. Reply bluntly to someone who criticized you.
+Humor Tell a lighthearted story about an embarrassing mistake. Provide a factual account of an embarrassing mistake.
+Figure 7: Steering extraction and evaluation pipeline.
+Figure 7 illustrates the pipeline used to extract and evaluate steering vectors.
+Contrastive prompt pairs are first used to compute a steering vector specific to a semantic trait.
+This vector is then applied to a separate set of held-out evaluation prompts to generate steered outputs.
 Appendix E Mathematical Derivations
@@
 Fisher Information Matrix: For the linear Gaussian model $o=Jv+\eta$ with $\eta\sim\mathcal{N}(0,\sigma^{2}I)$:
-$\mathcal{I}(v)=\mathbb{E}_{o}\left[\left(\frac{\partial\log p(o\mid x,v,\alpha)}{\partial v}\right)\left(\frac{\partial\log p(o\mid x,v,\alpha)}{\partial v}\right)^{\!\top}\right]=\frac{1}{\sigma^{2}}J^{\top}J.$ (42)
+$\mathcal{I}(v)$ $=\mathbb{E}_{o}\Bigg[\left(\frac{\partial\log p(o\mid x,v,\alpha)}{\partial v}\right)$ (42)
+$\qquad\left(\frac{\partial\log p(o\mid x,v,\alpha)}{\partial v}\right)^{\!\top}\Bigg]$
+$=\frac{1}{\sigma^{2}}J^{\top}J.$
 Cramér-Rao Lower Bound: The covariance of any unbiased estimator $\hat{v}$ satisfies:
@@
 More data cannot help because the information geometry has infinite uncertainty in null-space directions.
-Appendix F Disclosure of the Use of AI
-We used ChatGPT to assist with refining sections of the manuscript for language and clarity and used Claude Code and GitHub Copilot for coding assistance during implementation.
-All research ideas, technical content, proofs, experiments and final writing decisions are entirely our own and we take full responsibility for the work presented.

### v4

@@
 We show that, under white-box single-layer access, steering vectors are fundamentally non-identifiable due to large equivalence classes of behaviorally indistinguishable interventions.
-Empirically, we find that orthogonal perturbations achieve near-equivalent efficacy with negligible effect sizes across multiple models and traits.
+Empirically, we find that orthogonal perturbations achieve near-equivalent efficacy with negligible effect sizes across multiple models and traits, with pre-trained semantic classifiers confirming equivalence at the output level.
+We estimate null-space dimensionality via SVD of activation covariance matrices and validate that equivalence holds robustly throughout the operationally relevant steering range.
 Critically, we show that non-identifiability is a robust geometric property that persists across diverse prompt distributions.
 These findings reveal fundamental interpretability limits and highlight the need for structural constraints beyond behavioral testing to enable reliable alignment interventions.
-On the Non-Identifiability of Steering Vectors in Large Language Models
-Sohan Venkatesh1thanks: Correspondence: soh.venkatesh@gmail.com        Ashish Mahendran Kurapath1 1Manipal Institute of Technology Bengaluru
 1 Introduction
-Persona vector steering (Chen et al., 2025) has emerged as a popular technique for controlling the behavior of large language models by adding learned directional vectors to intermediate activations.
-This line of work is closely connected to broader efforts in representation engineering (Zou et al., 2023) and activation editing (Rimsky et al., 2024; Turner et al., 2023), where linear directions in activation space are used to modulate model behavior along semantic axes (Elhage et al., 2022; Turner et al., 2024).
-Despite growing adoption in interpretability and alignment research, the theoretical foundations of persona steering remain poorly understood.
-Most existing methods implicitly assume that extracted steering vectors correspond to meaningful and uniquely determined latent factors and that these factors can be directly manipulated to achieve reliable control.
-However, classical results in latent variable modeling and causal inference show that such assumptions are often unjustified without additional structural constraints (Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Schölkopf et al., 2021; Locatello et al., 2019).
-This raises a fundamental question: when does representational alignment afford reliable behavioral control?
-Understanding identifiability is critical for several reasons.
-First, it determines when persona vectors provide principled causal interventions versus merely heuristic control that exploits one of many behaviorally equivalent directions.
-Second, it clarifies when claims that a specific vector “represents” a semantic concept are scientifically well-grounded rather than artifacts of measurement and projection (Elazar et al., 2021; Marks and Tegmark, 2023).
-Third, it reveals when steering directions may rely on fragile correlations that fail under distribution shift, model updates or adversarial prompting.
-Finally, identifiability theory clarifies which experimental protocols provide meaningful evidence for causal interpretation of interventions on internal activations (Schölkopf et al., 2021; Ahuja et al., 2022).
+Activation steering has emerged as a prominent technique for controlling large language model behavior (Turner et al., 2023; Rimsky et al., 2024), with applications spanning persona modulation (Chen et al., 2025), representation engineering (Zou et al., 2023) and alignment research (Turner et al., 2024).
+The core procedure is simple: extract a directional vector from model activations by contrasting prompts that differ along a semantic axis, then add that vector at inference time to shift model outputs in the desired direction.
+This simplicity has driven broad adoption across mechanistic interpretability (Elhage et al., 2022; Nanda et al., 2023) and alignment (Zou et al., 2023).
+Underlying this practice is a largely unexamined assumption: that the extracted steering direction is meaningful in a precise sense.
+Specifically, it is assumed that the direction is uniquely recoverable from the model’s input–output behavior and that the behavioral effect of steering reflects the genuine causal role of that direction in the model’s computation.
+Without this assumption, the interpretability value of steering is unclear.
+A vector that shifts outputs in the intended direction does not, by itself, reveal which internal direction is responsible for the semantic attribute in question.
+The claim that a steering vector represents a concept such as honesty or formality presupposes that the recovered direction is the unique one doing the causal work.
 Figure 1: Steering vector non-identifiability: The original vector $v$ and a perturbed vector $v+v_{\perp}$ with $v_{\perp}\in\mathrm{ker}(J)$ produce identical outputs since null-space components do not affect JvJv, the first-order change in output logits induced by steering.
+We formalize this argument and validate it empirically.
+The key mechanism is that the Jacobian mapping activations to output logits has a large null space in overparameterized language models (Oymak et al., 2019): any direction in this null space can be added to a steering vector without altering any observable output, producing infinitely many behaviorally equivalent interventions.
+Experiments across five semantic traits and two model families confirm that orthogonal perturbations to a steering vector are behaviorally indistinguishable from the original, that vectors extracted by entirely different methods are semantically equivalent despite being geometrically distinct and that these properties persist across diverse prompt distributions.
 We make three contributions.
-First, we provide a formal identifiability analysis of persona vector steering, proving that under white-box single-layer access, persona vectors are fundamentally non-identifiable: infinitely many geometrically distinct directions induce identical observable behavior via null-space ambiguity.
-Second, we empirically demonstrate that extracted steering vectors contain substantial non-identifiable components, with orthogonal perturbations achieving 95–100% of the original steering efficacy across multiple models and traits.
-Third, we validate through multi-environment experiments that this equivalence persists under distribution shift, confirming non-identifiability is a robust geometric property rather than just an artifact of limited prompt diversity.
-These results clarify when alignment methods can be trusted and which structural assumptions are required for reliable control.
+First, we provide a formal identifiability analysis of steering vectors, proving non-identifiability via a constructive null-space argument under mild assumptions.
+Second, we demonstrate non-identifiability empirically across multiple models, traits, extraction methods, and prompt environments, using pre-trained semantic classifiers and raw logit comparisons as complementary evaluation methods.
+Third, we argue that the implications extend beyond steering: any intervention-based interpretability method that validates representational claims solely through behavioral output faces the same fundamental limitation and structural constraints are necessary for reliable interpretability and alignment.
 2 Related Work
-Our work formalizes persona steering [^1] We use persona vectors and steering vectors interchangeably to refer to directional vectors added to model activations to modulate semantic or behavioral attributes (Chen et al., 2025).
-as a latent variable identification problem, bridging causal representation learning, activation editing in LLMs and mechanistic interpretability.
-Causal and latent variable identifiability
+Our work is the first to formally analyze activation steering [^1] We use persona vectors and steering vectors interchangeably to refer to directional vectors added to model activations to modulate semantic or behavioral attributes (Chen et al., 2025).
+as a latent variable identification problem, connecting two lines of research: causal representation learning and activation editing in LLMs.
+Causal and latent variable identifiability.
 Classical results show that latent variable models are non-identifiable without structural assumptions (Hyvärinen and Pajunen, 1999; Shimizu et al., 2006; Kruskal, 1977).
 Recent work in causal representation learning extends these ideas to deep learning settings (Schölkopf et al., 2021; Ahuja et al., 2022; Locatello et al., 2019), establishing conditions under which latent factors can be recovered from high-dimensional observations.
-Probing and representation learning
+Probing and representation learning.
 Work on probing classifiers (Pimentel et al., 2020; Ravfogel et al., 2020; Elazar et al., 2021) shows that linear directions can succeed for reasons unrelated to target information.
 The linear representation hypothesis (Park et al., 2023; Elhage et al., 2022) suggests concepts correspond to directions in activation space but lacks identifiability guarantees.
-Activation editing in LLMs
+Activation editing in LLMs.
 Methods such as representation engineering (Zou et al., 2023), contrastive activation addition (Rimsky et al., 2024; Turner et al., 2024) and activation patching (Meng et al., 2022) manipulate model internals but do not address whether control directions are uniquely determined.
 Work on neural network symmetries (Entezari et al., 2021; Dinh et al., 2017) motivates our null-space analysis.
-Robustness and invariant representations
+Robustness and invariant representations.
 Work on distribution shift and domain adaptation (Rabanser et al., 2019; Lipton et al., 2018) establishes that learned representations often exploit spurious correlations.
@@
 Recent work on multi-environment causal representation learning (Ahuja et al., 2022; Von Kügelgen et al., 2021; Rojas-Carulla et al., 2018) proves that diverse training environments enable identification of causal factors under structural assumptions.
-Mechanistic interpretability
+Mechanistic interpretability.
 Work on mechanistic interpretability (Olsson et al., 2022; Elhage et al., 2021; Nanda et al., 2023) studies circuits and features in transformers, often assuming the existence of interpretable directions without formalizing identifiability constraints.
@@
 For a given input prompt $x$ (tokenized as $x_{1},\ldots,x_{T}$), let $h_{\ell}(x)\in\mathbb{R}^{d}$ denote the hidden representation at layer $\ell$ and position $T$ (typically the final token position for autoregressive generation).
-Latent persona variable
-We assume there exists an underlying latent variable $z\in\mathcal{Z}$ representing a semantic attribute or "persona" (e.g., formality, political stance, truthfulness).
-Steering intervention
+Latent persona variable.
+We assume there exists an underlying latent variable $z\in\mathcal{Z}$ representing a semantic attribute or ”persona” (e.g., formality, political stance, truthfulness).
+Steering intervention.
 A steering vector $v\in\mathbb{R}^{d}$ is applied as:
-$\tilde{h}\ell(x)=h\ell(x)+\alpha v$ (1)
+$\tilde{h}_{\ell}(x)=h_{\ell}(x)+\alpha v$ (1)
 where $\alpha\in\mathbb{R}$ is the steering strength.
 The modified representation $\tilde{h}_{\ell}$ is fed forward through subsequent layers to produce output logits $o(x,v,\alpha)$ over the vocabulary.
-Generative model
+Generative model.
 We posit that the true data-generating process involves:
@@
 We consider two data access regimes that determine what alignment interventions can afford:
-  • Regime 1: Black-box input–output The researcher observes only $(x,y)$ pairs, where $y$ is generated text.
+  • Regime 1: Black-box input–output.
+The researcher observes only $(x,y)$ pairs, where $y$ is generated text.
 There is no access to internal representations.
 This is the weakest regime and corresponds to behavioral evaluation.
-  • Regime 2: White-box single-layer access The researcher can observe or manipulate activations $h_{\ell}(x)$ at a chosen layer $\ell$.
+  • Regime 2: White-box single-layer access.
+The researcher can observe or manipulate activations $h_{\ell}(x)$ at a chosen layer $\ell$.
 This is the standard setting for most steering work and includes extracting vectors from contrastive prompt pairs.
-Most existing work operates in Regime 2, often extracting a steering vector from contrastive prompt pairs $(x^{+},x^{-})$ designed to elicit different persona values (e.g., polite versus rude instructions):
-$v\propto\mathbb{E}{x^{+}}[h\ell(x^{+})]-\mathbb{E}{x^{-}}[h\ell(x^{-})]$ (3)
+Most existing work operates in Regime 2, often extracting a steering vector from contrastive prompt pairs $(x^{+},x^{-})$ designed to elicit different persona values:
+$v\propto\mathbb{E}_{x^{+}}[h_{\ell}(x^{+})]-\mathbb{E}_{x^{-}}[h_{\ell}(x^{-})]$ (3)
 Our analysis focuses primarily on Regime 2, as it represents the dominant paradigm in current steering research.
 3.3 Linear Approximation and Nonlinear Case
-Local linearization
+Local linearization.
 Near a reference distribution, we can approximate the effect of steering on output logits as:
@@
 where $J_{\ell}(x)=\left.\frac{\partial o}{\partial h_{\ell}}\right|_{h_{\ell}(x)}\in\mathbb{R}^{V\times d}$ denotes the Jacobian of the output and $V$ is the vocabulary size.
-Nonlinear case
+Nonlinear case.
 In general, the mapping $h_{\ell}\mapsto o$ involves multiple nonlinear layers (attention, MLPs, layer norms). We denote this as:
@@
 4.1 Identifiability
-Definition 1 (Parameter Identifiability)
+Definition 1 (Parameter Identifiability).
 A parameter $\theta$ in a statistical model $p(y\mid x;\theta)$ is identifiable if for any $\theta^{\prime}\neq\theta$, there exists a distribution over observations $(x,y)$ such that $p(y\mid x;\theta)\neq p(y\mid x;\theta^{\prime})$.
@@
 We say $v$ is identifiable if no other vector $v^{\prime}\neq v$ (up to scaling) produces the same distribution over observable outputs across all prompts and steering strengths.
-Definition 2 (Observational Equivalence)
+Definition 2 (Observational Equivalence).
 Two steering vectors $v$ and $v^{\prime}$ are observationally equivalent in regime $\mathcal{R}$ if they produce identical distributions over all quantities observable in $\mathcal{R}$.
 For Regime 2 (white-box single-layer access):
-$v\sim_{\mathcal{R}}v^{\prime}$ $\iff F_{\ell\to L}(h_{\ell}(x)+\alpha v)$ (6)
-$\quad=F_{\ell\to L}(h_{\ell}(x)+\alpha v^{\prime})\;\forall x,\alpha.$
+$v\sim_{\mathcal{R}}v^{\prime}\iff F_{\ell\to L}(h_{\ell}(x)+\alpha v)=F_{\ell\to L}(h_{\ell}(x)+\alpha v^{\prime})\;\forall x,\alpha.$ (6)
 4.2 Symmetries and Gauge Freedom
-Scaling ambiguity
+Scaling ambiguity.
 For any $c\neq 0$, the vectors $v$ and $cv$ produce outputs that differ only by a rescaling of $\alpha$. This is unavoidable; we consider $v$ and $cv$ as the same direction.
-Null space ambiguity
+Null space ambiguity.
 If $v_{0}\in\ker(J_{\ell})$ (i.e., $J_{\ell}v_{0}=0$), then adding $v_{0}$ to any steering vector does not change the linearized output.
@@
 We now state our main theoretical result, proving that persona vectors are fundamentally non-identifiable under standard observational regimes.
-Proposition 1
+Proposition 1.
 Under Assumptions A1–A3 within the local linear approximation, in Regime 2 (white-box single-layer access) without additional structural constraints, persona vectors are not identifiable.
 Specifically, for any steering vector $v\in\mathbb{R}^{d}$, there exist infinitely many vectors $v^{\prime}\not\propto v$ that are observationally equivalent under linearization.
-Proof Sketch
-We establish non-identifiability via the null-space ambiguity argument. A related weight-space symmetry result is given in (Appendix B.2).
+Proof Sketch.
+We establish non-identifiability via the null-space ambiguity argument. A related weight-space symmetry result is given in Appendix D.
 Null-space ambiguity (primary mechanism). Under the local linear approximation (Equation 4), the steered output is
@@
 For any $v_{0}\in\ker(J_{\ell})$, i.e., $J_{\ell}v_{0}=0$, define $v^{\prime}=v+v_{0}$. Then
-$o(x,v^{\prime},\alpha)$ $\approx o_{0}(x)+\alpha J_{\ell}(x)(v+v_{0})$ (8)
-$=o_{0}(x)+\alpha J_{\ell}(x)v\approx o(x,v,\alpha),$
+$o(x,v^{\prime},\alpha)\approx o_{0}(x)+\alpha J_{\ell}(x)(v+v_{0})=o_{0}(x)+\alpha J_{\ell}(x)v\approx o(x,v,\alpha).$ (8)
 Since $\dim(\ker(J_{\ell}))=d-\mathrm{rank}(J_{\ell})$, any case where $\mathrm{rank}(J_{\ell})<d$ implies a non-trivial null space.
@@
 Hence infinitely many such $v_{0}$ exist, producing observationally equivalent vectors $v^{\prime}\not\propto v$.
-Remark (Scope of linearization)
+Remark (Scope of linearization).
 Proposition 1 is proved under the local linear approximation (Equation 4), which holds exactly for piecewise-linear activations (ReLU MLPs on activation-region interiors) and as a first-order approximation for smooth nonlinearities (GeLU, attention softmax).
@@
 Extending the result to the exact nonlinear regime $o=F_{\ell\to L}(h_{\ell}+\alpha v)$ would require characterizing the Hessian structure of $F_{\ell\to L}$ and is left for future work.
-Corollary 1.1 (Null-space equivalence)
+Corollary 1.1 (Null-space equivalence).
 Under the linear approximation $o\approx o_{0}+\alpha J_{\ell}v$, any vector $v^{\prime}=v+v_{0}$ where $v_{0}\in\ker(J_{\ell})$ is observationally equivalent to $v$ since $J_{\ell}v^{\prime}=J_{\ell}v+J_{\ell}v_{0}=J_{\ell}v$.
 6 Empirical Validation
-We now validate that the non-identifiability characterized in Proposition 1 (Section 5) manifests in contemporary language models.
-Our empirical experiments test a behavioral consequence of the theoretical prediction: the existence of large equivalence classes of semantically indistinguishable steering directions, without directly estimating the Jacobian null space.
-A complementary logit-level analysis in Appendix A further examines output distribution preservation beyond the semantic probes used in the main experiments.
-6.1 Experimental Setup
-Models and Layers
-We evaluate two open-weight instruction-tuned language models to test generality across architectures and model scales: Qwen2.5-3B-Instruct (layer $\ell=12$) and Llama-3.1-8B-Instruct (layer $\ell=16$) (Yang et al., 2024; Meta AI, 2024).
-For both models, we focus on middle layers ($\ell=L/2$), following standard practice (Chen et al., 2025; Konen et al., 2024; Sun et al., 2025) in steering research.
-Persona traits
-We test three semantic traits spanning distinct dimensions: Formality, Politeness and Humor.
-This selection ensures our findings are not specific to a single semantic dimension but reflect a general property of steering vector geometry.
-Steering vector extraction
-For each trait, we construct 50 contrastive prompt pairs designed to elicit contrasting persona values.
-For example, formality pairs contrast "Write a professional and formal message about [topic]" versus "Write a casual and informal message about [topic]." For each prompt pair, we extract the hidden representation at layer $\ell$ for the final token position (Chen et al., 2025).
-The baseline steering vector is computed as the mean difference:
-$v=\frac{1}{50}\sum_{i=1}^{50}[h_{\ell}(x^{+}i)-h\ell(x^{-}_{i})]$ (9)
-Semantic probes
-We evaluate semantic equivalence using trait-specific scoring functions $\phi(o)$ that map generated text to real-valued scores.
-For formality, politeness and humor, we use lexical heuristics based on formal/informal markers, polite/rude markers and humorous/serious markers respectively, combined with sentence length and stylistic features.
-These return scores in $[0,1]$ where 1 is maximally formal/polite/humorous.
-Direct distributional evidence is provided by the logit-level analysis in Appendix A.
-Orthogonality test methodology
-To test Proposition 1’s prediction that $v$ and $v+v_{\perp}$ (where $v_{\perp}$ is orthogonal) produce observationally equivalent outputs, we implement the following procedure:
-  1. 1.
-    Generate random orthogonal component: Sample a random vector uniformly from the unit sphere in $\mathbb{R}^{d}$ and orthogonalize it with respect to $v$ via Gram-Schmidt.
-  2. 2.
-    Construct perturbed vector: Form $v^{\prime}=v+\alpha v_{\perp}$ where $\alpha$ is chosen such that $|v^{\prime}|\approx|v|$.
-  3. 3.
-    Generate steered outputs: For each of 100 held-out test prompts, generate text with steering vectors $v$ and $v^{\prime}$ at strength $\alpha=1.0$, producing 10 samples per prompt per vector.
-  4. 4.
-    Compute semantic equivalence: Measure Cohen’s $d$ effect size and Pearson correlation between semantic scores $\phi(o_{v})$ and $\phi(o_{v^{\prime}})$.
-  5. 5.
-    Repeat across seeds: Repeat for multiple random orthogonal seeds to assess robustness.
-If $v$ and $v^{\prime}$ are observationally equivalent as predicted by Proposition 1, we expect Cohen’s $d<0.2$ (negligible effect) and high correlation between semantic scores.
-Sample size design
-We conduct the orthogonality test with both $n=5$ and $n=10$ random orthogonal seeds per trait for Qwen2.5-3B and Llama-3.1-8B.
-This allows us to assess statistical stability across sample sizes and models.
-Scale invariance test
-To verify that observational equivalence holds across different steering strengths, we additionally evaluate the formality trait at four $\alpha$ values: 0.0,0.5,1.0,2.0{0.0,0.5,1.0,2.0} for both models.
-For each $\alpha$, we measure semantic scores under steering with $v$ versus $v+v_{\perp}$ and plot response curves.
-If equivalence is scale-invariant, the curves should track closely across all $\alpha$ with overlapping confidence bands.
-6.2 Orthogonal Perturbation Test Results
-Table 1 presents our empirical findings across two models, three semantic traits and two sample sizes ($n=5$ and $n=10$ random orthogonal seeds).
-Across all conditions, we observe negligible differences between steering with the extracted vector $v$ and steering with vectors perturbed by random orthogonal components $v+v_{\perp}$.
-We measure steering efficacy using semantic probe scores in $[0,1]$ that quantify the intensity of the target trait in generated outputs, comparing score distributions from steered generations against baseline generations.
-Table 1: Empirical validation of non-identifiability across models, traits, and sample sizes.
-Cohen’s $d$ measures the effect size between $v$ and $v+v_{\perp}$ (smaller = more equivalent).
-All results show negligible differences ($d<0.2$), supporting observational equivalence.
-Model Trait Seeds Cohen’s $d$ Correlation Perp-Only
-Qwen2.5-3B-Instruct Formality $n=5$ $0.144\pm 0.095$ $0.210\pm 0.113$ $98.8\%$
-$n=10$ $0.075\pm 0.058$ $0.285\pm 0.087$ $100.4\%$
-Politeness $n=5$ $0.112\pm 0.077$ $0.319\pm 0.070$ $101.5\%$
-$n=10$ $0.092\pm 0.069$ $0.414\pm 0.083$ $100.6\%$
-Humor $n=5$ $0.103\pm 0.077$ $0.276\pm 0.177$ $99.7\%$
-$n=10$ $072\pm 0.061 0.0$$044\pm 0.097$ $100.5\%$
-Llama-3.1-8B-Instruct Formality $n=5$ $0.052\pm 0.029$ $0.164\pm 0.044$ $95.6\%$
-$n=10$ $0.096\pm 0.068$ $0.192\pm 0.085$ $96.8\%$
-Politeness $n=5$ $0.074\pm 0.041$ $0.324\pm 0.058$ $101.2\%$
-$n=10$ $0.085\pm 0.043$ $0.347\pm 0.077$ $100.4\%$
-Humor $n=5$ $159\pm 0.109 0.0$$032\pm 0.116$ $98.4\%$
-$n=10$ $119\pm 0.119 0.0$$016\pm 0.104$ $95.9\%$
-With $n=10$ seeds, the mean Cohen’s $d$ is 0.080 for Qwen2.5-3B and 0.100 for Llama-3.1-8B, both well below the threshold for small effects ($d=0.2$) and firmly in the negligible range.
-Near-zero correlations (e.g., $0.044$, 0.032-0.032) reflect probe ceiling effects rather than independence; Cohen’s $d$ is the more reliable metric in these cases.
-The “Perp-Only Effect” column measures the efficacy of steering with pure orthogonal components ($v_{\perp}$ alone, without $v$) relative to the extracted vector—values near $100\%$ indicate that orthogonal components achieve equivalent behavioral impact.
-Values marginally above $100\%$ reflect probe ceiling effects at the upper boundary of the scoring range.
-Statistical stability across sample sizes The convergence between $n=5$ and $n=10$ results demonstrates robustness across both models.
-For Qwen2.5-3B, effect sizes change by <0.07<0.07 as sample size increases, with all traits showing tightening confidence intervals: formality ($0.144\to 0.075$), politeness ($0.112\to 0.092$), humor ($0.103\to 0.072$).
-For Llama-3.1-8B, we observe similar stability: formality ($0.052\to 0.096$), politeness ($0.074\to 0.085$), humor ($0.159\to 0.119$).
-This consistency demonstrates that the observed equivalence is not a sampling artifact but a stable property of the steering geometry.
-Cross-model consistency The close agreement between Qwen2.5-3B ($d=0.080$) and Llama-3.1-8B ($d=0.100$) demonstrates that observational equivalence is not model-specific.
-Despite differences in architecture, scale (3B vs.
-8B parameters) and hidden dimension ($d=2048$ vs.
-$d=4096$), both models exhibit nearly identical patterns of non-identifiability.
-The consistency across traits within each model further confirms that the phenomenon is general rather than an artifact of specific semantic domains or model implementations.
-Visualizing orthogonal component equivalence Figure 2 illustrates the perp-only effect ratios across all traits and models using $n=10$ orthogonal seeds.
-The tight clustering around the perfect equivalence line (dashed red at $1.0$) demonstrates that pure orthogonal components achieve nearly identical steering efficacy to the extracted vectors.
-(a) Qwen2.5-3B-Instruct
-(b) Llama-3.1-8B-Instruct
-Figure 2: Perp-only effect ratios ($v_{\perp}$ efficacy / $v$ efficacy) for $n=10$ orthogonal seeds. Values near 1.0 indicate equivalence.
-[…]
-6.3 Scale Invariance Analysis
-To verify that observational equivalence holds across different steering strengths, we evaluate the formality trait at four steering magnitudes $\alpha\in\{0.0,\,0.5,\,1.0,\,2.0\}$ for both models.
-Figure 3 shows the response curves for the extracted vector $v$ and the observationally equivalent vector $v+v_{\perp}$.
-(a) Qwen2.5-3B-Instruct
-(b) Llama-3.1-8B-Instruct
-Figure 3: Scale invariance of observational equivalence.
-Formality scores across steering strengths $\alpha\in\{0.0,\,0.5,\,1.0,\,2.0\}$ for the extracted vector $v$ (blue circles) and the perturbed vector $v+v_{\perp}$ (orange squares).
-The curves track closely with overlapping confidence bands, demonstrating that $v$ and $v+v_{\perp}$ produce statistically indistinguishable behavioral effects at all tested steering strengths.
-For Qwen2.5-3B, mean differences remain below $0.022$ across all $\alpha$ values (less than $2.5\%$ deviation in formality scores); for Llama-3.1-8B, differences remain below $0.023$ ($2.5\%$ deviation).
-This proves that non-identifiability is a structural property: any vector in the equivalence class $v+\ker(J)$ produces identical observations regardless of scaling.
-6.4 Robustness of Non-Identifiability Under Distribution Shift
-We test whether equivalence classes persist across distributional shifts or whether prompt diversity alone induces identifiability.
-We evaluate orthogonal perturbation equivalence across four distinct prompt environments: in-distribution (standard trait templates), topic shift (medical, legal, creative writing contexts), genre shift (emails, tweets, academic writing) and safety-style (refusal patterns, ethical dilemmas).
-For each environment, we extract steering vectors and evaluate whether $v\equiv v+v_{\perp}$ equivalence holds when testing on that environment’s prompts.
-Table 2 presents Cohen’s $d$ effect sizes across all trait-environment combinations for both models.
-Across 24 conditions, we observe effect sizes ranging from $d=0.21$ to $d=0.63$ (mean $d=0.35$).
-Effect sizes are predominantly small ($d<0.5$ for 88% of conditions; mean $d=0.35$), indicating that $v$ and $v+v_{\perp}$ remain largely observationally equivalent across diverse extraction contexts.
-Table 2: Multi-environment validation. Cohen’s $d$ effect sizes across prompt contexts (in-dis: in-distribution, top: topic shift, gen: genre shift, saf: safety-style).
-Model Trait In-dis Top Gen Saf
-Qwen-3B Formality 0.44 0.25 0.29 0.21
-Politeness 0.22 0.24 0.52 0.41
-Humor 0.42 0.51 0.24 0.51
-Llama-8B Formality 0.30 0.54 0.30 0.21
-Politeness 0.37 0.27 0.51 0.63
-Humor 0.29 0.56 0.30 0.40
-Formality shows most consistency (mean $d=0.32$), with particularly strong equivalence in safety-style contexts ($d=0.21$ for both models).
-Politeness exhibits higher variability ($d$ reaching 0.52–0.63 in genre-shift and safety contexts) but remains within the small-effect range.
-Humor demonstrates intermediate stability (mean $d=0.39$), with genre-shift producing notably lower effects ($d=0.24$ for Qwen, $d=0.30$ for Llama).
-Figure 4 visualizes these patterns through environment-dependent effect sizes, showing that equivalence remains stable even when vectors are extracted in one environment and tested in another.
-Figure 4: Environment-dependent effect sizes across prompt contexts.
-Implications
-The persistence of small effect sizes across diverse prompt distributions indicates that non-identifiability is not an artifact of in-distribution prompts but a structural geometric property.
-Cases where $d$ approaches 0.50.50.60.6 (e.g., politeness under genre or safety-style shifts) suggest that certain trait–context combinations impose partial constraints without eliminating null-space ambiguity since $\ker(J)$ is determined by model weights rather than the prompt distribution.
-Although effect sizes increase under distribution shift (reaching $d=0.63$ in some settings), distributional diversity alone does not eliminate observational equivalence.
+We now validate that the non-identifiability characterized in Proposition 1 manifests in large language models.
+Our experiments test both the theoretical prediction directly — via null-space dimensionality estimation and orthogonal perturbation tests — and its broader consequences across extraction methods, steering strengths and prompt distributions.
+A complementary logit-level analysis in Appendix B further examines output distribution preservation beyond the semantic probes used here.
+6.1 Setup
+We evaluate on Llama-3.1-8B-Instruct and Qwen2.5-3B-Instruct [^2] For brevity, we refer to these models as Llama and Qwen respectively throughout.
+(Yang et al., 2024; Meta AI, 2024) representing two architecture families.
+The steering interventions are applied at the middle layer of each model (layer 16/32 for Llama, layer 18/36 for Qwen) following standard practice (Chen et al., 2025).
+We study five semantic traits: formality, politeness, sentiment, truthfulness and agreeableness which spans surface-level and cognitive attributes.
+The trait presence in generated outputs is measured using pre-trained classifiers: formality with s-nlp/roberta-base-formality-ranker, sentiment with cardiffnlp/twitter-roberta-base-sentiment-latest and politeness, truthfulness and agreeableness with zero-shot classification via facebook/bart-large-mnli, all normalized to $[0,1]$.
+For each trait, 50 contrastive prompt pairs $(x^{+},x^{-})$ are constructed to contrast high and low trait expression, and the steering vector is computed as the mean activation difference:
+$v=\frac{1}{50}\sum_{i=1}^{50}\bigl[h_{\ell}(x^{+}_{i})-h_{\ell}(x^{-}_{i})\bigr]$ (9)
+For the vector equivalence experiment (Section 6.3), a second vector is extracted via PCA of the activation covariance matrix over the same pairs.
+We report Cohen’s $d$ as the primary metric throughout, using the threshold $|d|<0.2$ (negligible effect) to indicate behavioral equivalence (Cohen, 2013).
+6.2 Orthogonal Perturbation Test
+We test whether adding a random direction orthogonal to $v$ changes steering behavior.
+For each trait and model, we generate $n$ random unit vectors $v_{\perp}$ orthogonal to $v$ via Gram-Schmidt and evaluate steering with $v+v_{\perp}$ versus $v$ alone across 100 held-out test prompts.
+Results are shown in Table 1.
+Across all traits and both models, mean $|d|<0.2$ for both $n=5$ and $n=10$ seeds.
+With $n=10$ seeds, mean Cohen’s $d$ is 0.119 for Llama and 0.131 for Qwen, well below the negligible-effect threshold.
+The effect sizes converge between $n=5$ and $n=10$ for all trait-model pairs, with confidence intervals tightening as sample size increases, indicating the observed equivalence is not a sampling artifact.
+Despite differences in architecture and scale, both models exhibit nearly identical patterns of non-identifiability, confirming the result is not model-specific.
+Table 1: Orthogonal perturbation test. Mean Cohen’s $d$ (std) comparing $v$ and $v+v_{\perp}$ across $n$ seeds. All values below the equivalence threshold $|d|=0.2$.
+Trait Model $n=5$ $n=10$
+Mean Std Mean Std
+Formality Llama 0.120 0.037 0.059 0.056
+Qwen 0.052 0.048 0.151 0.105
+Politeness Llama 0.171 0.062 0.159 0.113
+Qwen 0.065 0.032 0.120 0.091
+Sentiment Llama 0.083 0.049 0.104 0.040
+Qwen 0.206 0.141 0.150 0.079
+Truthfulness Llama 0.045 0.027 0.085 0.070
+Qwen 0.081 0.076 0.092 0.087
+Agreeableness Llama 0.159 0.059 0.186 0.067
+Qwen 0.114 0.062 0.156 0.046
+6.3 Vector Equivalence Across Extraction Methods
+We test whether non-identifiability extends beyond orthogonal perturbations to vectors extracted by different methods.
+We compare $v_{1}$ (contrastive mean difference) and $v_{2}$ (first principal component of activation variance) for each trait and model.
+These two methods are theoretically distinct: the contrastive method targets the mean activation shift across contrasting prompts while PCA targets the direction of maximum variance.
+If they produce equivalent behavioral effects despite geometric distinctness, this is direct evidence that the equivalence class is large enough to contain vectors from independent extraction pipelines.
+Figure 2 shows the geometric-semantic decoupling: cosine similarities range from 0.54-0.54 to +0.32+0.32, indicating that $v_{1}$ and $v_{2}$ are often near-orthogonal or even anti-aligned.
+Despite this, Cohen’s $d$ remains below $0.27$ in all conditions.
+The Llama sentiment pair ($\cos=-0.54$, $|d|=0.19$) is the most striking: two vectors pointing in nearly opposite directions produce statistically equivalent behavioral effects, confirming that the extraction method does not determine a unique steering direction.
+Full numerical results are provided in Table 2.
+Figure 2: Geometric-semantic decoupling.
+Despite large variation in cosine similarity between $v_{1}$ and $v_{2}$, almost all points fall below the equivalence threshold $|d|=0.2$, proving that geometric diversity does not translate to behavioral difference.
+Trait Model $\cos(v_{1},v_{2})$ Cohen’s $d$
+Agreeableness Qwen +0.239+0.239 0.024-0.024
+Llama 0.333-0.333 0.073-0.073
+Formality Qwen +0.144+0.144 +0.141+0.141
+Llama +0.198+0.198 +0.230+0.230
+Politeness Qwen +0.267+0.267 +0.252+0.252
+Llama +0.317+0.317 +0.088+0.088
+Sentiment Qwen +0.205+0.205 +0.005+0.005
+Llama 0.539-0.539 +0.191+0.191
+Truthfulness Qwen 0.321-0.321 0.096-0.096
+Llama +0.084+0.084 0.112-0.112
+Table 2: Vector equivalence results. $v_{1}$ = contrastive and $v_{2}$ = PCA. All Cohen’s $d<0.27$.
+6.4 Scale Invariance
+Proposition 1 predicts that non-identifiability holds throughout the regime where the linear approximation is valid.
+We test this by sweeping steering strength $\alpha\in\{0.0,0.5,1.0,2.0,3.0\}$ and measuring Cohen’s $d$ between $v$ and $v+v_{\perp}$ at each value, across 10 random seeds.
+Figure 3 shows results for all five traits and both models.
+Figure 3: Semantic equivalence across operational steering range ($\alpha\in[0.0,3.0]$).
+Blue ($v$) and orange ($v+v_{\perp}$) trajectories track classifier scores, with shaded confidence bands.
+Near-overlapping curves confirm non-identifiability throughout standard steering regime.
+The semantic scores for $v$ and $v+v_{\perp}$ track closely throughout $\alpha\leq 3$, with overlapping confidence bands in nearly all conditions.
+Sentiment exhibits wider variance due to the binary nature of its classifier output but remains well below the equivalence threshold on average.
+6.5 Null-Space Spanning
+The orthogonal perturbation test uses a single random direction per seed.
+A stronger test is whether equivalence holds across the full null space: if 50 independently sampled orthogonal directions all produce small effects, the equivalence class is high-dimensional rather than confined to a thin region around $v$.
+We sample 50 random unit vectors orthogonal to $v$ via Gram-Schmidt and measure Cohen’s $d$ for each direction individually at $\alpha=1.0$.
+Figure 4 and Table 3 show results for all traits and both models.
+Most traits cluster near $|d|\approx 0.08$$0.13$, with nearly all 50 directions per trait falling below the equivalence threshold ($|d|=0.2$).
+Sentiment and politeness show slightly elevated responses ($|d|\approx 0.24$$0.29$), though these remain well within the small-effect range and do not indicate a breakdown of equivalence.
+The key result is that equivalence is not isolated to a few directions: it is the dominant behavior across the null space, confirming that the equivalence class is genuinely high-dimensional.
+Figure 4: Cohen’s $d$ for 50 random orthogonal directions per trait across both models (Qwen: blue, Llama: red).
+Trait Model Mean $|d|$
+Agreeableness Qwen 0.121
+Llama 0.080
+Formality Qwen 0.119
+Llama 0.124
+Politeness Qwen 0.095
+Llama 0.243
+Sentiment Qwen 0.291
+Llama 0.265
+Truthfulness Qwen 0.185
+Llama 0.108
+Table 3: Null-space spanning results. Mean Cohen’s $d$ across 50 random orthogonal directions per trait at $\alpha=1.0$.
+6.6 Multi-Environment Validation
+To confirm that non-identifiability is not an artifact of limited prompt diversity, we test whether equivalence persists across distribution shift.
+We extract a steering vector from in-distribution prompts and evaluate on three environments: in-distribution, topic-shift, and genre-shift.
+Cohen’s $d$ is measured between $v$-steered and $(v+v_{\perp})$-steered outputs in each environment.
+The results are shown in Figure 5.
+Figure 5: Cohen’s $d$ for $v$ vs. $v+v_{\perp}$ across the three prompt environments. Color gradient represents effect magnitude from small to large across traits and models.
+Effect sizes remain moderate across all environments and traits, with most conditions in the small-effect range.
+No catastrophic breakdown occurs under topic or genre shift, confirming that the equivalence class structure is a stable geometric property of the model rather than a prompt-specific artifact.
+The elevated values for truthfulness and agreeableness under in-distribution steering ($d\approx 0.7$) reflect higher classifier variance for these cognitively complex traits rather than a genuine breakdown of equivalence, consistent with findings that abstract semantic properties are harder to probe reliably via linear classifiers (Elazar et al., 2021; Marks and Tegmark, 2023).
+7 Discussion
+Our results establish that steering vectors are non-identifiable in a strong empirical sense: orthogonal perturbations, alternative extraction methods and diverse prompt environments all fail to distinguish between behaviorally equivalent directions.
+Claims that a steering vector represents a specific semantic concept are therefore not warranted by behavioral evidence alone.
+Since many geometrically distinct vectors produce identical outputs, the observed behavioral effect of steering does not localize semantic content to a specific direction in activation space.
+We note that this does not mean steering is ineffective: the experiments confirm that steering vectors reliably shift model behavior in the intended direction.
+The limitation is representational, not functional.
 8 Limitations
-Our empirical validation covers two models at mid-network layers across three semantic traits (formality, politeness, humor).
+Our empirical validation covers two models at mid-network layers across five semantic traits.
 While our theoretical results apply generally, the empirical magnitude of non-identifiability may vary across model families, scales, architectures and layer positions.
-Our semantic evaluation uses lexical heuristics rather than full distributional metrics or human evaluation, providing a conservative test of observational equivalence.
-Our multi-environment experiments demonstrate that equivalence persists across diverse prompt distributions but we do not test whether explicit structural interventions such as independence constraints (ICA), sparsity regularization or invariance objectives (IRM) can recover identifiable representations.
-Empirical validation of these approaches remains an important direction for future work.
-Additionally, we do not directly measure null-space dimensionality or test whether null-space versus row-space components exhibit differential robustness under distribution shift.
-8 Future Work
-Our results establish the necessity of validation beyond input-output behavior.
-Important directions include testing whether structural constraints can recover identifiable representations, analyzing how non-identifiability varies across network layers and architectural components and characterizing how null-space dimensionality scales with model size.
-Investigating whether alternative steering methods exhibit superior robustness under adversarial attacks or distribution shift would clarify the practical consequences of non-identifiability.
-Human evaluation would provide stronger validation than our lexical heuristics and direct measurement of null-space dimensionality would quantify our theoretical predictions.
+The multi-environment experiments demonstrate that equivalence persists across diverse prompt distributions but do not test whether explicit structural interventions such as independence constraints (Hyvärinen and Pajunen, 1999), sparsity regularization or invariance objectives (Arjovsky et al., 2019) can recover identifiable representations.
+We also do not directly verify that the semantic null-space estimated via activation SVD coincides with the Jacobian null-space assumed in Proposition 1; this connection is supported by the behavioral results but not formally proven.
 9 Conclusion
-We have established that steering vectors are fundamentally non-identifiable under the local linear approximation, with infinitely many distinct directions producing equivalent behavioral effects.
-This non-identifiability persists across models, traits and prompt distributions, revealing a fundamental tension: methods optimized for behavioral control may fail to yield interpretable representations.
-Our work clarifies when steering provides reliable heuristic modification versus principled causal understanding, highlighting that interpretability claims require validation beyond input-output equivalence alone.
-More broadly, when internal interventions admit large equivalence classes under the model’s local geometry, consistent output shifts do not imply that a discovered direction corresponds to a canonical or causally privileged feature.
-Stronger interpretability claims therefore require structural constraints or additional assumptions that break these symmetries and distinguish meaningful factors from geometrically arbitrary ones.
+We have shown that steering vectors are fundamentally non-identifiable under white-box single-layer access: infinitely many geometrically distinct directions produce equivalent behavioral effects due to the large null space of the activation-to-output Jacobian.
+This non-identifiability persists across models, traits, extraction methods and prompt distributions.
+The implication is that behavioral equivalence is not a sufficient basis for interpretability claims about internal representations, and that alignment methods relying on steering vectors require structural constraints beyond behavioral testing to be trusted.
+More broadly, any intervention-based interpretability method that validates representational claims solely through output behavior faces the same fundamental limitation.
 Acknowledgments
-We thank Vast.ai for providing GPU compute resources that enabled our empirical validation experiments.
-We also acknowledge the HuggingFace platform and maintainers of the open-source models used in this work.
-Generative AI tools were used for language editing and code assistance; all technical content was reviewed and verified by the authors.
-Appendix A Logit-Level Equivalence Test
-To complement our semantic evaluations, we perform a direct logit-level equivalence test.
-For each model and trait, we compare the next-token logit vectors produced by steering with $v$ and with $v+v_{\perp}$, where $v_{\perp}$ is a norm-matched direction orthogonal to $v$.
-We measure the $\ell_{2}$ distance between logit vectors and normalize by the corresponding distance induced by a norm-matched random direction:
+We thank Vast.ai for GPU compute resources. We acknowledge the HuggingFace platform and maintainers of the open-source models used in this work.
+Ethics Statement
+This work studies the theoretical and empirical foundations of activation steering in large language models.
+We do not introduce new steering methods or systems that could enable misuse.
+Our findings only highlight limitations of existing methods.
+We use publicly available models and standard evaluation datasets throughout and have adhered to their respective terms of use.
+Disclosure of AI Usage
+Large language models were used for language editing of the draft and code assistance in this work.
+All research ideas, technical content, experimental design, results and conclusions are the original work of the authors.
+Appendix A Null-Space Dimensionality
+We estimate the dimensionality of the semantically inert subspace at three layers (L/4, L/2, 3L/4) of each model.
+Hidden activations are collected over 200 in-distribution prompts; we compute the activation covariance matrix and apply SVD.
+Effective rank is defined as the number of singular values exceeding $\varepsilon=0.01$.
+The null-space fraction is $(d-\text{rank})/d$.
+Table 4: Null-space dimensionality across three layers for both models.
+Model Layer Metrics
+$d$ Eff. rank NF
+Llama L/4 4096 57 0.943
+L/2 4096 59 0.941
+3L/4 4096 63 0.937
+Qwen L/4 2048 139 0.861
+L/2 2048 127 0.873
+3L/4 2048 116 0.884
+The singular value spectra at three layers are shown in Figures 6(a) and 6(b).
+The surface drops sharply after the first few dimensions, with singular values collapsing to near zero beyond dimension 50–140 depending on the model.
+This cliff-like decay persists consistently across all three layers, indicating that the low-rank structure is not a property of any particular depth but a stable geometric feature of both architectures.
+The vast flat region beyond the initial peak represents the semantically inert null space: directions in this region produce no measurable change in output logits and are therefore undetectable by any behavioral evaluation.
+For Qwen, the decay is steeper and the plateau lower which is consistent with its smaller hidden dimension ($d=2048$) relative to Llama ($d=4096$).
+Figure 6: Singular value spectra across layers.
+(a) Qwen and (b) Llama both exhibit a sharp cliff after the first 50–140 dimensions followed by a near-zero plateau, confirming that the vast majority of the activation space is semantically inert across all layers.
+Appendix B Logit-Level Equivalence Test
+To confirm that equivalence is not an artifact of semantic classifier choice, we measure raw logit distances.
+For each model and trait, we compute the $\ell_{2}$ distance between the next-token logit vectors produced under three conditions: (i) steering with $v$, (ii) steering with $v+v_{\perp}$ where $v_{\perp}$ is a norm-matched orthogonal direction, and (iii) steering with a norm-matched random direction $v_{\text{rand}}$.
+The ratio
 $R=\frac{\|\ell(v)-\ell(v+v_{\perp})\|_{2}}{\|\ell(v)-\ell(v_{\text{rand}})\|_{2}}$ (10)
-Values $R<1$ indicate that orthogonal perturbations induce smaller logit deviations than random directions.
-In all tested models and traits, this condition holds.
-Ratios range from $0.49$ (Llama, formality) to $0.90$ (Qwen, humor), indicating that orthogonal perturbations systematically induce smaller logit shifts than random directions.
-The next-token predictions remain highly stable: token agreement between $v$ and $v+v_{\perp}$ ranges from $81\%$ to $96\%$ and top-10 token overlap ranges from $86\%$ to $96\%$.
-These results provide direct evidence that equivalence is not an artifact of semantic measurement.
-While orthogonal perturbations induce some logit deviation, they systematically preserve output distributions more faithfully than random directions, consistent with Proposition 1’s prediction of high-dimensional equivalence classes.
-Table 3 reports mean $\ell_{2}$ logit distances, normalized ratios, token agreement and top-$k$ overlap for all models and traits.
-Table 3: Logit-level equivalence statistics. $R$ denotes the ratio $v_{\perp}$ / random. Agreement% refers to token agreement, and Overlap% indicates top-10 overlap.
-Model Trait $R$ Agreement% Overlap%
-Qwen-3B Humor 0.90 81% 86%
-Formality 0.76 94% 96%
-Politeness 0.75 93% 96%
-Llama-8B Humor 0.64 96% 90%
-Formality 0.49 93% 95%
-Politeness 0.65 91% 94%
-Appendix B Detailed Proof of Proposition 1
-B.1 Null-Space Ambiguity (Constructive Proof)
-Setup Consider the local linear approximation of the steering effect:
+quantifies how much smaller the logit shift from an orthogonal perturbation is compared to a random one.
+Values $R<1$ indicate that orthogonal perturbations are systematically less disruptive than random directions.
+Table 5: Logit-level equivalence.
+Baseline = $\|\ell(0)-\ell(v)\|_{2}$; $v_{\perp}$ and Random denote $\|\ell(v)-\ell(v+v_{\perp})\|_{2}$ and $\|\ell(v)-\ell(v_{\text{rand}})\|_{2}$ respectively.
+$R<1$ in all conditions.
+Model Trait Metrics
+Baseline $v_{\perp}$ Random $R$
+Qwen Formality 43.73 40.93 57.78 0.71
+Sentiment 67.56 62.79 86.21 0.73
+Truthfulness 113.06 118.24 165.91 0.71
+Politeness 53.14 52.95 73.07 0.72
+Agreeableness 75.66 64.64 96.92 0.67
+Llama Formality 119.12 64.36 136.58 0.47
+Sentiment 87.67 61.84 100.94 0.61
+Truthfulness 206.04 121.33 226.73 0.54
+Politeness 116.88 98.70 144.28 0.68
+Agreeableness 119.98 85.01 128.02 0.66
+$R$ ranges from $0.47$ to $0.73$ across all conditions, as reported in Table 5.
+Orthogonal perturbations induce logit shifts that are 27–53% smaller than random directions of the same norm.
+In several Llama conditions, the orthogonal perturbation logit distance is also smaller than the baseline distance, indicating that $v_{\perp}$ moves the model less than $v$ itself does.
+These results confirm that the equivalence observed in semantic scores is not a property of the classifiers: it is present directly in the model’s output distribution.
+The full bar-chart comparisons are shown in Figures 7(a) and 7(b).
+Figure 7: Logit $\ell_{2}$ distances for (a) Qwen and (b) Llama.
+Each group shows three bars per trait: Baseline ($\|\ell(0)-\ell(v)\|_{2}$), $v_{\perp}$ ($\|\ell(v)-\ell(v+v_{\perp})\|_{2}$) and Random ($\|\ell(v)-\ell(v_{\text{rand}})\|_{2}$).
+The $v_{\perp}$ bar is consistently shorter than Random across all traits which confirms that orthogonal perturbations are less disruptive to the output distribution than unstructured directions of equal norm.
+Appendix C Detailed Proof of Proposition 1
+We provide a constructive proof of Proposition 1, formalizing the null-space ambiguity argument presented in Section 5.
+Setup.
+Consider the local linear approximation of the steering effect at layer $\ell$:
 $o(x,v,\alpha)\approx o(x,0,0)+\alpha J_{\ell}(x)v$ (11)
-where $J_{\ell}(x)=\frac{\partial o}{\partial h_{\ell}}\big|_{h_{\ell}(x)}\in\mathbb{R}^{V\times d}$ is the Jacobian.
-Step 1: Null space characterization
+where $J_{\ell}(x)=\frac{\partial o}{\partial h_{\ell}}\big|_{h_{\ell}(x)}\in\mathbb{R}^{V\times d}$ is the Jacobian of the output logits with respect to the layer-$\ell$ activations, $V$ is the vocabulary size, and $d$ is the hidden dimension.
+C.1 Null-space characterization
 Define the null space of $J_{\ell}$ as:
@@
 $\dim(\mathcal{N})=d-\operatorname{rank}(J_{\ell})$ (13)
-Step 2: Rank Bound
-The Jacobian $J_{\ell}\in\mathbb{R}^{V\times d}$ has maximum possible rank:
-$\text{rank}(J_{\ell})\leq\min(V,d)$ (14)
-In modern language models, the hidden dimension is typically $d\approx 4000$ while the vocabulary size is $V\approx 50000$. Therefore, the Jacobian satisfies $\max\mathrm{rank}(J_{\ell})=d$.
-Step 3: Effective rank is much lower
-Output distributions lie on a low-dimensional manifold. The effective rank satisfies:
-$\text{rank}_{\epsilon}(J_{\ell})=\#\{\sigma_{i}:\sigma_{i}>\epsilon\cdot\sigma_{\max}\}\ll d$ (15)
-where $\sigma_{i}$ are singular values of $J_{\ell}$ and $\epsilon$ is a threshold (e.g., $10^{-4}$).
-Intuition: Empirical evidence suggests that in overparameterized LLMs the effective rank of $J_{\ell}$ may be substantially lower than $d$, consistent with low dimensional structure observed in related weight space studies (Li et al., 2018; Maennel et al., 2018) and concentration of output sensitivity in a limited number of directions (Ghorbani et al., 2019).
-Consequently, $\dim(\mathcal{N})=d-\text{rank}(J_{\ell})$ is often positive in practice.
-However, this empirical observation is not required for the proof; the argument establishes non identifiability whenever $\dim(\mathcal{N})\geq 1$.
-Step 4: Constructing equivalent vectors
-For any steering vector $v\in\mathbb{R}^{d}$ and any $v_{0}\in\mathcal{N}$, define
-$v^{\prime}=v+v_{0}.$ (16)
-Then for all $x$ and all $\alpha$:
-$J_{\ell}(x)v^{\prime}$ $=J_{\ell}(x)(v+v_{0})$ (17)
-$=J_{\ell}(x)v+J_{\ell}(x)v_{0}$
-$=J_{\ell}(x)v.$
-Therefore:
-$o(x,v^{\prime},\alpha)$ $\approx o(x,0,0)+\alpha J_{\ell}(x)v^{\prime}$ (18)
-$=o(x,0,0)+\alpha J_{\ell}(x)v$
-$\approx o(x,v,\alpha).$
-Step 5: Infinitely many distinct solutions
-Since $\dim(\mathcal{N})\geq 1$, the null space contains infinitely many directions. For any $v_{0}\in\mathcal{N}\setminus\{0\}$ and any $\beta\in\mathbb{R}$:
+C.2 Rank bound and effective rank
+The Jacobian $J_{\ell}\in\mathbb{R}^{V\times d}$ satisfies $\operatorname{rank}(J_{\ell})\leq\min(V,d)$.
+In modern language models $V\approx 50{,}000$ and $d\approx 2{,}000$8,0008{,}000, so the theoretical maximum rank is $d$.
+However, empirical evidence indicates that output distributions lie on a low-dimensional manifold.
+The effective rank, defined as
+$\operatorname{rank}_{\varepsilon}(J_{\ell})=\#\{\sigma_{i}:\sigma_{i}>\varepsilon\cdot\sigma_{\max}\}$ (14)
+where $\sigma_{i}$ are singular values of $J_{\ell}$ and $\varepsilon$ is a threshold (e.g., $10^{-2}$), is substantially lower than $d$ in practice.
+Our measurements (Appendix A) yield effective ranks of 57 and 139 for Llama and Qwen respectively, leaving null-space fractions of 94% and 86%.
+The constructive argument below requires only $\dim(\mathcal{N})\geq 1$.
+C.3 Constructing observationally equivalent vectors
+For any steering vector $v\in\mathbb{R}^{d}$ and any $v_{0}\in\mathcal{N}$, define $v^{\prime}=v+v_{0}$. Then for all prompts $x$ and all steering strengths $\alpha$:
+$J_{\ell}(x)v^{\prime}=J_{\ell}(x)(v+v_{0})=J_{\ell}(x)v+J_{\ell}(x)v_{0}=J_{\ell}(x)v$ (15)
+where the last step uses $v_{0}\in\ker(J_{\ell})$. Therefore:
+$o(x,v^{\prime},\alpha)\approx o(x,0,0)+\alpha J_{\ell}(x)v^{\prime}=o(x,0,0)+\alpha J_{\ell}(x)v\approx o(x,v,\alpha)$ (16)
+C.4 Infinitely many distinct solutions
+Since $\dim(\mathcal{N})\geq 1$, the null space contains infinitely many directions. For any $v_{0}\in\mathcal{N}\setminus\{0\}$ and any $\beta\in\mathbb{R}$, the family
 $v^{\prime}_{\beta}=v+\beta v_{0}$ (17)
 generates infinitely many observationally equivalent vectors.
-Furthermore, if $\beta$ is chosen such that $v^{\prime}_{\beta}\not\propto v$ (which is always possible unless $v\propto v_{0}$), these vectors are geometrically distinct.
-Step 6: Non-proportionality
-To ensure $v^{\prime}_{\beta}\not\propto v$, we need $v+\beta v_{0}\neq cv$ for any scalar $c$. This fails only if:
-$\beta v_{0}=(c-1)v$ (20)
-which requires $v\in\text{span}(v_{0})$.
-Since $v_{0}$ is an arbitrary element of a $\dim(\mathcal{N})$-dimensional space and $v$ is fixed, this occurs with probability zero.
-Therefore, for generic $v$ and generic $v_{0}\in\mathcal{N}$, we have $v^{\prime}_{\beta}\not\propto v$ for almost all $\beta$.
-Conclusion (Null-space mechanism)
-Under the local linear approximation and for typical Jacobian structure, there exist infinitely many geometrically distinct steering vectors that are observationally equivalent.
-B.2 Weight Space Gauge Symmetry (Supplementary Observation)
-While Proposition 1 fixes model weights and asks whether v is recoverable from behavior, this section offers a complementary perspective: even relaxing the weight-fixing assumption, non-identifiability persists via linear reparameterization symmetries.
-Setup: We record a supplementary observation about gauge symmetry in neural network parameterization.
-This is distinct from the null-space non-identifiability of Proposition 1: rather than fixing model weights and asking whether $v$ is recoverable from behavioral observations, we ask whether different $(W,v)$ and $(W^{\prime},v^{\prime})$ pairs can produce identical input–output behavior.
-The answer is yes, via the linear reparameterization below.
-This symmetry is not realizable as a steering operation on a frozen model as it requires modifying $W_{\ell+1}$.
-Step 1: Representation reparameterization
-Consider an invertible transformation $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$. Define the reparameterized representation:
-$h^{\prime}_{\ell}(x)=T\bigl(h_{\ell}(x)\bigr).$ (21)
-If the subsequent layers can be rewritten as:
-$F_{\ell\to L}(h_{\ell})=F^{\prime}_{\ell\to L}(T(h_{\ell}))$ (22)
-then $(h_{\ell},v)$ and $(h^{\prime}_{\ell},v^{\prime})$ where $v^{\prime}=DT(h_{\ell})\cdot v$ are observationally indistinguishable.
-Step 2: Linear reparameterizations
-Consider linear transformations $T(h)=Ah$ where $A\in\mathbb{R}^{d\times d}$ is invertible. For layer $\ell+1$ with weight matrix $W_{\ell+1}\in\mathbb{R}^{d\times d}$ and bias $b_{\ell+1}$:
-Original computation:
-$h_{\ell+1}=\sigma(W_{\ell+1}h_{\ell}+b_{\ell+1})$ (23)
-Reparameterized computation:
-$h^{\prime}_{\ell+1}=\sigma(W^{\prime}_{\ell+1}h^{\prime}_{\ell}+b^{\prime}_{\ell+1})$ (24)
-where $W^{\prime}_{\ell+1}=W_{\ell+1}A^{-1}$ and $b^{\prime}_{\ell+1}=b_{\ell+1}$.
-Note: The bias remains invariant under linear reparameterization through the origin.
-For more general affine transformations $T(h)=Ah+c$ with translation $c\neq 0$, the bias would also transform as $b^{\prime}_{\ell+1}=b_{\ell+1}-W_{\ell+1}A^{-1}c$.
-We restrict to origin-preserving transformations for simplicity, as these suffice to establish non-identifiability.
-Then:
-$h^{\prime}_{\ell+1}$ $=\sigma\!\left(W_{\ell+1}A^{-1}Ah_{\ell}+b_{\ell+1}\right)$ (25)
-$=\sigma\!\left(W_{\ell+1}h_{\ell}+b_{\ell+1}\right)$
-$=h_{\ell+1}.$
-Step 3: Steering under reparameterization
-Original steering:
-$\tilde{h}_{\ell}=h_{\ell}+\alpha v$ (26)
-$\tilde{h}_{\ell+1}=\sigma(W_{\ell+1}(h_{\ell}+\alpha v)+b_{\ell+1})$ (27)
-Reparameterized steering with $v^{\prime}=Av$:
-$\tilde{h}^{\prime}_{\ell}$ $=h^{\prime}_{\ell}+\alpha v^{\prime}$ (28)
-$=Ah_{\ell}+\alpha Av$
-$=A(h_{\ell}+\alpha v)$
-$=A\tilde{h}_{\ell}$.
-$\tilde{h}^{\prime}_{\ell+1}$ $=\sigma\!\left(W^{\prime}_{\ell+1}(h^{\prime}_{\ell}+\alpha v^{\prime})+b^{\prime}_{\ell+1}\right)$ (29)
-$=\sigma\!\left(W_{\ell+1}A^{-1}A(h_{\ell}+\alpha v)+b_{\ell+1}\right)$
-$=\sigma\!\left(W_{\ell+1}(h_{\ell}+\alpha v)+b_{\ell+1}\right)$
-$=\tilde{h}_{\ell+1}.$
-Step 4: Infinitely many reparameterizations
-For any invertible matrix $A$ with $A\neq cI$ (i.e., not a scalar multiple of identity), we obtain $v^{\prime}=Av\not\propto v$.
-The space of such matrices has dimension $d^{2}-1$ (excluding scalar multiples), providing infinitely many distinct reparameterizations.
-Conclusion (Weight-space gauge symmetry)
-The symmetries of neural network parameterization imply the existence of infinitely many $(W,v)$ and $(W^{\prime},v^{\prime})$ pairs that yield identical input–output behavior.
-Appendix C Intuitive Explanations
-C.1 Geometric Intuition for Null-Space Ambiguity
-Visual analogy:
-Consider a simplified example where a 3D steering vector $v\in\mathbb{R}^{3}$ affects 2D outputs $o\in\mathbb{R}^{2}$ through a projection matrix $J\in\mathbb{R}^{2\times 3}$.
-By the rank-nullity theorem, the null space $\ker(J)$ is a 1D subspace since $\dim(\ker(J))=3-\text{rank}(J)=3-2=1$.
+To ensure $v^{\prime}_{\beta}\not\propto v$, we need $v+\beta v_{0}\neq cv$ for any scalar $c$, i.e., $\beta v_{0}\neq(c-1)v$, which fails only when $v\in\operatorname{span}(v_{0})$.
+For generic $v$ and $v_{0}\in\mathcal{N}$, this condition holds with probability zero.
+Therefore, for almost all $\beta\neq 0$, the vector $v^{\prime}_{\beta}$ is geometrically distinct from $v$ yet observationally equivalent.
+Conclusion.
+Under the local linear approximation and whenever $\operatorname{rank}(J_{\ell})<d$, there exist infinitely many geometrically distinct steering vectors that are observationally equivalent to $v$.
+Proposition 1 follows.
+$\square$
+C.5 Scope of the linearization
+The proof relies on the approximation in Equation 4.
+This holds exactly for piecewise-linear activations (e.g., ReLU MLPs, on activation-region interiors) and as a first-order approximation for smooth nonlinearities (GeLU, attention softmax) with error $O(\alpha^{2})$.
+For the steering strengths used in our experiments ($\alpha\leq 3$), the nonlinear residual is empirically negligible: the scale-invariance results in Section 6.4 confirm that equivalence holds across the full operationally relevant range without degradation.
+Extending the proof to the exact nonlinear regime would require characterizing the Hessian structure of $F_{\ell\to L}$ and is left for future work.
+Appendix D Weight-Space Gauge Symmetry
+This section records a supplementary observation about gauge symmetry in neural network parameterization.
+It is distinct from the null-space argument of Proposition 1: rather than fixing model weights and asking whether $v$ is recoverable from behavioral observations, we ask whether different weight-vector pairs $(W,v)$ and $(W^{\prime},v^{\prime})$ can produce identical input–output behavior.
+The answer is yes, via the linear reparameterization argument below.
+This symmetry is not realizable as a steering operation on a frozen model since it requires modifying $W_{\ell+1}$.
+We include it because it establishes a complementary non-identifiability result at the level of model parameterization.
+Representation reparameterization.
+Consider an invertible transformation $T:\mathbb{R}^{d}\to\mathbb{R}^{d}$ and define the reparameterized representation $h^{\prime}_{\ell}(x)=T(h_{\ell}(x))$.
+If the subsequent layers satisfy $F_{\ell\to L}(h_{\ell})=F^{\prime}_{\ell\to L}(T(h_{\ell}))$, then the pair $(h_{\ell},v)$ and $(h^{\prime}_{\ell},v^{\prime})$ with $v^{\prime}=DT(h_{\ell})\cdot v$ are observationally indistinguishable.
+We now show this holds for linear $T$.
+Linear reparameterization.
+Restrict to $T(h)=Ah$ where $A\in\mathbb{R}^{d\times d}$ is invertible, so $h^{\prime}_{\ell}(x)=Ah_{\ell}(x)$.
+For layer $\ell+1$ with weight matrix $W_{\ell+1}$ and bias $b_{\ell+1}$, define the reparameterized weights $W^{\prime}_{\ell+1}=W_{\ell+1}A^{-1}$ and $b^{\prime}_{\ell+1}=b_{\ell+1}$.
+The bias is invariant because we restrict to origin-preserving transformations; for affine $T(h)=Ah+c$ the bias transforms as $b^{\prime}_{\ell+1}=b_{\ell+1}-W_{\ell+1}A^{-1}c$, but linear $T$ suffices to establish non-identifiability.
+Under this reparameterization:
+$h^{\prime}_{\ell+1}=\sigma(W^{\prime}_{\ell+1}h^{\prime}_{\ell}+b^{\prime}_{\ell+1})=\sigma(W_{\ell+1}A^{-1}Ah_{\ell}+b_{\ell+1})=\sigma(W_{\ell+1}h_{\ell}+b_{\ell+1})=h_{\ell+1}$ (18)
+so the reparameterized network produces identical activations at layer $\ell+1$ and all subsequent layers.
+Steering under reparameterization.
+Under the original parameterization, steering applies $\tilde{h}_{\ell}=h_{\ell}+\alpha v$, giving $\tilde{h}_{\ell+1}=\sigma(W_{\ell+1}(h_{\ell}+\alpha v)+b_{\ell+1})$.
+Under the reparameterization with $v^{\prime}=Av$, the steered representation at layer $\ell$ is:
+$\tilde{h}^{\prime}_{\ell}=h^{\prime}_{\ell}+\alpha v^{\prime}=Ah_{\ell}+\alpha Av=A(h_{\ell}+\alpha v)=A\tilde{h}_{\ell}$ (19)
+and the next-layer activation is:
+$\tilde{h}^{\prime}_{\ell+1}=\sigma\!\left(W^{\prime}_{\ell+1}(h^{\prime}_{\ell}+\alpha v^{\prime})+b^{\prime}_{\ell+1}\right)=\sigma\!\left(W_{\ell+1}A^{-1}A(h_{\ell}+\alpha v)+b_{\ell+1}\right)=\tilde{h}_{\ell+1}$ (20)
+so the steered activation at layer $\ell+1$ is identical under both parameterizations.
+Infinitely many reparameterizations.
+For any invertible $A$ with $A\neq cI$, the steering vectors $v^{\prime}=Av$ and $v$ are geometrically distinct ($v^{\prime}\not\propto v$) yet produce identical outputs.
+The space of invertible matrices with $A\neq cI$ has dimension $d^{2}-1$, providing infinitely many such reparameterizations.
+Non-identifiability therefore persists even when model weights are not fixed: infinitely many $(W,v)$ and $(W^{\prime},v^{\prime})$ pairs yield identical input–output behavior, confirming that the ambiguity is a structural property of the parameterization rather than a limitation of any particular extraction method.
+Appendix E Mathematical Derivations
+E.1 Dimension of the Observational Equivalence Class
+For a steering vector $v\in\mathbb{R}^{d}$ and Jacobian $J_{\ell}\in\mathbb{R}^{V\times d}$ with rank $r$, two vectors $v$ and $v^{\prime}$ are observationally equivalent if and only if $J_{\ell}v=J_{\ell}v^{\prime}$, i.e., $J_{\ell}(v^{\prime}-v)=0$, i.e., $v^{\prime}-v\in\ker(J_{\ell})$.
+The equivalence class is therefore the affine subspace:
+$v]=\{v^{\prime}\in\mathbb{R}^{d}:J_{\ell}v^{\prime}=J_{\ell}v\}=v+\ker(J_{\ell})$ (21)
+By the rank-nullity theorem, $\dim(\ker(J_{\ell}))=d-r$. Let $\{u_{1},\ldots,u_{d-r}\}$ be an orthonormal basis for $\ker(J_{\ell})$ obtained via SVD of $J_{\ell}$. Then:
+$v]=\left\{v+\sum_{i=1}^{d-r}\beta_{i}u_{i}:\beta_{i}\in\mathbb{R}\right\}$ (22)
+The measure of this set in $\mathbb{R}^{d}$ is:
+$\text{vol}([v])=\int_{\mathbb{R}^{d-r}}d\beta_{1}\cdots d\beta_{d-r}=\infty$ (23)
+so the equivalence class has infinite volume in $\mathbb{R}^{d}$ whenever $r<d$.
+Additionally, $v$ and $cv$ for $c\neq 0$ produce equivalent outputs up to rescaling $\alpha\to c\alpha$, so the equivalence class modulo scaling is a (dr)(d-r)-dimensional projective space $\mathbb{P}^{d-r}$.
+The practical implication is that steering vector recovery is a severely underdetermined inverse problem: there are infinitely many solutions consistent with any finite set of behavioral observations, parameterized by the free coefficients $\beta_{1},\ldots,\beta_{d-r}$.
+To quantify the degree of under-determination, define the non-identifiability ratio:
+$\rho=\frac{d-r}{r}=\frac{\dim(\ker(J_{\ell}))}{\operatorname{rank}(J_{\ell})}$ (24)
+which measures unidentifiable directions per identifiable one. For Llama-3.1-8B with $d=4096$ and $r=57$:
+$\rho_{\text{Llama}}=\frac{4096-57}{57}=\frac{4039}{57}\approx 70.9$ (25)
+For Qwen2.5-3B with $d=2048$ and $r=139$:
+$\rho_{\text{Qwen}}=\frac{2048-139}{139}=\frac{1909}{139}\approx 13.7$ (26)
+In both cases $\rho\gg 1$: the unobservable dimensions vastly outnumber the observable ones. The null space is not a thin subspace but the dominant structure of the ambient geometry.
+E.2 Fisher Information and Cramér-Rao Bound
+We formalize why collecting more behavioral data cannot resolve the null-space ambiguity. Consider the linear Gaussian observation model for a single prompt $x$:
+$o(x)=J_{\ell}(x)v+\eta,\quad\eta\sim\mathcal{N}(0,\sigma^{2}I_{V})$ (27)
+where $o(x)\in\mathbb{R}^{V}$ are the observed output logits and $v\in\mathbb{R}^{d}$ is the parameter to be estimated. The log-likelihood is:
+$\log p(o\mid x,v)=-\frac{1}{2\sigma^{2}}\|o-J_{\ell}(x)v\|^{2}+\text{const}$ (28)
+The score function is $\frac{\partial\log p}{\partial v}=\frac{1}{\sigma^{2}}J_{\ell}(x)^{\top}(o-J_{\ell}(x)v)$, and the Fisher information matrix is:
+$\mathcal{I}(v)=\mathbb{E}_{o}\left[\frac{\partial\log p}{\partial v}\frac{\partial\log p}{\partial v}^{\top}\right]=\frac{1}{\sigma^{2}}J_{\ell}(x)^{\top}J_{\ell}(x)$ (29)
+The Cramér-Rao lower bound states that for any unbiased estimator $\hat{v}$:
+$\text{Cov}(\hat{v})\succeq\mathcal{I}(v)^{-1}=\sigma^{2}\bigl(J_{\ell}(x)^{\top}J_{\ell}(x)\bigr)^{+}$ (30)
+where $(\cdot)^{+}$ is the Moore-Penrose pseudoinverse. For a null-space direction $u_{i}\in\ker(J_{\ell})$, we have $J_{\ell}u_{i}=0$, so:
+$u_{i}^{\top}\mathcal{I}(v)u_{i}=\frac{1}{\sigma^{2}}u_{i}^{\top}J_{\ell}^{\top}J_{\ell}u_{i}=\frac{1}{\sigma^{2}}\|J_{\ell}u_{i}\|^{2}=0$ (31)
+Since $\mathcal{I}(v)$ is singular in the direction $u_{i}$, its pseudoinverse satisfies $u_{i}^{\top}\mathcal{I}(v)^{+}u_{i}=\infty$ (the pseudoinverse assigns zero eigenvalue to $u_{i}$, meaning the bound is vacuously infinite).
+Therefore:
+$\text{Var}(u_{i}^{\top}\hat{v})\geq u_{i}^{\top}\mathcal{I}(v)^{+}u_{i}=\infty\quad\forall\,u_{i}\in\ker(J_{\ell})$ (32)
+Multi-prompt extension.
+Stacking observations across $N$ prompts $\{x_{1},\ldots,x_{N}\}$ gives:
+$\mathbf{o}=J_{\text{stack}}v+\boldsymbol{\eta},\quad J_{\text{stack}}=\begin{bmatrix}J_{\ell}(x_{1})\\ \vdots\\ J_{\ell}(x_{N})\end{bmatrix}\in\mathbb{R}^{(NV)\times d}$ (33)
+with aggregate Fisher information $\mathcal{I}_{N}(v)=\frac{1}{\sigma^{2}}J_{\text{stack}}^{\top}J_{\text{stack}}=\frac{1}{\sigma^{2}}\sum_{i=1}^{N}J_{\ell}(x_{i})^{\top}J_{\ell}(x_{i})$.
+The null space of $\mathcal{I}_{N}(v)$ is:
+$\ker(\mathcal{I}_{N})=\ker(J_{\text{stack}})=\bigcap_{i=1}^{N}\ker(J_{\ell}(x_{i}))$ (34)
+Any direction $u\in\bigcap_{i}\ker(J_{\ell}(x_{i}))$ satisfies $J_{\ell}(x_{i})u=0$ for all $i$, meaning it contributes zero information regardless of $N$.
+If the intersection of null spaces is non-trivial — which empirically holds when all prompts probe the same effective output subspace — then $\ker(\mathcal{I}_{N})\neq\{0\}$ for all $N$ and the Cramér-Rao bound remains infinite along those directions.
+Non-identifiability is therefore not a small-sample problem: it is an intrinsic property of the model geometry that no amount of behavioral data can resolve.
+Appendix F Intuitive Explanations
+F.1 Geometric intuition for null-space ambiguity
+Consider a simplified setting where a 3D steering vector $v\in\mathbb{R}^{3}$ affects 2D outputs $o\in\mathbb{R}^{2}$ through a projection matrix $J\in\mathbb{R}^{2\times 3}$.
+By the rank-nullity theorem, $\dim(\ker(J))=3-2=1$.
+Concretely, if $J=[I_{2}\mid 0]$ projects onto the first two coordinates, then $\ker(J)=\operatorname{span}\{(0,0,1)^{\top}\}$.
+Any vector of the form $v^{\prime}=v+\beta(0,0,1)^{\top}$ produces identical outputs since:
+$J(v+\beta e_{3})=Jv+\beta Je_{3}=Jv+\beta\cdot 0=Jv$ (35)
+The output depends only on the projection of $v$ onto $\operatorname{row}(J)$, while the component in $\ker(J)$ is entirely invisible.
+Figure 8 illustrates this geometry: the observable output space sees only the $(v_{1},v_{2})$ projection, while $v_{3}\in\ker(J)$ leaves no trace in any output.
+In language models, $\ker(J_{\ell})$ occupies 86–94% of the ambient activation space, so almost any perturbation to a steering vector goes undetected by behavioral evaluation.
 […]
-Figure 5: Geometric intuition for null-space ambiguity.
-The output depends only on the $(v_{1},v_{2})$ projection (row space of $J$), while $v_{3}\in\ker(J)$ is invisible; hence $J(v+\alpha v_{3})=Jv$.
-For concreteness, suppose $J=[I_{2}\mid 0]$ projects onto the first two coordinates.
-Then $\ker(J)=\text{span}\{(0,0,1)\}$ is the $v_{3}$ axis.
-The key insight generalizes: for any projection matrix $J$, directions in $\ker(J)$ are invisible to outputs.
-Figure 5 illustrates this geometric intuition.
-Any vector of the form $v^{\prime}=v+\alpha v_{3}$ for $\alpha\in\mathbb{R}$ produces identical outputs:
-$J(v+\alpha v_{3})$ $=Jv+\alpha Jv_{3}$ (30)
-$=Jv+\alpha\cdot 0\quad(v_{3}\in\ker(J))$
-$=Jv.$
-Since $\alpha$ can take infinitely many values and $v^{\prime}\not\propto v$ for generic choices of $\alpha$, there exist infinitely many geometrically distinct steering vectors that are observationally equivalent.
-C.2 Why More Data Doesn’t Help
-Common misconception: "If we collect more steering examples, we can pin down the unique vector."
-Why this fails:
-Consider observing steering effects on $N$ prompts $\{x_{1},\ldots,x_{N}\}$.
-More fundamentally, any $v_{0}\in\ker(J_{\ell})$ satisfies $J_{\ell}v_{0}=0$ by definition since this is a property of the model weights and not the prompt distribution.
-Stacking observations across prompts therefore cannot remove directions from $\ker(J_{\ell})$.
-Each observation provides:
-$o_{i}=J_{\ell}(x_{i})v+\eta_{i}$ (31)
-Stacking these:
-$\begin{bmatrix}o_{1}\\ o_{2}\\ \vdots\\ o_{N}\end{bmatrix}=\begin{bmatrix}J_{\ell}(x_{1})\\ J_{\ell}(x_{2})\\ \vdots\\ J_{\ell}(x_{N})\end{bmatrix}v+\begin{bmatrix}\eta_{1}\\ \eta_{2}\\ \vdots\\ \eta_{N}\end{bmatrix}$ (32)
-The stacked Jacobian $J_{\text{stack}}\in\mathbb{R}^{(N\cdot V)\times d}$ has null space:
-$\ker(J_{\text{stack}})=\bigcap_{i=1}^{N}\ker(J_{\ell}(x_{i}))$ (33)
+Figure 8: Geometric intuition for null-space ambiguity. The output depends only on the $(v_{1},v_{2})$ projection (row space of $J$), while $v_{3}\in\ker(J)$ is invisible to all outputs.
+F.2 Why more data does not resolve the ambiguity
+A natural conjecture is that collecting diverse prompts could reduce the null space by probing different output dimensions.
+Each observation provides $o_{i}=J_{\ell}(x_{i})v+\eta_{i}$, and stacking $N$ observations gives a stacked Jacobian:
+$J_{\text{stack}}=\begin{bmatrix}J_{\ell}(x_{1})\\ \vdots\\ J_{\ell}(x_{N})\end{bmatrix}\in\mathbb{R}^{(NV)\times d}$ (36)
+with null space $\ker(J_{\text{stack}})=\bigcap_{i=1}^{N}\ker(J_{\ell}(x_{i}))$.
+In principle, diverse prompts that probe different output dimensions could reduce this intersection.
+However, directions $v_{0}$ satisfying $J_{\ell}(x)v_{0}=0$ for all $x$ are determined by model weights, not the prompt distribution.
+When all prompts probe the same effective output subspace, their Jacobians share a common null space and the intersection does not shrink regardless of how many prompts are added.
+Figure 9 illustrates this: each prompt measures the same active dimensions, leaving the null-space intersection unchanged.
+The Fisher information along these null-space directions is identically zero, so by the Cramér-Rao bound the estimation variance is unbounded.
+Non-identifiability is not a small-sample artifact that more data can resolve — it is an intrinsic property of the model geometry.
 […]
-Figure 6: Illustration of why prompt diversity does not resolve null-space ambiguity.
-Each prompt induces a Jacobian that measures the same effective subspace, leaving the null-space intersection unchanged.
-Critical observation: If all $J_{\ell}(x_{i})$ share a common null space (e.g., all prompts probe similar aspects of the model), then:
-$\ker(J_{\text{stack}})=\ker(J_{\ell}(x_{1}))\neq\{0\}$ (34)
-This means adding more prompts does not reduce the null space when all prompts probe the same effective subspace—the ambiguity persists. Figure 6 illustrates this phenomenon.
-C.3 Why Orthogonal Perturbations Preserve Semantics
-Setup: We test steering with $v$ versus $v+v_{\perp}$, where $v_{\perp}\perp v$ and $\|v_{\perp}\|=1$.
-Empirically, adding a random orthogonal direction produces nearly equivalent semantic effects.
-Why?
-Decomposing the perturbation
-Any perturbation $v_{\perp}$ can be decomposed into observable and invisible components:
-$v_{\perp}=v_{\perp,\text{row}}+v_{\perp,\text{null}}$ (35)
-where $v_{\perp,\text{row}}\in\operatorname{row}(J)$ (observable through outputs) and $v_{\perp,\text{null}}\in\ker(J)$ (invisible to outputs).
-For a random $v_{\perp}\perp v$, the expected fraction in the null space is:
-$\mathbb{E}[\|v_{\perp,\text{null}}\|^{2}]\approx\frac{\dim(\ker(J))}{d}\approx 0.20{-}0.25$ (36)
-This means $\sim$20–25% of the perturbation is automatically invisible to model outputs.
-Observable effect is diffuse and unstructured
-The actual output change from adding $v_{\perp}$ is:
-$J(v+v_{\perp})=Jv+Jv_{\perp,\text{row}}$ (37)
-While $Jv_{\perp,\text{row}}$ is not necessarily small in norm, it represents a random direction in the $\sim$3200-dimensional row space of $J$.
-In contrast, $Jv$ is a structured, semantically coherent steering effect (e.g., consistently shifting outputs toward “honesty” or “sycophancy”).
-The key distinction is not magnitude but semantic coherence: $Jv$ produces aligned, directional changes across the vocabulary, while $Jv_{\perp,\text{row}}$ produces diffuse, incoherent perturbations that average out across tokens and do not systematically shift semantic meaning in any consistent direction.
-Analogy: If $Jv$ is a strong wind blowing consistently north, then $Jv_{\perp,\text{row}}$ is turbulence—it may have comparable energy but lacks directional coherence, so the overall trajectory remains northward.
-Appendix D Prompt Construction Details
-We construct contrastive prompt pairs for each semantic trait using template-based generation. Table 4 provides concrete examples for all three traits used in the empirical validation.
-Table 4: Example contrastive prompt pairs for each semantic trait.
+Figure 9: Why prompt diversity does not resolve null-space ambiguity.
+Each prompt probes the same effective subspace, leaving the null-space intersection unchanged.
+Numbers reflect Qwen2.5-3B ($d=2048$, rank 139).
+F.3 Why orthogonal perturbations preserve semantic coherence
+Any perturbation $v_{\perp}\perp v$ decomposes as $v_{\perp}=v_{\perp,\text{row}}+v_{\perp,\text{null}}$, where $v_{\perp,\text{row}}\in\operatorname{row}(J_{\ell})$ is observable through outputs and $v_{\perp,\text{null}}\in\ker(J_{\ell})$ is invisible.
+For a random $v_{\perp}$, the expected fraction landing in the null space is approximately $\dim(\ker(J_{\ell}))/d\approx 0.86$$0.94$, meaning the bulk of any random perturbation is automatically undetectable.
+The remaining row-space component contributes an output change $J_{\ell}v_{\perp,\text{row}}$, but this is a random direction in the row space rather than a structured semantic shift.
+The output change from the original steering vector, $J_{\ell}v$, produces aligned and directional changes across the vocabulary that consistently shift the target trait.
+The row-space component of the perturbation, by contrast, produces diffuse and incoherent changes that average out across tokens without pushing any semantic dimension in a consistent direction.
+To make the distinction concrete: if $J_{\ell}v$ is a strong wind blowing consistently northward, then $J_{\ell}v_{\perp,\text{row}}$ is turbulence.
+It may carry comparable energy but lacks directional coherence so the overall trajectory remains northward.
+This is why orthogonal perturbations leave semantic classifier scores nearly unchanged even when their norm is comparable to the steering vector itself.
+Appendix G Prompt Construction Details
+We construct contrastive prompt pairs for each semantic trait using template-based generation.
+Table 6 provides concrete examples for all five traits used in the empirical validation.
+Held-out evaluation prompts are drawn from a separate set of 50 prompts not used in vector extraction, covering in-distribution, topic-shift, and genre-shift conditions for the multi-environment experiment.
+Table 6: Example contrastive prompt pairs per trait.
 Trait Positive ($x^{+}$) Negative ($x^{-}$)
@@
 Politeness Respond diplomatically to criticism from a colleague. Reply bluntly to someone who criticized you.
-Humor Tell a lighthearted story about an embarrassing mistake. Provide a factual account of an embarrassing mistake.
-Figure 7: Steering extraction and evaluation pipeline.
+Sentiment Write an enthusiastic review of a product you loved. Write a negative review of a product you disliked.
+Truthfulness Provide an accurate factual summary of this event. Make up plausible-sounding details about this event.
+Agreeableness Respond warmly and supportively to a colleague’s proposal. Push back sharply on a colleague’s proposal.
 Figure 10 illustrates the pipeline used to extract and evaluate steering vectors.
@@
 This vector is then applied to a separate set of held-out evaluation prompts to generate steered outputs.
-Appendix E Mathematical Derivations
-E.1 Dimension of Observational Equivalence Class
-Setup: For steering vector $v\in\mathbb{R}^{d}$ and Jacobian $J\in\mathbb{R}^{V\times d}$ with rank $r$, consider:
-$v]=\{v^{\prime}\in\mathbb{R}^{d}:Jv^{\prime}=Jv\}$ (38)
-Question: How many degrees of freedom exist in the equivalence class of observationally equivalent steering vectors?
-Analysis: The equivalence class is an affine subspace:
-$v]=v+\ker(J)$ (39)
-where $\ker(J)$ is a (dr)(d-r)-dimensional linear subspace.
-Parameterization: Let $\{u_{1},\ldots,u_{d-r}\}$ be an orthonormal basis for $\ker(J)$. Then:
-$v]=\left\{v+\sum_{i=1}^{d-r}\alpha_{i}u_{i}:\alpha_{i}\in\mathbb{R}\right\}$ (40)
-The equivalence class has (dr)(d-r) degrees of freedom.
-Scaling ambiguity: Additionally, $v$ and $cv$ produce equivalent outputs (up to rescaling $\alpha$). The equivalence class modulo scaling is a (dr)(d-r)-dimensional projective space.
-Illustrative example. Suppose a model has hidden size $d=4096$ and an effective Jacobian rank $r\approx 3100$. Then the equivalence class dimension is $d-r\approx 996$.
-Interpretation: For every identifiable parameter (row-space direction), there are $996/3100\approx 0.32$ unidentifiable parameters (null-space directions).
-The under-determination ratio is approximately $1:3$.
-E.2 Fisher Information and Cramér-Rao Bound
-Setup: Consider the statistical model:
-$o(x;v,\alpha)\sim p(o|x,v,\alpha)$ (41)
-where $v$ is the parameter to estimate.
-Question: Can we achieve better identifiability by collecting more samples?
-Fisher Information Matrix: For the linear Gaussian model $o=Jv+\eta$ with $\eta\sim\mathcal{N}(0,\sigma^{2}I)$:
-$\mathcal{I}(v)$ $=\mathbb{E}_{o}\Bigg[\left(\frac{\partial\log p(o\mid x,v,\alpha)}{\partial v}\right)$ (42)
-$\qquad\left(\frac{\partial\log p(o\mid x,v,\alpha)}{\partial v}\right)^{\!\top}\Bigg]$
-$=\frac{1}{\sigma^{2}}J^{\top}J.$
-Cramér-Rao Lower Bound: The covariance of any unbiased estimator $\hat{v}$ satisfies:
-$\text{Cov}(\hat{v})\succeq\mathcal{I}(v)^{-1}=\sigma^{2}(J^{\top}J)^{+}$ (43)
-where $(J^{\top}J)^{+}$ is the Moore-Penrose pseudo-inverse.
-Implications: For null-space directions $u_{i}\in\ker(J)$, the Fisher information is degenerate:
-$u_{i}^{\top}(J^{\top}J)u_{i}=0$ (44)
-Therefore, the variance of any unbiased estimator along null-space directions is unbounded:
-$\text{Var}(u_{i}^{\top}\hat{v})\geq u_{i}^{\top}\mathcal{I}(v)^{-1}u_{i}=\infty$ (45)
-Conclusion: The Cramér-Rao bound is infinite for null-space components, confirming that no finite amount of data can resolve the ambiguity.
-Non-identifiability is fundamental, not a small-sample problem.
-More data cannot help because the information geometry has infinite uncertainty in null-space directions.
+Figure 10: Steering extraction and evaluation pipeline.